### v1

# Robotic Sorting Systems: Robot Management and Layout Design Optimization

Abstract
In the contemporary logistics industry, automation plays a pivotal role in enhancing production efficiency and expanding industrial scale. Autonomous mobile robots, in particular, have become integral to the modernization efforts in warehouses. One noteworthy application in robotic warehousing is the robotic sorting system (RSS), distinguished by its characteristics such as cost-effectiveness, simplicity, scalability, and adaptable throughput control. While previous research has focused on analyzing the efficiency of RSS, it often assumed an ideal robot management system ignoring potential queuing delays by assuming constant travel times. This study relaxes this assumption and explores the quantitative relationship between RSS configuration parameters and system throughput. We introduce a novel robot traffic management method, named the rhythmic control for sorting scenario (RC-S), for RSS operations, equipped with an estimation formula establishing the relationship between system performance and configurations. Simulations validate that RC-S reduces average service time by 10.3% compared to the classical cooperative A* algorithm, while also improving throughput and runtime. Based on the performance analysis of RC-S, we further develop a layout optimization model for RSS, considering RSS configuration, desired throughput, and costs, to minimize expenses and determine the best layout. Numerical studies show that at lower throughput levels, facility costs dominate, while at higher throughput levels, labor costs prevail. Additionally, due to traffic efficiency limitations, RSS is well-suited for small-scale operations like end-of-supply-chain distribution centers. Keywords: Logistics; Robotic sorting system; Robot management system; Performance evaluation; Layout design
1 Introduction
Improved logistics and delivery services have fueled the rapid growth of e-commerce in the 21st century. The development of stable supply chains, same-day or next-day delivery options, and hassle-free return policies increase consumer confidence in online shopping (BusinessWire,, 2023). In order to handle the increasing volume of online orders, a new generation of warehouses specifically catering to individual customers has become a hot topic for logistics companies. This kind of warehouse efficiently meets the demand for small orders with tight delivery schedules through the implementation of automated equipments (Boysen et al., 2019b, ). For instance, robotic mobile fulfill system (RMFS), which employs a rack-moving mechanism, is widely used in intra-warehouse logistics, such as Amazon’s KIVA systems (Wurman et al.,, 2008). The main activities performed in warehouse include: (1) receiving, (2) transfer and put away, (3) order picking/selection, (4) accumulation/sorting, (5) cross-docking, and (6) shipping (De Koster et al.,, 2007).
In this paper, we focus on the sorting process. In a distribution center or warehouse where batching and/or zoning policies are implemented within the order picking process, sorting is required to split the batch and consolidate items per customer order or destination for shipping. Conventional sorting systems commonly employ sorting conveyors equipped with different sorter technologies. The sorters above move steadily, sequentially pass through each outbound station, and release the loaded parcels at suitable locations (Boysen et al., 2019a, ). These systems are highly appreciated for its efficiency and stability, while suffering from the inflexibility and significant space occupation (Boysen et al.,, 2023).
(a) A dual-layer RSS solution
(b) Conveyor-belt sorting robot
Figure 1: The application of RSS (source: www.geekplus.com)
A novel sorting system utilizing mobile robots attracts increasing attention in recent years, which is called the robotic sorting system (RSS) (Zou et al.,, 2021). Figure 1(a) presents a dual-layer implementation of RSS. The site is dotted with considerable number of outlets, each of which is associated with a specific sorting category, gathering all the packages of that category. The hardware for the outlet can be a bin, a cage cart, or a chute on the floor, which connects to the packing equipment on the lower level. The loading stations are located on the periphery of the site, where workers are responsible for loading parcels onto the robots. Sorting robots are the devices responsible for carrying out the sorting operations, serving as the conveyor belts in conventional sorting systems. A top-mounted tray or conveyor device allows the robot to load or unload shipments. Figure 1(b) shows a specific model of sorting robot developed by Geekplus, a global technology company specialized in smart logistics and robots. The operational cycle of a sorting robot is as follows:
  1. 1.
    The robot receives a parcel at a loading station, along with the destination information;
  2. 2.
    The robot moves along aisles to the designated outlet;
  3. 3.
    When reaching the target outlet, the robot releases the loaded parcel;
  4. 4.
    The robot returns to the waiting zone behind a loading station and becomes idle.
The independent and modular nature of robotic operations enables the system design process to become more flexible. During off-hours break, warehouse managers can easily modify the control software to change the positions of loading stations, correspondences between outlets and destinations, and the topology of the aisle network (Xu et al.,, 2022). Compared with conveyor belt equipment, RSS possesses the advantage of scalability and throughput flexibility, which are highly suited to the high-demand and dynamic nature of the current e-commerce environment (Azadeh et al.,, 2017). To promote this advantage, robotics companies introduce an innovative business model called robots-as-a-service, in which robotics companies offer customers the opportunity to rent robots for use in manufacturing. As a result, the customers no longer have to contend with the high costs typically associated with purchasing advanced automation tools outright. In RSS, warehouse manager can adjust the number of robots based on sorting demand fluctuations, thereby reducing operations costs.
However, the flexibility of robot system also brings additional operational challenge, namely the multi-agent path finding (MAPF) problem, which aims to resolve the potential conflicts among the trajectories of robots. The MAPF problem has been proven to be NP-hard (Yu and LaValle,, 2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in the instability of RSS efficiency. Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in an RSS, particularly when dealing with a substantial parcel flow. Existing studies often overlook issues related to robot coordination, making the idealized assumption that there are no conflicts or deadlocks. Additionally, these studies tend to assume a fixed route travel time excluding the potential queuing delays. As a result, their outcomes may become distorted, especially in situations of high traffic load, as the actual travel time of the robots deviates from the free-flow travel time.
Given the existing gap, this paper aims to address critical challenges in designing an efficient RSS, focusing on three key areas. First, we seek to ensure the system operates both stably and efficiently in a setting where multiple robots collaborate, a non-trivial challenge given the complexities of robot interaction. Second, the development of a precise model to quantify congestion effects within the RSS is essential. Such a model will enable accurate throughput estimation, which is vital for understanding the relationship between the size of the facility and its sorting capacity. Third, we emphasize the need for management strategies that exploit the flexibility of RSS to adapt to fluctuating demands cost-effectively. These targeted areas of improvement aim to enhance the overall performance and scalability of RSS implementations.
To address the aforementioned challenges and achieve the system optimum, we first explore the potential of centralized control strategy for sorting robots. To fully utilize the global information of the robot control system, taking into account robot operational patterns, and safety distance regulations, we adopt a network-level control strategy for the online coordination of autonomous vehicles and propose a collision-free spatio-temporal path planning method for RSS. Compared to existing MAPF methods, it does not require high communication frequency and tight latency between the central host and each robot. Additionally, we derive an efficient heuristic algorithm to enhance system scalability. Second, building upon this control strategy, under any given system configurations, we conduct a quantitative assessment of the average travel distance of robots and the maximum throughput of the RSS. We introduce an estimation formula that elucidates the relationship between system performance metrics and configuration settings. Finally, we propose an optimization model aimed at minimizing the initial investment and average operations costs under variable demand. The solution includes layout design and resource allocation recommendations. Experiments are conducted to provide several insights about the RSS, such as the cost structure of the RSS across varying unit price conditions. To the best of our knowledge, this paper is among the first to incorporate the traffic issues of robots during operations into the configuration design of the RSS system. Our intention is to bridge operational and strategic planning in multi-robot coordination contexts, and hence contribute a comprehensive theoretical framework to support decision-making in such complex systems.
The remainder of this paper is organized as follows. In section 2, we provide an extensive review of relevant research on RSS systems and conventional sorting systems, highlighting the contributions of this research. Section 3 offers a detailed problem description. In section 4, we introduce an innovative traffic management framework to coordinate multiple robots in a warehouse setting. Based on the proposed framework, section 5 presents a system throughput estimation formula, which provides quantitative measurement of layout configuration’s impact to system efficiency. In section 6, we demonstrate the superiority of our proposed traffic management framework compared with benchmarks, as well as validate the throughput estimation formula, through comprehensive experiments. Moving on to section 7, we propose a layout optimization model to minimize the total costs and introduce an efficient solution approach. Section 8 presents a sensitivity analysis to investigate the optimal layout design and the corresponding cost structure under different unit cost, and insights distilled from the results. Finally, section 9 provides a summary of the entire paper.
2 Literature Review
The cost savings in labor through automation are particularly crucial in the picking process, and higher sorting depth enables parcel companies to offer convenient and low-cost same-day delivery services (Dekhne et al.,, 2019). However, the operations research community has scarcely explored the realm of robotized sorting systems, which represent the latest advance of warehouse automation. Among the existing literature, studies mainly focus on small parcel sorting systems in distribution centers. Boysen et al., (2023) conducted research on order sorting system, which involves additional tasks including Piece-to-Order assignment and Order-to-Collection Point assignment. Therefore, the primary focus of the referred study is not solely on scheduling and control of robots, and the travel time of robots to a specific outlet is assumed to be constant.Shi et al., (2021)’s study, on the other hand, investigates a human-robot hybrid sorting system, in which the scale of robot sorting is fixed, while the manual capacity is adjustable according to the demand. Regrettably, in this context, the considerable flexibility inherent in the robotic system has not been fully harnessed.
In addition to the two aforementioned references, the remaining literature on robotic sortation, including this study, reorients the focus back to the original RSS itself. This body of work delves into the comprehensive assessment of transportation efficiency within the robot fleet under various system configurations. In the majority of these studies, robots’ sorting tasks are modeled as services in queuing network models, and the total throughput is the key performance evaluation metrics (Zou and Chen,, 2020; Zi and Gao, 2020a, ; Zi and Gao, 2020b, ; Zou et al.,, 2021; Xu et al.,, 2022). Zi and Gao, 2020a ; Zi and Gao, 2020b assumed that each robot moves independently. Zou and Chen, (2020) introduced a semi-open queuing network (SOQN) model and subsequently improved its theoretical aspects in their follow-up research (Zou et al.,, 2021). The SOQN models in these two studies take into account the queuing behavior of robots at loading stations and outlets, partially addressing the issue of robot occupancy on the aisle. Furthermore, their second study compares how the minimum total costs of the system varied under different network topology and different unit cost of system components, given the target sorting efficiency. Xu et al., (2022) further investigated an RSS with a parcel-to-loading-station assignment mechanism, which could be regarded as a pre-sorting strategy. To simplify the model, the study removes the constraint on the number of robots, resulting in an open queuing network (OQN) model. Experimental results demonstrates that the introduction of pre-sorting effectively reduces the average travel distance of the robots but might lead to extra congestion upstream at the same time. Additionally, Liu et al., 2019b and Tan et al., (2021) focused on the task assignment problem with the objective of minimizing the sorting makespan. Both adopts a travel time model (closed-form travel time expression, defined by Azadeh et al., (2017)), and the second research formulates a mixed-integer programming model. Compared to queuing network models, travel time models exhibit significant distortion when the number of robots is high, as they allow multiple robots to simultaneously occupy a facility.
Table 1: Literature summary on RSS investigation
Reference System Modeling Objective Congestion & Deadlock
Liu et al., 2019b parcel sortation TT, MIP MS, C $\times$
[…]
Zi and Gao, 2020a parcel sortation Q T $\times$
Zi and Gao, 2020b parcel sortation Q, MIP T, C $\times$
[…]
Tan et al., (2021) parcel sortation TT, MIP MS $\times$
Xu et al., (2022) parcel sortation OQN T $\times$
Boysen et al., (2023) order sortation TT, MIP T $\times$
This paper parcel sortation TT, MIP T, TC $\checkmark$
[…]
Throughout the existing research, a limited number of studies consider the floor space cost while others relax the constraints on aisle resources. Furthermore, finding collision-free paths for multiple vehicles has been proved to be an NP-hard problem (Surynek,, 2010), and some state-of-the-art algorithms (such as the CROTPN-based model in Sun et al.,, 2018) also struggle to completely eliminate the issue of road congestion resulting from an increasing number of robots. However, except the study by Shi et al., (2021), which employs a fitting method to predict the positive correlation between traffic flow and congestion, most of the studies neglect the mutual influence among multiple robots in aisle occupation for a compact system, leading to an overestimation of aisle capacity. Table 1 presents a comprehensive overview of the literature reviewed on the RSS, highlighting its relevance to our paper.
Next, we review the performance analysis and layout design of sorting systems based on conveyors. The study by Russell and Meller, 2003a comprehensively investigates the layout design problem in an order fulfillment system. Their descriptive model analyzes the impact of the number of workers and workstations on the system’s capacity, considering the costs of labor and equipment, as well as providing appropriate system configuration based on target throughput requirement. The study also validates their analytical results through simulation experiments and concluded the advantages of conveyor belt sorting compared to manual sorting. Bartholdi and Gue, (2000) studied the scenario of a cross-docking terminal, where freight needs to undergo sorting on a dragline conveyor system from strip doors to stack doors. They accounted for three types of congestion in the system and employed a simulated annealing method to examine sorting efficiency and labor costs across various layouts. Besides, the postal service industry faces greater challenges, including higher volume of shipments and tighter time window. Fedtke and Boysen, (2017) examined how design alternatives within the distribution center affect the performance of sorting system. The authors devised an integer programming model to optimize the minimum service distance under various configurations. Additionally, they developed heuristic algorithms for solutions. Similar methodologies have been applied in several other researches (Briskorn et al.,, 2017; Zenker and Boysen,, 2018). Readers are recommended to refer to Boysen et al., 2019a for an in-depth understanding of sortation conveyor systems.
In addition to sortation systems, there are a sufficient number of studies on RMFS system in modern order picking process, which also consider the multi-robot management problem. An important operational metric of order picking is the proportion of effective working time for workers, and thus, related studies often utilize queuing network models to characterize the uncertainty in RMFS. Lamballais et al., (2017) proposed a SOQN model and provided insights into how performance is affected by factors such as the length-to-width ratio, workstation locations, number of robots. This modeling approach effectively captures the system bottleneck associated with robot blocking at workstations, which is validated through simulation. This modeling approach has proven instrumental in addressing challenges related to battery recovery, robot assignment, inventory allocation, and multi-deep layout (Zou et al.,, 2018; Roy et al.,, 2019; Tessensohn et al.,, 2020; Yang et al.,, 2022). Certain studies have relaxed the constraint on the number of robots, depending on the focus of their research (open queuing network: Yuan and Gong,, 2017, closed queuing network: Nigam et al.,, 2014; Roy et al.,, 2019). However, some researchers have recognized the limitations of queuing models in capturing the impact of aisle design on system throughput, and have chosen to address the problem from the perspective of MAPF problem in their studies. Wagner and Choset, (2011) proposed a novel M* algorithm that allowed for the planning of a larger number of robot trajectories compared to the original A* algorithm. However, they faced challenges due to significant memory requirements and time complexity, and could not guarantee a solution. Some researchers designed priority-based algorithms and integrated task assignment to reduce waiting time (Nguyen et al.,, 2019; Liu et al., 2019a, ), sacrificing algorithm optimality in exchange for stability and computational efficiency. Li et al., (2021) introduced an RHCR algorithm framework to tackle the windowed MAPF problem, providing an efficient solution for managing newly arrived orders in warehouse operations. The algorithm demonstrates scalability up to 1,000 agents and yields high-throughput solutions. However, despite their effectiveness, MAPF-based studies have encountered challenges in deriving a closed-form expression for throughput estimation, a critical element for robust system configurations. In summary, there are few studies on RMFS that comprehensively address both operational robot management problems and strategic warehouse design, with considering their interplay. We aim for the findings and framework of this study not only to be applicable to the RSS system, but also to be extendable to these multi-robot systems in future studies.
3 Problem Description
Figure 2: Layout of RSS
In an RSS, three zones are deployed from center to periphery in the field (Wang et al.,, 2021): 1) sorting zone; 2) loading zone; 3) waiting zone, as shown in Figure 2. Parcels are loaded to robots by workers in the loading zone. Robots drop off each parcel to its target outlet in the sorting zone to fulfill a sorting task. Waiting zone is the space where robots idle and queue behind loading stations. The area of every zone is determined in the design phase, and remains unchanged regardless of sorting demands, number of robots and workers. Let $W_{l}$ and $W_{w}$ denote the width of loading zone and waiting zone, respectively, which are considered to be constant in this study. The decision of site planning mainly focuses on the design of sorting zone, particularly the configuration of the aisle network.
Without loss of generality, we make the following assumptions in this study:
  • The parcels arrive in batch manner. When a parcel is loaded on a robot, the next parcel will arrive at the loading station immediately. That is, the loading time is negligible, and this parameter is considered implicitly in the loading rate of worker.
  • The arrival rate of parcels destinating for each outlet at each loading station is the same [^2] When the sorting demand is unevenly distributed, this assumption can be achieved by increasing or decreasing the number of outlets corresponding to each destination.;
  • The time for robots dropping parcels into outlets is negligible;
  • Each outlet allows at most four robots in different directions releasing parcels simultaneously, and the drop-off operation by robots around one outlet does not affect each other;
  • Each robot can only carry one parcel during one delivery process;
  • Workers are evenly distributed in the loading zone;
  • The transportation of parcels after being collected in the outlet is not considered in this research, which may affect the operations stage for RSSs with different physical structure; samples of single-tier and two-tier design are provided by Zou et al., (2021);
  • The battery charging needs of robots are not considered.
(a) Sample gird layout of a compact $4+4$ aisle network
(b) Topology of sample network
Figure 3: Illustration of aisle network in sorting zone
To better evaluate the performance of this system with high loads and complicated sorting demands, the microscopic traffic management of robots in the sorting zone needs careful investigation. Notably, we use the term ”aisle” instead of ”road” in the following content to clarify that each link in the network is single-lane. A sample grid network with compact aisle layout is depicted in Figure 3(a). It consists of $n_{v}$ vertical aisles and $n_{h}$ horizontal aisles. Unit cells are evenly distributed along each aisle and serve for robot localization (in some robotic warehouses, distinct unit cells can be identified by QR codes). The unit cell is a square with a side length of $D$. We assume the length of a robot is not greater than $D$. It allows robot to spin turn in a grid. Outlets collect the parcels dropped by robots. Outlets are evenly distributed over the sorting zone and divided by aisles, and each outlet occupies one grid. This layout ensures that robots move along aisles can directly drop parcels into outlets without additional turning, while also achieving sufficient density. Figure 3(b) shows the topology of sample layout in Figure 3(a). Conflict nodes (blue dot) connect the vertical and horizontal aisles, and unloading nodes (red hexagon) are the points where robots release parcels. To improve traffic efficiency, all aisles are one-way and the directions of any two adjacent aisles are opposite, which is a common strategy in real. To ensure the connectivity of the map, each loading station in loading zone has independent entrance and exit, and the number of horizontal and vertical aisles should be even. Hence, the number of loading stations is no more than the number of aisles.
Figure 4: Common types of conflicts: (a) a swapping conflict, (b) a vertex conflict, (c) a following conflict, and (d) a cycle conflict
The sorting robots share the same kinematic parameters, featuring a maximum velocity of $v_{max}$ and maximum acceleration/deceleration of $c_{max}$. Additionally, these robots possess the capability of in-place rotation, characterized by an average angular velocity denoted as $\omega_{r}$, with no consideration for turning acceleration or deceleration. To ensure collision-free movement of robots, the robot management system needs to avoid conflicts between different robots during path finding. Common conflicts, as illustrated in Figure 4, include the following four types (Stern et al.,, 2019): (a) Swapping conflict occurs when two robots are planned to swap locations at the same time; (b) Vertex conflict occurs when more than two robots are planned to occupy the same vertex; (c) Following conflict occurs when one robot is planned to occupy a vertex that was occupied by another robot in the previous timestep; (d) Cycle conflict occurs among multiple robots, with each robot having a following conflict with another robot. Among these conflicts, both swapping conflicts and cycle conflicts necessitate a re-planning of paths. The occurrence of either can result in significant losses in system efficiency.
Planing problems related to RSS can be divided into two stages: the site planning stage and operations stage. In the site planning stage, decision maker concentrates on the division of zones for different functions in the warehouse. The area of an RSS should be large enough to fulfill the designated throughput level in peak season, and the number of outlets should be no fewer than the number of individual sorting destinations. The decision variables of site planning stage are the number of horizontal and vertical aisles $n_{h},n_{v}$ and the number of loading stations $n_{l}$; in the operations stage, decision maker should adaptively adjust the number of workers and robots to minimize operations costs, given real-time sorting demands. As a warehouse is usually rented for years, it is advisable to keep it unchanged during this stage. The decision variables of operations stage are the number of workers and robots in peak season $n_{w}^{H},n_{r}^{H}$ and in off-peak season $n_{w}^{L},n_{r}^{L}$, where $L$ and $H$ in superscript indicate that the sorting demand is at low and high levels, respectively; and $w$ and $r$ in subscript represent worker and robot, respectively.
The goal of RSS planning problem is to minimize the total costs, which consists of two parts: the facility costs $C_{f}$, and the operations costs $C_{o}$. The size of the site holds a trade-off between $C_{f}$ and $C_{o}$: With increased demand, a small sorting zone may lead to congestion, thereby increasing the service time for robots and raising the rental cost of robots in operations costs; whereas, expanding the size of RSS would escalate the facility costs. In sections 4-6, we will respectively present the traffic management of robots, conduct system performance analysis under the proposed traffic management framework, and validate the proposed algorithm and performance estimation formula. In section 7, a layout design optimization model is formulated to minimize the total costs of RSS with given demand.
4 Management of High-density Robot Traffic in RSS
Managing a large number of robots within a dense network poses significant challenges, especially when striving to balance system throughput performance with the computational load of algorithms under unpredictable traffic demands. In order to address multi-robot path finding and eliminate traffic gridlock within an admissible computational duration, this study adopts an innovative autonomous vehicle management scheme, called Rhythmic Control (RC) (Lin et al.,, 2021). It enables the planning of uninterrupted scheduling for robot fleets. In this section, we start with a motivation example to illustrate why introducing rhythm into robot management is necessary. Subsequently, we summarize the key concepts inherited from the original RC. By leveraging the node-reservation idea in classical MAPF solver, we propose a new centralized control framework for sorting scenario, i.e., the RC-S scheme. RC-S presents high orderliness in practical performance, serving as the foundation for theoretical analysis of system efficiency, which will be discussed in detail in section 5.
4.1 Motivation example
In conveyor-based sorting systems, there is always a mainline through which goods move. When passing through the target outlets, the goods are pushed out by barriers to complete the sorting process. We aim to combine the stability of conveyor-based sorting systems with the flexibility and cost-effectiveness of RSS. Consider a conveyor system comprising two distinct types of conveyor segments: Type A, with a set length and flow speed, and Type B, featuring a turntable capable of altering directions as depicted in Figure 5(a). It is assumed that the transit time for goods on segment A is odd, whereas for conveyor B, this time is essentially negligible. Positioned along the mainline’s direction, the indices of segment B are sequentially labeled $1,2,3,$ and so forth. The orientation of each B segment coincides with the mainline when its index shares the same parity as the current time; if not, it shifts to align with the branch line. Consequently, by initiating the loading of goods onto the main line exclusively at even-numbered intervals, the goods will seamlessly proceed along the mainline. To divert them from the mainline, we could activate a barrier on segment B, blocking the goods for one unit of time when they reach it. By controlling the state of all the barriers, we can divert the goods to different branch lines. The production rhythm of such line is two units of time.
(a) Mainline-branch structure
(b) Network of conveyor segments
Figure 5: Conveyor system with two types of conveyor segment
The rhythmic manner remains applicable when applying the aforementioned control method to a four-line network in Figure 5(b), where adjacent lines always have opposite directions. Each line maintains consistent throughput when all the barriers are inactive. Activating a barrier at the appropriate time allows goods to switch to another line. However, this may conflict with the original rhythm of goods on the new line. We need a scheduling and reservation mechanism to avoid such conflicts.
4.2 Virtual platoon and cycle
If we use robots to replace the functionality of conveyors in the motivation example, then the problem becomes a MAPF issue. Lin et al., (2021) investigated a method of incorporating rhythm into traffic management at the macro-traffic level and introduced the concept of virtual platoon (VP). VPs represent the spatiotemporal slot that is generated in a rhythmic manner and is available for being occupied by vehicles. As shown in Figure 6(a), VPs move along a designated spatiotemporal trajectory and maintain a constant distance between each other, resembling an endless series of train carriages connected by invisible ropes of equal length in each road. We discretize time based on the generation time of VPs at the entrance, and we refer to each segment as a cycle. Each entrance generates VPs at a predetermined pace, noted as the cycle length, which is equal to the time gap between the arrival time of two consecutive VPs at each node in the network. By staggering the entry times of VPs at different entrances within the same cycle, VPs pass one intersection alternately with relatively short headway and the intersection capacity could be maximized. Figure 6(b) shows the spatiotemporal trajectory of VPs in a four-lane network, without any collision or deceleration.
Whenever a vehicle needs to pass through the network, the system reserves one unit of capacity from the corresponding VPs, and then releases it with an appropriate start cycle. Then that vehicle enters the network, following the movement of certain VPs within each route segments to reach its destination, and exits the network. Through the reservation mechanism, RC ensures that the number of vehicles within each VP does not exceed its predetermined capacity. Therefore, the trajectories are conflict-free without any blockage.
(a) Virtual platoons in a four-lane network
(b) Trajectory of virtual platoons in different directions. (Lin et al.,, 2021)
Figure 6: Concept of Rhythmic Control
RC is essentially a traffic control method, akin to a combination of globally coordinated traffic signal control and road network perimeter control. Within this framework, it is essential to clarify the movement patterns of each entity for integration into a sorting-specific robot management system. For detailed technical information on the RC strategy, we refer readers to Lin et al., (2021). The following points provide a foundational overview critical to our study on the implementation of the RC-S scheme:
  • A VP represents a time–space slot that is available for being occupied by vehicles and is generated in a rhythmic manner in each entrance of the network.
  • VPs move in a straight line at a constant speed while maintaining a fixed distance, ensuring that the total number of VPs present in the road network remains the same at any given moment.
  • After entering the network, vehicles need to continuously move along with a specific VP or transfer from one VP to another until exiting through an exit.
  • At any given moment, the number of vehicles accommodated by a VP cannot exceed its capacity. If the required VP capacity for vehicle travel is insufficient, the vehicle needs to queue and wait at the entrance until entering the network in a subsequent cycle.
4.3 Concept of RC-S
In order to embody RC in the RSS, we regulate the behavior of robots, and propose the scheme of rhythmic control for sorting scenario, abbreviated as RC-S. RC-S can be regarded as a centralized MAPF algorithm. To focus on the behavior of each individual robot, we discretize the map into grids, defining each grid where robots can pass through as a node. Given $(n_{v},n_{h})$, the network in sorting zone is represented by a directed graph $\mathcal{G}=(\mathcal{V}_{c},\mathcal{V}_{u},\mathcal{A})$, where $\mathcal{V}_{c}$, $\mathcal{V}_{u}$ and $\mathcal{A}$ stand for the set of conflict nodes, unloading nodes and links, respectively. To reduce traffic conflicts and deadlocks, we adopt a one-way network, where aisles from different directions are arranged in an interleaved pattern, as shown in Figure 7. Furthermore, the behaviors of robots are limited to three actions: moving straight, dropping off parcel without stopping, and left turning. The joint movement inside a virtual platoon is not practical for robots. To avoid collision within a platoon, the capacity of each VP within a cycle is set to one robot. Let $\tau_{e}$ and $\tau_{c}$ denote the fixed travel time of VPs on each unit cell in the grid map and the cycle length of rhythmic control, respectively. Under RC-S, robot enters the network through one aisle with fixed interval $\tau_{c}$, thus the upper bound of robot flow on each grid is $1/\tau_{c}$. In each cycle, the distribution of VPs could be divided into four phases (Figure 7(a)-(d)), and the duration of each phase is $\tau_{e}$. We have:
$\tau_{c}=4\tau_{e}$ (1)
Figure 7: Illustration of RC-S. (a)-(d) are the four phases in one cycle.
$\tau_{e}$ and $\tau_{c}$ are determined by the maximum velocity of robot $v_{max}$ and the maximum loading rate of each loading station $r_{ls}$, respectively:
$\tau_{e}\geq\frac{D}{v_{max}}$ (2)
$\tau_{c}\geq\frac{1}{r_{ls}}$ (3)
where $D$ is the length of each grid. At phase(a) of a cycle, new VPs are released at the entrance of each aisle; at phase(c), VPs leave sorting zone through the exit of each aisle. Thus, if we do not count the VPs at the exit of each aisle, then the number of VPs in the network remains the same throughout the entire sorting process. Intersections are occupied by VPs at phases(a) and (c); each outlet could be served by four VPs in four different directions at phases(b) or (d). The fixed number of VPs and the regularity in their phases allow us to manage and analyze this system from a global perspective. The speed of each VP could be derived by:
$v_{VP}$ $=\frac{D}{\tau_{e}}$ (4)
Figure 8: Illustration of turning and transfer process for one robot (red dot)
In addition, auxiliary road for vehicle turning in the original RC is removed, since the robot is capable to spin at the intersection grid for turning. The process of turning and transfer between VPs are demonstrated in Figure 8, which takes six consecutive phases. As shown in the figure, the robot to turn first decelerates to zero, stops following its original VP, rotates 90 degrees at the intersection, joins a new VP, and accelerate to the speed of VP. Turning action entails four phases for both the original and subsequent VPs, totaling eight phases that are unavailable to other robots. However, the actual motion spans merely six consecutive phases, thus leading to a reduction in traffic capacity.
During this process, the parameters of RC-S needs to satisfy the following conditions:
$d(\tau_{e},v_{VP},v_{max},c_{max})\geq 2D$ (5)
$\tau_{e}\geq\frac{\pi}{2\omega_{r}}$ (6)
Constraint (5) implies that the robot can catch up with the VP in a turn, where $d(\tau_{e},v_{VP},v_{max},c_{max})$ denotes the maximum travel distance a robot can cover in acceleration/deceleration process. The expression is derived in detail in Appendix B; Constraint (6) imposes limits on the rotation speed. In summary, incorporating expressions (1)-(6) and the kinematic parameters of the robots, we could determine the parameters for RC-S.
The trajectory of robot in one delivery task should start from one loading station, pass the target outlet corresponding to the destination of parcel, and end at another loading station. We define feasible path in RC-S:
DEFINITION 1.
A feasible path in RC-S is a sequence of VPs and their occupying cycles. The trajectories traveled by these VPs during the occupied cycles are connected end-to-end, linking two loading stations at the periphery of the network.
When a robot performs a sortation task, it moves from one loading station $i$ to another loading station $j$, $i,j\in\mathcal{S}$, where $\mathcal{S}$ is the set of all loading stations. Let $\mathcal{O}$ represent the set of outlets. Then, we denote the set of all feasible path as $\mathcal{R}=\cup_{i\in\mathcal{S},j\in\mathcal{S},k\in\mathcal{O}}\mathcal{R}(i,k,j)$, where $\mathcal{R}(i,k,j)$ is the subset containing the paths start from loading station $i$, pass by outlet $k$ and end at loading station $j$. To simplify the notation, we use $\mathcal{R}_{i,k}$ to represent the set of all feasible paths originating from station $i$ and passing by outlet $k$, and ^$\hat{\mathcal{R}}_{i}$ to represent the set of all feasible paths heading for station $i$, that is, $\mathcal{R}_{i,k}=\cup_{j\in\mathcal{S}}\mathcal{R}(i,k,j)$ and ^$\hat{\mathcal{R}}_{i}=\cup_{j\in\mathcal{S},k\in\mathcal{O}}\mathcal{R}(j,k,i)$. In each cycle, the feasible path assignment problem (FPA) for all the loaded robots is formulated as follows.
$\underset{x_{i,k}^{r},\hat{x}_{i,k}}{\min}\ \sum_{i\in\mathcal{S}}\sum_{k\in\mathcal{O}}(\sum_{r\in\mathcal{R}_{i,k}}c_{i,k}^{r}x_{i,k}^{r}+\hat{c}_{i,k}\hat{x}_{i,k})$
$s.t.$ $\sum_{i\in\mathcal{S}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}\delta_{i,k}^{r,\nu,e}x_{i,k}^{r}\leq N_{\nu}^{e}$ $\forall\nu\in\mathcal{V}_{c}\cup\mathcal{V}_{u},e\in\mathcal{E}$ (7)
$\hat{x}_{i,k}+\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=d_{i,k}$ $\forall i\in\mathcal{S},k\in\mathcal{O}$ (8)
$\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=\sum_{j\in\mathcal{S}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{j,k}\cap\hat{\mathcal{R}}_{i}}x_{j,k}^{r}$ $\forall i\in\mathcal{S}$ (9)
$x_{i,k}^{r},\hat{x}_{i,k}\in\{0,1\}$ $\forall i\in\mathcal{S},k\in\mathcal{O},r\in\mathcal{R}_{i,k}$ (10)
There are two sets of variables: $x_{i,k}^{r}$ equals one if the feasible path $r\in\mathcal{R}_{i,k}$ is reserved in current cycle by a robot, and zero otherwise; $\hat{x}_{i,k}$ equals one if the robot at loading station $i$ is heading for outlet $k$ and is kept waiting until the next cycle, and zero otherwise. In the objective function, $c_{i,k}^{r}$ is the travel time on path $r\in\mathcal{R}_{i,k}$, and $\hat{c}_{i,k}$ is the penalty of delay. The objective function is to minimize the total costs in this cycle. Constraint (7) indicates that in each cycle, at most one robot can be present at node $\nu$, where $N_{\nu}^{e}$ is the pre-set remaining capacity in node $\nu$ in cycle $e$, $N_{\nu}^{e}\leq 1$. Parameter $\delta_{i,k}^{r,\nu,e}$ equals one if feasible path $r$ will occupy node $\nu$ in cycle $e$ and zero otherwise. Constraint (8) describes the path allocation results for a robot at loading station $i$, where $d_{i,k}$ equals one if the robot is heading for outlet $k$, and zero otherwise. Constraint (9) is the conservation of flow constraint, ensuring a stable number of robots at each loading station.
In large-scale network, the number of feasible paths rapidly increases. It is difficult to solve FPA within a reasonable amount of time. Here we propose a heuristic algorithm to generate a feasible solution in a relatively short time, as shown in Algorithm 1. To expand the search depth, feasible paths starting at some future cycle are allowed to be reserved. We further define the entry cycle:
DEFINITION 2.
The entry cycle of a specific feasible path is the earliest cycle at which all VPs in the path are unoccupied.
So far, we have identified the essential components required to integrate heuristics into the RC-S framework: recording the status of VPs ever existed, and for each delivery task, finding a feasible path with an entry cycle, and reservingthe VPs in the space-time map. To improve robot transporting efficiency and ensure the accessibility of each outlet, we allow up to three turns in each delivery route under this scheme. Compared with classic MAPF methods, it avoids potential path conflicts from the design stage and greatly reduces the search range while maximizing intersection capacity. The optimal strategy of RC-S requires selecting a combination of feasible path and entry time for each robot, and checking the occupancy status of VPs.
Algorithm 1 A heuristic method to implement RC-S
0:  Set the maximum working cycle $N_{c}$. Use a reservation table to record the status of VPs on the map within $N_{p}$ cycles. Set the maximum search range in each cycle as $N_{t}$. Calculate feasible path set $\mathcal{R}$.
0:  Set $t=1$;
0:  At the beginning of cycle $t$, add the first robot of each queue at entrance to the candidate list.
0:  If the candidate list is empty, go to step 6; else, find the robot with the longest waiting time in the candidate list, obtain its origin $i$ and target outlet $k$. Calculate the entry cycle of each feasible path in $\mathcal{R}(i,j,k)$, and choose the feasible path with the smallest entry cycle.
0:  If the entry cycle is less than $N_{t}$, reserve the VPs according to the feasible path; else, add the waiting time of the robot by $\tau_{c}$.
0:  Remove the robot from the candidate list and go to step 3.
0:  $t=t+1$. If $t>N_{c}$, end; else, add new VPs at the entrance of each aisle, labeled as unoccupied in the reservation table. Remove the VPs which have left the network. go to step 2.
Algorithm 1 has a complexity of $O(N_{t}\cdot\lvert R\rvert)$ in each cycle. Moreover, when the algorithm is forcibly terminated midway, it can still produce a feasible solution, where robots that are not served will wait until the next cycle. The performance of Algorithm 1 will be validated in section 6.
5 Performance Analysis of RC-S
In a multi-robot system, the mutual influence among robots often leads to suboptimal utilization of roadway resources. In an ideal scenario, without any robot making turns and merely traveling from a loading station to a designated station, the flow of robots would be uniformly distributed across each link, maximizing network capacity. However, under such circumstances, it becomes unfeasible for a single loading station to serve multiple outlets—a common requirement in sortation scenarios, especially those without pre-sorting. Therefore, we depict the network capacity reduction mechanism by analyzing the characteristics of RC-S, and investigate how the number of operational loading stations and the total number of robots affect the traffic.
Two important metrics in traffic analysis (Daganzo,, 2010) are the total vehicular distance traveled per second of operation $\bar{m}(veh\cdot m/s)$, and the average travel distance $\bar{l}$. The relationship among system throughput $\tilde{T}_{O}$ and these two metrics is shown in equation (11):
$\tilde{T_{O}}=\frac{\bar{m}}{\bar{l}}$ (11)
Under the logic of RC-S, the first metric $\bar{m}$ is numerically consistent with the product of the number of occupied VPs, denoted by $n^{occupied}_{VP}$, and the speed of each VP, denoted by $v_{VP}$:
$\bar{m}=n^{occupied}_{VP}\cdot v_{VP}$ (12)
The speed of each VP could be obtained by expression (4)-(6) with the parameters of RC-S, while the number of occupied VPs is influenced by two factors: workforce factor $\kappa$, representing the impact of the number of workers on the quantity of VPs that can be occupied; attenuation factor $\beta$, representing the utilization of VPs by the feasible paths generated by RC-S. In addition, the number of available robots $n_{r}$ also limits the upper bound of occupied VPs. Above all, the expression for calculating $n^{occupied}_{VP}$ can be formulated as:
$n^{occupied}_{VP}=\min\{\kappa\cdot\beta\cdot n_{VP},n_{r}\}$ (13)
where $n_{VP}$ denotes the total number of VPs in the network, which is constant given the scale of the network. According to the phase analysis in section 4.3, the RC-S scheme stipulates that half of the aisle cells are occupied by VPs at any given time. Therefore, we obtain:
$n_{VP}=\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2}$ (14)
Now we investigate the forms of the two factors.
Workforce factor
In the operational context, the relationship between a VP and its respective entrance mirrors the relationship between a conveyor belt and its driving electric motor. A VP becomes eligible for occupation only when its associated loading station, situated at the aisle’s entrance, is operational—indicative of the corresponding worker being actively engaged in tasks. The quantity of such active VPs is a critical determinant of the maximum capacity for robot throughput. Take, for instance, a scenario where 8 workers are stationed within an RSS featuring a $6+6$ aisle network. To ensure even distribution of traffic flow and mitigate congestion at the aisle network’s periphery, these workers are strategically positioned along the four edges, as depicted in Figure 9. For ease of discussion, the notation $\alpha$ will be employed to denote the ratio of active workers to the network’s maximum operational capacity in discussions that follow.
$\alpha=\frac{n_{w}}{n_{h}+n_{v}}$ (15)
Figure 9: The impact of number of workers on the VPs that can be occupied
The top 4 horizontal aisles and left 4 vertical aisles are served by workers, and thus robots could be loaded with parcels at the corresponding entrances. We merge the areas covered by these aisles and obtain the blue polygon. The ratio of area, namely the workforce factor $\kappa$, could be obtained by:
$\kappa=1-{(1-\alpha)}^{2}$ (16)
Attenuation factor
We now examine the path reservation strategy within the RC-S framework. In a first-come-first-served basis, the system allocates to each newly requested sorting task an available robot, assigning it the currently shortest viable route. This robot then proceeds through the network along a predefined spatio-temporal path, ensuring a collision- and queue-free delivery within the aisle network. However, this spatio-temporal reservation method exhibits a degree of ”inflexibility” because it obligates the robot to adhere strictly to its assigned spatio-temporal trajectory, precluding any alterations like deceleration or temporary halts. Such inflexibility can lead to underutilization of traffic capacity, as VPs that are momentarily unoccupied cannot always be sequentially linked to construct a continuous path if they do not intersect at the same point in time, thereby diminishing the network’s traffic capacity. To quantify the impact of this reduction, we introduce an attenuation factor $\beta$, representing the proportion of network throughput decrease relative to its potential maximum.
Figure 10: The ”pile up” process of spatiotemporal occupation under RC.
Figure 11: The loss of capacity in different network.
The attenuation factor $\beta$ rises with the increase of $(n_{h}+n_{v})$. Under the RC-S protocol, the aisle’s spatiotemporal occupation showcases a ”pile-up” effect, as depicted in Figure 10. The figure’s horizontal axis marks the spatial position within an aisle, while the vertical axis tracks discretized time intervals. Each robot’s transit through this path necessitates occupying a VP, visualized as a blue bar in the illustration. The collective occupation by different robots generates the ”pile-up” phenomenon. The capacity loss at any moment is represented by the unoccupied area beneath the upper boundary of this accumulation.
When comparing networks of varying scales, i.e., those with shorter versus longer aisle lengths, we maintain a continuous approach: assuming that new sorting tasks emerge with starting points evenly spaced along the boundary and target destinations uniformly spread across the network. In a stable operation, the flow distribution of paths under consistent control logic remains analogous across different network sizes, with variations only in total path lengths. Accordingly, the pattern of unoccupied spaces beneath the upper boundary remains consistent across scales. Figure 11 illustrates this by comparing capacity losses across aisles of diverse lengths. Longer aisles tend to exhibit extended periods of spatiotemporal occupation, potentially exacerbating capacity loss — the unoccupied intervals shown in Figure 11. In consequence, we can reasonably infer that the absolute amount of lost capacity increases quadratically with the number of aisles $(n_{h}+n_{v})$, and the proportion of the lost traffic capacity should grow linearly (because the number of available VPs increases linearly with the length of aisle). In summary, the expression of attenuation factor $\beta$ is as follows:
$\beta(n_{h},n_{v})=\frac{1}{a+b\cdot(n_{h}+n_{v})}$ (17)
where $a$ and $b$ are two parameters. Parameter $a$ is slightly greater than 1 and $b$ is relatively small, indicating that for a small-scale network, the attenuation factor should be close to 1 because there are fewer conflict points in one feasible path. We use simulation to obtain the real performance of RC-S and apply linear regression to calibrate the value of the parameters in the attenuation coefficient. Experiments show that a proper value is $a=1.4$, $b=0.012$. Note that, due to the presence of the continuity assumption, there is a certain bias in $\beta$ when the number of aisles is small. Specifically, in smaller networks, this factor may overestimate the system’s throughput, especially in scenarios where the activation rate of loading stations is low. This is because some path shapes may not exist in such networks. Experimental validation of this effect will be presented in the subsequent section. To this end, we derive the final expression of number of occupied VPs:
$n^{occupied}_{VP}(\alpha,n_{r})=\min\{\frac{1-{(1-\alpha)}^{2}}{a+b(n_{h}+n_{v})}\cdot\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2},\ n_{r}\}$ (18)
LEMMA 1.
$n^{occupied}_{VP}(\alpha,n_{r})$ is non-decreasing and concave for all $\alpha$ and $n_{r}$ in its domain.
The proof of Lemma 1 can be accomplished by verifying that the partial derivatives of the two terms in the minimum function are non-negative and the Hessian matrices are both semi-negative definite. Lemma 1 indicates that opening more loading stations will increase the capacity of VPs within the network, thereby raising the upper limit of total throughput. However, the marginal returns of constructing new loading stations will gradually diminish.
Average travel distance
We now derive the average travel distance of a sorting task, as the remaining part of the throughput estimation formula. Based on the mechanism of RC-S, an appropriate selection of paths is based on the principles of evenly distributed sorting demand, workload balancing of workers, and minimizing travel distance. Workload balancing principle ensures that loading stations do not experience starvation, thus preserving the stability of the system. It also aligns better with real-world scenarios. To maintain a balanced flow of robots between loading stations, we arrange the workers at the centers of each side, covering a length equal to $\alpha$ times the side length. Figure 12 illustrates this layout, where outlets are divided into four types of areas based on their positions. Sorting demand are uniformly and continuously distributed across the entire area. As mentioned earlier, each time a robot turns, it occupies additional space in the VP fleet, so it is necessary to restrict the maximum permissible robot’s turning times in one delivery. However, from the diagram, it can be observed that to ensure the outlets in area 3 are reachable from each loading station, paths with at least three turns are required. According to our simulation experiments, introducing paths with four or more turns did not effectively improve the sorting throughput. Therefore, we prioritize paths with fewer turns and do not consider paths that require more than three turns. We denote the path length serving area $i$ as $l_{i}$, with the proportion of this area represented by $p_{i}$. We will next present the rationale for their calculation and the formulas. For a detailed derivation process, please refer to the appendix B.
Figure 12: Dividing areas according to the layout
l
Figure 13: Service paths for different categories
Firstly, for area 1, servicing the outlets requires one change of direction, as shown in Figure 13(a1). Based on the balance of workload, the turning points should have an equal probability of occurring within the central area. We then perform a weighted accumulation based on the length and then take the average. Consequently, we obtain expressions for the area proportion and the expected path length:
$p_{1}$ $=\alpha(1-\alpha)$ (19)
$E[l_{1}]$ $=2D\cdot[\frac{\frac{9+a^{2}}{6}(n_{h}^{2}+n_{v}^{2})-n_{h}n_{v}-\frac{1}{3}}{n_{h}+n_{v}}+1]$ (20)
For paths that pass by the areas of area 2, their endpoints are located at either the lower or upper loading stations, as shown in Figure 13(b1) and (b2), respectively. Let $(x_{1},y_{1})$, $(x_{2},y_{2})$ denote the coordinates of the two turning points in a path. $x_{1}$ and $x_{2}$ are uniformly distributed over the length covered by the workers, while the $y_{1}$ and $y_{2}$ are identical and uniformly distributed across the area. Notably, when x1 = x2, the path actually does not involve any turns. Let the robot choose the nearest top or bottom exit point when making the second turn. The expressions for $p_{2}$ and $E[l_{2}]$ are as follows:
$p_{2}$ $=\alpha(1-\alpha)$ (21)
$E[l_{2}]$ $=2D\cdot[\frac{\alpha(n_{h}^{2}+n_{v}^{2})}{3(n_{h}+n_{v})}-\frac{2}{3\alpha(n_{h}+n_{v})}+\frac{3n_{h}n_{v}}{2(n_{h}+n_{v})}]$ (22)
For Area 3, as previously mentioned, only paths involving three turns can service it. Area 3 includes four individual small areas. The two small areas on the bottom, as their service paths shown in Figure 13(c1), have a length equal to the paths with one turn in Figure 13(a1); the two areas on the top, with service paths depicted in Figure 13(c2), involve two extra segments of path in the y-direction (or in the x-direction when the starting point is on the left or right side). As a result, the expressions for $p_{3}$ and $E[l_{3}]$ are as follows:
$p_{3}$ $={(1-\alpha)}^{2}$ (23)
$E[l_{3}]$ $=2D\cdot(\frac{n_{h}+n_{v}}{2}+\frac{(1+\alpha)}{4}\cdot\frac{n_{h}n_{v}}{n_{h}+n_{v}})$ (24)
Finally, for Area 4, which is in the center, we assume it is equally served by the paths of area 1 and area 2.
$p_{4}$ $={\alpha}^{2}$ (25)
$E[l_{4}]$ $=\frac{1}{2}E[l_{1}]+\frac{1}{2}E[l_{2}]$ (26)
In summary, the average travel distance of a sorting task can be obtained by the following weighted summation:
$\bar{l}(n_{h},n_{v},\alpha)=\sum_{i=1}^{4}p_{i}E[l_{i}]$ (27)
Now, we further investigate the properties of equation (27).
LEMMA 2.
[…]
Proofs for Lemma 2 is provided in Appendix B. The condition within the lemma indicates that the conclusion is applicable to an RSS where the length-to-width ratio does not exceed two. This is a reasonable setup in practice, as excessively elongated sites accommodate fewer outlets under the same area. Lemma 2(i) states that the efficiency of sorting by robots exhibits a trend of first increasing and then decreasing with the increase in the number of workstations. This is because, under the worker load balancing criterion, the average distance from the corner loading stations to each outlet is greater than the stations located in the middle of the four sides. Lemma 2(ii) suggests that, under the permissible conditions, designing the aisle network in a square shape reduce the demand for robots. As aforementioned in section 3, the density of outlets remains constant, thus the distribution of outlets in RSS does not need to be altered as well. Lastly, Lemma 2(iii) provides the range of average distance to be covered in a single delivery within RC-S. This will serve as a foundation for the efficiency analysis in the following sections.
Combining equations (15)-(27), the estimation formula of sorting throughput under given configuration $(n_{h},n_{v},n_{w},n_{r})$ is as follows:
$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})=\frac{D\cdot n^{occupied}_{VP}(\frac{n_{w}}{n_{h}+n_{v}},n_{r})}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{w}}{n_{h}+n_{v}})}$ (28)
If we fix the values of $n_{h}$ and $n_{v}$, the upper bound of system throughput could be calculated by setting the value of $n_{r}$ to infinity and assigning workers to all loading stations:
$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})\leq\frac{D\cdot n^{occupied}_{VP}(\frac{n_{l}}{n_{h}+n_{v}},\infty)}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$ (29)
We use $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ to denote this upper bound in the right side of inequality (29). Here we investigate the properties of this bound.
PROPOSITION 1.
$\frac{\partial\tilde{T}_{M}(\cdot)}{\partial n_{l}}$ has one zero point within the valid range of $n_{l}$. Denote the value as $n_{l}^{(0)}$. Then $\tilde{T}_{M}(\cdot)$ is monotonically increasing in the interval $(0,n_{l}^{(0)}]$ and monotonically decreasing in the interval $(n_{l}^{(0)},n_{h}+n_{v}]$.
The proof of Proposition 1 can be derived by incorporating Lemma 1 and Lemma 2, which are depicted in detail in Appendix B. Proposition 1 reveals that under the RC-S robot control scheme, the system’s maximum operational efficiency has a critical point, and this critical point is not achieved when all loading stations are open. This may seem counter-intuitive, but remember that we have assumptions of workload balancing and uniformly distributed sorting demands. When any short-sighted dispatch strategy is employed at the loading stations, robots dispatched from the corners will negatively impact the overall throughput of the entire system.
6 Numerical Validation
6.1 Comparative analysis: RC-S versus benchmark traffic management framework
In this section, we compare our proposed RC-S against cooperative A* (CA*) (Silver,, 2021) for traffic management in a warehouse, to better illustrate the superiority of RC-S in terms of AGV service distance and system throughput. CA* is based on a simple prioritized-planning scheme: each agent is assigned a unique priority, and based on the order of priorities, algorithm find the shortest path for each agent that avoids conflicts with agents of higher priority. We use SIPP (Phillips and Likhachev,, 2011), an efficient variant of location-time A*, as the lower-level solver of CA*. Moreover,other well-known MAPF algorithm frameworks, such as PBS (Ma et al.,, 2019) and RHCR (Li et al.,, 2021), are found less efficient compared to CA* in large-scale settings ($n_{r}>100$) of our experiments, primarily due to the increased likelihood of cycle conflicts within the RSS framework, and thus are excluded from this comparative analysis.
Table 2: RC-S parameters in simulation
$D(m)$ $\tau_{e}(s)$ $\tau_{c}(s)$ $v_{p}(m/s)$
1 0.5 2 2
We compare the performance in two scenarios, 12+12 aisles and 20+20 aisles. To ensure uniformity in the experimental setup, we use directed maps in all tests to prevent swapping conflicts. We activate all possible loading station locations for the experiments, and parcel targets are evenly distributed among outlets. The RC-S parameters, detailed in Table 2, are based on sorting robot data from the Geekplus company. For a fair comparison of algorithm performance, robots move at a constant speed of 2m/s, matching the speed of VPs in RC-S. The simulations do not consider acceleration and deceleration phases. We conduct 10 rounds of replication for each test. In each experiment, we set a warm-up period of 10 minutes and then record the system performance for 50 minutes of continuous operations.
Tables 3 and 4 report the system throughput, average service time and average run-time. The comparative results indicate that the runtime of RC-S is consistently lower than that of CA*, and in larger-scale cases, it is even less than $1/10$ of the latter. Moreover, RC-S always assigns shorter service paths to robots, resulting in an average service time reduction of 10.3% compared to CA*. In most experiments, RC-S demonstrates superior throughput compared to CA*, with the exception observed in a 20+20 aisles scenario with $n_{r}=400$. This observation underscores the suitability of the RC-S algorithm for environments like RSS, characterized by a compact network and numerous outlets as obstacles.
Table 3: Results on 12+12 aisles directed map
$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time (ms)
RC-S CA* RC-S CA* RC-S CA*
40 8.81 4.72 11.70 14.17 5.26 5.99
80 14.94 9.03 12.90 14.57 6.32 16.85
120 16.99 12.40 13.81 15.24 7.52 35.37
160 17.58 14.32 14.15 15.79 8.07 55.31
200 17.81 14.94 14.32 16.03 8.35 63.36
Table 4: Results on 20+20 aisles directed map
$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time (ms)
RC-S CA* RC-S CA* RC-S CA*
50 7.41 3.94 19.34 21.75 14.38 11.61
100 14.12 7.79 19.92 21.95 14.96 29.94
200 21.04 15.07 20.99 22.54 19.32 99.61
300 22.34 21.18 21.48 23.33 21.30 226.76
400 22.91 24.37 21.97 23.92 22.08 363.45
6.2 Throughput estimation formula validation
This section validates the throughput estimation formula proposed in section 5. We consider 5 scenarios with different network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set as the number of activated VPs in the aisle network, plus 5 per loading station to ensure the stability of operation. Other experimental parameters are specified in section 6.1 in detail. We conduct 20 rounds of replication for each scenario. In each experiment, we set a warm-up period of one hour and then record the system performance for three hours of continuous operation.
Figure 14: Analytical formula compared with simulation results.
Figure 15: Robot flow distribution in simulation, $n_{h}=n_{v}=20$
Figure 14 shows the trend of the average service distance as a function of the number of active workstations when $n_{h}=n_{v}=20$. It indicates that the formula and simulation results present the similar trend of first decreasing and then increasing. This suggests that the service frequency of each robot is highest when about half of the loading stations are activated. Figure 14 shows the relative error between formula and simulation results under five scenarios. It can be seen that the error is always within ±$\pm 8\%$, indicating a good fit.
Figure 14 shows the trend of the system throughput as a function of the number of activated workstations when $n_{h}=n_{v}=20$. The trend demonstrates a clear marginal effect of the number of activated loading stations: when the system is operating at the maximum capacity, there is a certain degree of resource waste. Figure 14 shows the relative error between formula and simulation results for five cases, which remains within ±$\pm 15\%$. Figure 15 depicts the average traffic flow distribution during a stable operation of the system in simulation with the maximum number of workers. The darker the color of the aisle segments, the higher the traffic load. We notice that the corner loading stations have almost no incoming or outgoing traffic flow. This indicates that Algorithm 1, in practical execution, cannot effectively balance the distribution of robots in the loading stations in the extreme situation. Consequently, it incurs a certain efficiency loss, resulting in a decrease in the throughput, as demonstrated in Proposition 1.
7 Layout Design Optimization
The layout design problem aims at seeking for a balance between facility costs and long-term operating costs. The facility costs $C_{f}$ and operations costs $C_{o}$ are calculated as follows.
$C_{f}=P_{s}\cdot[2D\cdot(n_{h}-1)+W_{w}+W_{l}]\cdot[2D\cdot(n_{v}-1)+W_{w}+W_{l}]+P_{l}\cdot n_{l}$ (30)
$C_{o}=P_{w}\cdot[\theta\cdot n_{w}^{H}+(1-\theta)\cdot n_{w}^{L}]+P_{r}\cdot[\theta\cdot n_{r}^{H}+(1-\theta)\cdot n_{r}^{L}]$ (31)
Parameter $P_{s}$ and $P_{l}$ in equation (30) denote the discounted site rental cost per square meter and equipment cost per loading station, while $M_{w}$ and $M_{r}$ in equation (31) denote the discounted unit labor cost and unit rental cost of robot. We assume that there are two typical values of sorting demand, namely the average off-peak season demand ${T}^{L}$ and the average peak season demand ${T}^{H}$, with $\theta$ denoting the ratio of peak seasons.
The objective of the problem is as follows:
$C_{d}({T}^{L},{T}^{H},\theta,N_{o})=C_{f}+C_{o}$ (32)
where parameter $N_{o}$ denotes the minimum number of outlets that should be covered. The trade-off between these two types of costs in the objective function is the focus of the site planning stage. We formulate the layout design problem (LDP) as following integer programming model.
$\underset{n_{h},n_{v},n_{l},n_{w}^{H},n_{r}^{H},n_{w}^{L},n_{r}^{L}}{\min}\ C_{d}$
[…]
$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{H},n_{r}^{H})\geq{T}^{H}$ (33)
$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{L},n_{r}^{L})\geq{T}^{L}$ (34)
$(n_{h}-1)(n_{v}-1)\geq N_{o}$ (35)
$n_{l}\leq n_{h}+n_{v}$ (36)
$n_{w}^{H}\leq n_{l}$ (37)
$n_{w}^{L}\leq n_{l}$ (38)
$n_{h}=2k_{h}\quad\forall k_{h}\in\mathcal{Z}_{+}$ (39)
$n_{v}=2k_{v}\quad\forall k_{v}\in\mathcal{Z}_{+}$ (40)
$n_{h},n_{v},n_{w}^{H},n_{r}^{H},n_{w}^{L},n_{r}^{L}\in\mathcal{Z}_{+}^{(6)}$ (41)
where constraints (33) and (34) ensure that the solution meets predicted sorting requirements. Constraint (35) restricts the lower bound of the site size to accommodate an adequate number of outlets. Constraint (36) represents the maximum number of loading stations accommodated in the aisle network under RC-S. Constraints (37) and (38) limit the maximum number of workers to be stationed. By the design of the system in section 3, we add constraints (39)-(40) to ensure the number of aisles even. The remaining constraints include the expressions to calculate the cost of each component and the estimation formula of system efficiency.
It is difficult to solve LDP because of the non-linearity of the inequality constraints, especially for large-scale problems. In this research, we apply the method of penalty successive linear programming (PSLP) to solve the LP-relaxation of LDP, which exhibits good robustness and convergence properties for large-scale problems (Bazaraa et al.,, 2013). Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem along with the penalty function, and utilizes the concept of trust region (updated at each iteration) to control the step size. To ensure the constraints continuous and differentiable, we introduce additional constraints requiring that the number of robots is less than the number of available VPs, shown as follows.
$n_{r}^{H}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{H}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (42)
$n_{r}^{L}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{L}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (43)
At the end of the algorithm, the optimal solution of the relaxed problem is converted into a feasible solution for the original problem by rounding up. The details of PSLP algorithm are shown in appendix C.
8 Numerical Examples
In this section, we conduct a sensitivity analysis on the site rental cost per square meter $P_{s}$, and labor cost per man-month $P_{w}$, focusing on the trend of total costs under different scenarios as the sorting throughput level changes. Due to increased competition in the robot market, prices remain relatively stable across different companies. Consequently, we do not investigate variations in robot rental cost during the experiments, nor do we examine equipment cost at loading stations. We consider a 5-year investment plan with a monthly interest rate $\gamma_{0}=0.5\%$, and then the discounted costs of different components in LDP are calculated by:
$P_{i}=\sum_{t=0}^{60}\frac{M_{i}}{{(1+\gamma_{0})}^{t}}\ ,\quad i=s,l,w,r$ (44)
where $M_{i}$ denotes the unit cost of component $i$ in a month. We estimate the monthly warehousing rental cost using JD Logistics’ financial reports for the first three quarters of 2023. Labor cost is sourced from data related to warehousing job recruitment in Beijing, obtained from website for job posts (58.com). The data regarding the equipment cost of loading stations and the rental cost of robots are obtained from Geekplus company. The ratio of peak month $\theta$ is set to 1/6, which means that every quarter has a two-week-long promotion campaign in average. Meanwhile, we set ${T}^{L}=0.8\cdot{T}^{H}$ to ensure that there is no enormous space waste during the off-peak period. The default parameter of experiments is shown in table 5 and the configurations of RC-S keeps the same with table 2. Our algorithms are coded in Python 3.11 with the solver Gurobi Optimizer 9.0, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM. All processes are run single-threaded.
Table 5: Default parameter in sensitivity analysis
$M_{s}(CNY/m^{2}\cdot mo)$ $M_{l}(CNY/mo)$ $M_{w}(CNY/mo)$ $M_{r}(CNY/mo)$ $W_{w}(m)$ $W_{l}(m)$
10 400 5,000 200 5 5
8.1 Site rental cost
We begin by investigating the impact of site rental cost under two different scenarios: one with $N_{o}=100$ (scenario 1) and another with $N_{o}=400$ (scenario 2). The monthly rental cost of site floor space per square meter varies from $10\ CHY$ to $30\ CHY$. The results are presented in Figure 16(a) and Figure 16(b). The findings reveal that:
(a) $N_{o}=100$ (scenario 1)
(b) $N_{o}=400$ (scenario 2)
Figure 16: Comparison of total costs and facility costs proportion on different site rental cost
Table 6: Results of scenario $M_{s}=10,N_{o}=100$
${T}^{H}$ $(n_{h},n_{v})$ $(n_{w}^{L},n_{r}^{L})$ $(n_{w}^{H},n_{r}^{H})$ $C_{f}$ $C_{o}$ $C_{d}$ Proportion of
(×$(\times{10}^{6})$ (×$(\times{10}^{6})$ (×$(\times{10}^{6})$ site rental cost (%)
3000 (10,12) (2,10) (3,13) 0.53 0.67 1.20 38.80
6000 (10,12) (4,21) (5,26) 0.57 1.31 1.88 24.78
9000 (10,12) (7,33) (9,42) 0.65 2.26 2.92 15.96
12000 (12,14) (9,51) (13,67) 0.87 3.07 3.94 15.20
15000 (14,16) (12,75) (18,98) 1.12 4.20 5.32 14.07
18000 (18,18) (15,109) (23,142) 1.48 5.44 6.92 14.54
21000 (20,22) (18,147) (28,194) 1.88 6.72 8.60 15.09
24000 (24,26) (21,199) (32,261) 2.41 8.11 10.52 16.60
27000 (26,26) (25,236) (44,315) 2.79 9.91 12.70 14.74
30000 (30,30) (29,302) (49,401) 3.42 11.72 15.14 15.88
  • When the target throughput level is low, the total costs in scenario 1 is significantly lower than that in scenario 2. This is because, to accommodate more outlets, the latter requires renting a larger initial space. In both scenarios, the curve of the optimal system cost exhibits a turning point $T^{*}$. When $T^{H}>T^{*}$, the cost escalation rate increase significantly. The turning point indicates that the initial space in the current scenario is no longer sufficient to accommodate a larger throughput, necessitating an expansion of the aisle network scale. In scenario 1, this turning point occurs at $T^{*}=9,000$; whereas in scenario 2, the turning point is at $T^{*}=21,000$.
  • The proportion of facility costs increases as $M_{s}$ rises, particularly remarkable at lower sorting throughput levels. As the throughput level becomes higher, this proportion gradually decreases and stabilizes after the turning point $T^{*}$ in each case. By observing the system configurations in each dataset, we find that this phenomenon arises because, at higher throughput demands, efficiency improvements necessitate expanding the space, which leads to the increase in the site rental cost. Table 6 shows the optimal system configuration and corresponding costs in the case of $M_{s}=10$ and $N_{o}=100$. When $T^{H}>T^{*}$, the optimal solutions show not only assigning more workers and robots to loading stations but also simultaneously expanding the size of the network to alleviate traffic pressure.
  • Comparing the data for $T^{H}=15,000$ and $T^{H}=30,000$ in Table 6, the total costs for the latter are nearly three times those for the former. This indicates that due to the simultaneous growth of $C_{s}$ and $C_{o}$, the total costs growth rate of the RSS system exceeds the throughput growth rate.
From the results of study in site rental cost, the derived managerial insights are summarized as follows.
INSIGHT 1.
The advantage of RSS is evident in its low initial investment in scenarios with both throughput demand and sortation category low.
INSIGHT 2.
High density lead to decreased efficiency of workers. Instead of running RSS at full capacity, it is more cost-effective to appropriately expand the site size and reduce the proportion of activated loading stations.
8.2 Labor cost
In a similar manner as our previous experiments, we now explore the influence of different labor costs on the total costs in two scenarios with $N_{o}=100$ and $N_{o}=400$. The monthly labor cost of each worker varies from $5,000\ CHY$ to $12,000\ CHY$. The proportions of labor costs and robot rental costs are both visualized in Figure 17(a) and Figure 17(b). Our observations are concluded as follows:
(a) $N_{o}=100$ (scenario 1)
(b) $N_{o}=400$ (scenario 2)
Figure 17: Comparison of total costs and labor costs proportion on different unit labor cost
  • As $T^{H}$ increases, the proportion of labor costs gradually rises and eventually stabilizes. In every scenario, the proportion of labor costs in the optimal solution for $T^{H}=30,000$ exceeds 50%.
  • The rental cost of robots gradually increases with the level of throughput in each scenario. This is because, to maintain traffic efficiency, the number of robots needs to increase with the expansion of the network. Meanwhile, as discussed in the performance analysis in section 5, the rate of traffic capacity loss also increases, resulting in a considerable robot rental cost.
  • Comparing the two scenarios, it is evident that in larger-scale sorting sites (scenario 2), the initial increase in the proportion of robot rental cost is faster, making it challenging for RSS systems to balance the advantages of high scalability and low cost.
From the result of study in labor cost, we summarize the following managerial insight:
INSIGHT 3.
At lower levels of throughput, facility costs are the primary expense of RSS, whereas at higher levels of throughput, labor costs becomes the primary expense.
Additionally, the cost structure undergoes notably changes with varying sorting demands. In scenarios with high sorting demand, the proportion and growth rate of robot rental cost become more significant. This reveals a distinctive feature of RSS systems: Robots, as carriers of parcels, become less efficient in larger sites because they have to cover increased distances. Merely enhancing management methods is insufficient to tackle the challenge of diminishing traffic efficiency. Today’s commonly used conveyor sorters, equipped with high-power motors, necessitate substantial initial investments but experience only slight increments in total costs as sorting demands grow (Zou et al.,, 2021). Furthermore, the optimal system configurations of a conveyor sorter are not easily affected by fluctuations in the unit price of cost components (Russell and Meller, 2003b, ), which makes it suitable for long-term operation in large-scale sorting scenarios. Therefore, we conclude the following insight regarding the application of RSS:
INSIGHT 4.
Due to the traffic issues of robots, RSS is suitable for application in small-scale scenarios, such as distribution centers at the end of the supply chain.
9 Conclusion
This study conducts a comprehensive analysis of the RSS in modern warehouses, encompassing the underlying robot traffic management, efficiency analysis, as well as the cost composition and layout design of the system. By incorporating the throughput estimation, this research sheds light on the prediction of performance and resource allocation in modern warehouses or factories that utilize numerous robots, promoting the ongoing development of automation in logistics industry.
In the operations stage, efficient robot scheduling emerges as a crucial element for unlocking the business value of RSS. We propose an innovate RC-S scheme, which serves as an framework for managing a considerable number of robots simultaneously and efficiently. We provide a detailed description of the composition of phases and the method for setting parameters in RC-S. Additionally, we present a mathematical programming model FPA to minimize the travel costs at each cycle, and a heuristic algorithm as online solver. Subsequently, a theoretical analysis of the efficiency of RC-S is conducted, exploring the impact of various system configurations on throughput. In the validation section, we first compare our proposed traffic management framework with other common framework based on A*. The simulation results indicate that our control method achieves a high level of performance and computational efficiency. Furthermore, we validate the throughput estimation formula. Results show that the analytical model fits the real traffic of robots well.
In the site planning stage, warehouse managers need to make decisions that involve a trade-off between the initial site investment $C_{s}$, and the discounted expected operations costs $C_{o}$. We analyze the cost composition of the RSS, and propose a layout design optimization model LDP that minimizes the total system costs. Through a detailed sensitivity analysis examining the cost dynamics of site rental and labor, our findings demonstrate a distinct cost distribution based on throughput levels. Specifically, at lower throughput, the rental cost for the site represents a considerable proportion of the total costs. However, at higher throughput, the expenditure on labor emerges as the predominant cost element. Results also validate the discussions on the properties of our model. The key insights can guide managers in understanding the investment and returns of applying the RSS, thereby reducing the total costs of the warehouse.
For future work, there are two directions to explore. Firstly, optimizing layout design considering non-uniform demand and various outlet distributions would further improve the model’s effectiveness. Secondly, extending the analysis to cover the entire process from order fulfillment to sorting in modern warehouses would provide valuable insights for enhancing system performance.
Appendix A: Main notations
Table A-1: Summary of main abbreviations
Abbreviations Explanation
RSS robotic sorting system
RMFS robotic mobile fulfill system
MAPF multi-agent path finding problem
SOQN semi-open queueing network
OQN open queueing network
RC-S rhythmic control for sorting scenario
VP virtual platoon
FPA feasible path assignment problem
LDP layout design problem
PSLP penalty successive linear programming
Table A-2: Notations and explanation
Notations Explanation
Sets
$\mathcal{G}$
graph constituted of all nodes and links
$\mathcal{V}_{c}$
set of all conflict nodes
$\mathcal{V}_{u}$
set of all unloading nodes
$\mathcal{A}$
set of all links
$\mathcal{S}$
set of all loading stations
$\mathcal{O}$
set of all outlets
$\mathcal{E}$
set of all cycles
$\mathcal{R}$
set of all feasible path
$\mathcal{R}(i,j,k)$
set of feasible path connecting loading station $i$ and $j$, passing outlet $k$
$\mathcal{R}_{i},k$
set of feasible path starting at loading station $i$ and passing by outlet $k$
^$\hat{\mathcal{R}}_{i}$
set of feasible path ending at loading station $i$
$\mathcal{I}_{x}$
set of locations of loading stations on the top and bottom sides
$\mathcal{I}_{y}$
set of locations of loading stations on the left and right sides
Parameters
$D$
length of each grid in the sorting zone
$W_{w}$
width of waiting zone
$W_{l}$
width of loading zone
$C_{f}$
the facility costs of RSS
$C_{o}$
the operations cost of RSS
$C_{d}$
the total costs of RSS
$P_{f}$
cost of facility per square floor space
$P_{l}$
cost of equipment in one loading station
$P_{w}$
cost of labor per person per month
$P_{r}$
cost of robot per vehicle per month
${T}_{H},{T}_{L}$
target throughput level in peak/off-peak season
$\theta$
Ratio of peak seasons
$N_{o}$
minimum number of outlets required to meet the sorting category demands
$\gamma$
weight of operations costs in optimization model
$\tau_{c}$
time interval of VPs in RC-S
$\tau_{e}$
travel time of VPs on each link in RC-S
$v_{VP}$
speed of VPs
$v_{max}$
maximum speed of robots
$c_{max}$
maximum acceleration/deceleration of robots
$r_{ls}$
maximum loading rate of a loading station
$c_{i,k}^{r}$
travel time of feasible path $r\in R_{i,k}$
$\hat{c}_{i,k}$
penalty of delay of a sorting task from loading station $i$ to outlet $k$
$\delta_{i,k}^{r,\nu,e}$
incidence between feasible path $r\in R_{i,k}$, node $\nu$ and cycle $e$
$N_{\nu}^{e}$
remaining capacity in node $\nu$ in cycle $e$
$d_{i,k}$
sorting demand from loading station $i$ to outlet $k$
Variables
$x_{i,k}^{r}$
decision on whether feasible path $r\in R_{i,k}$ is reserved in current cycle
$\hat{x}_{i,k}$
decision on whether the robot located at loading station $i$ should wait until the next cycle, with the target outlet $k$
$n_{h},n_{v}$
number of horizontal and vertical aisles in sorting zone
$n_{l}$
number of loading stations in loading zone
$n_{w}^{H},n_{w}^{L}$
number of workers in peak/off-peak season
$n_{r}^{H},n_{r}^{L}$
number of robots in peak/off-peak season
Appendix B: Proofs
B.1 Derivation of Maximum Travel Distance in Constraint 5
We first examine the acceleration process of the robot after a turn. It needs to cover a distance of $2D$ within $2\tau_{e}$, reaching a final velocity of $v_{VP}$. We discuss two cases regarding the maximum travel distance the robot can cover while satisfying the final velocity $v_{VP}$, which are shown in Figure B-B-1 and B-B-1, respectively.
Figure B-1: Speed curves in the acceleration process
CASE 1: $2v_{max}-v_{VP}\geq 2\tau_{e}c_{max}$
The speed curve $v_{1}(t)$ and the associated travel distance $d_{1}$ are formulated as follows:
$v_{1}(t)=\left\{\begin{aligned} &c_{max}t,\quad t\leq\tau_{e}+\frac{v_{VP}}{2c_{max}}\\ &v_{VP}+c_{max}(2\tau_{e}-t),\quad\tau_{e}+\frac{v_{VP}}{2c_{max}}<t\leq 2\tau_{e}\\ \end{aligned}\right.$
$d_{1}=\int_{0}^{2\tau_{e}}v_{1}(t)=\frac{v_{VP}^{2}}{4c_{max}}+\tau_{e}v_{VP}+c_{max}\tau_{e}^{2}$
CASE 2: $2v_{max}-v_{VP}<2\tau_{e}c_{max}$
The speed curve $v_{1}(t)$ and the associated travel distance $d_{1}$ are formulated as follows:
$v_{2}(t)=\left\{\begin{aligned} &c_{max}t,\quad t\leq\frac{v_{max}}{c_{max}}\\ &v_{max},\quad\frac{v_{max}}{c_{max}}<t\leq 2\tau_{e}-\frac{v_{max}-v_{VP}}{c_{max}}\\ &v_{VP}+c_{max}(2\tau_{e}-t),\quad 2\tau_{e}-\frac{v_{max}-v_{VP}}{c_{max}}<t\leq 2\tau_{e}\\ \end{aligned}\right.$
$d_{2}=\int_{0}^{2\tau_{e}}v_{2}(t)=2\tau_{e}v_{max}-\frac{2v_{max}^{2}-2v_{max}v_{VP}+v_{VP}^{2}}{2c_{max}}$
Above all, the maximum travel distance is expressed by:
$d(\tau_{e},v_{VP},v_{max},c_{max})=\left\{\begin{aligned} &d_{1},\quad 2v_{max}-v_{VP}\geq 2\tau_{e}c_{max}\\ &d_{2},\quad 2v_{max}-v_{VP}<2\tau_{e}c_{max}\\ \end{aligned}\right.$
Similarly, the analysis for the acceleration process before a turn is consistent with that of the deceleration process, yielding the same result.
$\square$
B.2 Derivation of Average Travel Distance in Section 5
Let $I_{x}$ represents the set of x-coordinates of loading stations on the top and bottom sides, while $I_{y}$ represents the set of y-coordinates of loading stations on the left and right sides. Similarly, $W_{x}$ and $W_{y}$ denote the set of all coordinates of potential locations on the top and bottom sides, and on the left and right sides, respectively.
Area 1: Select turning points with equal probability in the central area and perform a weighted sum of all path lengths, with weights proportional to the route lengths.
$E[l_{1}]$ $=$ $\frac{2D}{\lvert I_{x}\rvert\cdot\lvert I_{y}\rvert}\cdot\sum_{x\in I_{x}}\sum_{y\in I_{y}}\{\frac{{(x+y)}^{2}}{2(n_{h}+n_{v})}+\frac{{[x+(n_{h}-y)]}^{2}}{2(n_{h}+n_{v})}+\frac{{[(n_{v}-x)+y]}^{2}}{2(n_{h}+n_{v})}$
$+\frac{{[(n_{v}-x)+(n_{h}-y)]}^{2}}{2(n_{h}+n_{v})}\}$
$=$ $2D\cdot[\frac{\frac{9+a^{2}}{6}(n_{h}^{2}+n_{v}^{2})-n_{h}n_{v}-\frac{1}{3}}{n_{h}+n_{v}}+1]$
Area 2: Calculate the average path lengths for starting points on the top and bottom sides and on the left and right sides, separately. According to the number of loading stations, perform a weighted sum. Based on the distance between the first turning point and the starting point, paths are divided into those that return to the original side and those that reach the opposite side, with each scenario having a probability of 1/2. For the former, the vertical movement distance is $n_{v}/2$ or $n_{h}/2$, while for the latter, it is $n_{v}$ or $n_{h}$, depending on the start point.
$E[l_{2}]$ $=$ $2D\cdot\big{\{}\frac{n_{v}}{n_{h}+n_{v}}\sum_{x\in I_{x}}\sum_{z\in I_{x}}\frac{\frac{1}{2}(\frac{n_{h}}{2}+\lvert z-x\rvert)+\frac{1}{2}(n_{h}+\lvert z-x\rvert)}{{\lvert I_{x}\rvert}^{2}}$
$+\frac{n_{h}}{n_{h}+n_{v}}\sum_{y\in I_{y}}\sum_{z\in I_{y}}\frac{\frac{1}{2}(\frac{n_{v}}{2}+\lvert z-y\rvert)+\frac{1}{2}(n_{v}+\lvert z-y\rvert)}{{\lvert I_{y}\rvert}^{2}}\big{\}}$
$=$ $2D\cdot[\frac{\alpha(n_{h}^{2}+n_{v}^{2})}{3(n_{h}+n_{v})}-\frac{2}{3\alpha(n_{h}+n_{v})}+\frac{3n_{h}n_{v}}{2(n_{h}+n_{v})}]$
Area 3: We first calculate the basic average path length. In accordance with workload balance, the starting and ending points of the path should be equally likely to fall on each workstation. Then add the detour lengths for the two different scenarios multiplied by their respective probabilities, according to the path allocation rules.
$E[l_{3}]$ $=$ $2D\cdot\sum_{x\in I_{x}}\sum_{y\in I_{y}}[\frac{x+y}{\lvert I_{x}\rvert\cdot\lvert I_{y}\rvert}+\frac{n_{h}}{2(n_{h}+n_{v})}\cdot\frac{\sum_{x\in W_{x}\backslash I_{x}}2(x-\frac{\alpha n_{v}}{2})}{\lvert W_{x}\backslash I_{x}\rvert}$
$+\frac{n_{v}}{2(n_{h}+n_{v})}\cdot\frac{\sum_{y\in W_{y}\backslash I_{y}}2(y-\frac{\alpha n_{h}}{2})}{\lvert W_{y}\backslash I_{y}\rvert}]$
$=$ $2D\cdot(\frac{n_{h}+n_{v}}{2}+\frac{(1+\alpha)}{4}\cdot\frac{n_{h}n_{v}}{n_{h}+n_{v}})$
B.3 Proof of Lemma 2
We first prove (i). By equation (27), the expression of average travel distance of a sorting task is as follows:
$\bar{l}(n_{h},n_{v},\alpha)=$ $2D\cdot[\frac{n_{h}^{2}+n_{v}^{2}}{n_{h}+n_{v}}\cdot\frac{-\alpha^{4}-5\alpha^{2}+18\alpha}{12}+\frac{n_{h}n_{v}}{n_{h}+n_{v}}\cdot\frac{-\alpha^{2}+2\alpha}{4}$
$+\frac{1}{n_{h}+n_{v}}\cdot\frac{3\alpha^{3}-\alpha^{2}-3\alpha-5}{12}+(n_{h}+n_{v})\cdot\frac{\alpha^{2}-2\alpha+1}{2}+\frac{-\alpha^{2}+2\alpha}{2}]$
where $\alpha$ is the ratio of the number of workers to the number of aisle entrances, $\alpha\in(0,1]$. We consider the partial derivative:
$\frac{\partial\bar{l}(n_{h},n_{v},\alpha)}{\partial\alpha}=$ $2D\cdot[\frac{n_{h}^{2}+n_{v}^{2}}{n_{h}+n_{v}}\cdot\frac{-4\alpha^{3}-10\alpha+18}{12}+\frac{n_{h}n_{v}}{n_{h}+n_{v}}\cdot\frac{-\alpha+1}{2}$
$+\frac{1}{n_{h}+n_{v}}\cdot\frac{9\alpha^{2}-2\alpha-3}{12}+(n_{h}+n_{v})(\alpha-1)-\alpha+1]$
At the two endpoints of the range of $\alpha$, the derivative has the value:
$\lim_{\alpha\xrightarrow{}0}\frac{\partial\bar{l}(n_{h},n_{v},\alpha)}{\partial\alpha}$ $=2D\cdot[\frac{3(n_{h}^{2}+n_{v}^{2})}{2(n_{h}+n_{v})}+\frac{n_{h}n_{v}}{2(n_{h}+n_{v})}-\frac{1}{4(n_{h}+n_{v})}-(n_{h}+n_{v})+1]$
$\leq 2D\cdot[\frac{n_{h}^{2}+n_{v}^{2}-3n_{h}n_{v}}{2(n_{h}+n_{v})}+1]<0$
$\leq 2D\cdot[-\frac{1}{4}min\{n_{h},n_{v}\}+1]\leq 0$
$\lim_{\alpha\xrightarrow{}1}\frac{\partial\bar{l}(n_{h},n_{v},\alpha)}{\partial\alpha}$ $=2D\cdot[\frac{n_{h}^{2}+n_{v}^{2}}{3(n_{h}+n_{v})}+\frac{1}{3(n_{h}+n_{v})}-1]$
$\geq 2D\cdot[\frac{1}{3}min\{n_{h},n_{v}\}-1]>0$
Similarly, we could calculate the second-order derivative and obtain that $\frac{\partial^{2}\bar{l}(n_{h},n_{v},\alpha)}{\partial\alpha^{2}}\geq 0$. Above all, we can prove that $\bar{l}(n_{h},n_{v},\alpha)$ decreases initially and then increases.
Next, we prove the advantage of square-shape site in (ii), namely the network with $n_{h}=n_{v}$. Without loss of generality, we assume $n_{h}=k\cdot n_{v}$, $k\geq 1$. Our objective is to show that the average travel distance is minimum when $k=1$. Let the area of the sorting zone be $S\cdot 4D^{2}$, then $n_{h}=\sqrt{kS}$, $n_{v}=\sqrt{S/k}$. Consider the derivative:
$\frac{\partial\bar{l}(S,k,\alpha)}{\partial k}=$ $2D\cdot[\frac{k^{2}-1}{2k^{\frac{3}{2}}}\cdot\sqrt{S}\cdot\frac{-\alpha^{4}-5\alpha^{2}+18\alpha}{6}+\frac{k-1}{2k^{\frac{1}{2}}{(k+1)}^{2}}\cdot\sqrt{S}\cdot\frac{-\alpha^{2}+2\alpha}{4}$
$+\frac{k-1}{2k^{\frac{1}{2}}{(k+1)}^{2}}\cdot\frac{1}{\sqrt{S}}\cdot\frac{3\alpha^{3}-\alpha^{2}-3\alpha-5}{12}+\frac{k-1}{2k^{\frac{3}{2}}}\cdot\sqrt{S}\cdot\frac{\alpha^{2}-2\alpha+1}{2}]$
It is easy to prove that the right side is consistently non-negative. As a result, the derivative is non-negative when $k\geq 1$, thus $\bar{l}(S,k,\alpha)\geq\bar{l}(S,1,\alpha)$.
To prove (iii), we first obtain the upper bound of $\bar{l}(n_{h},n_{v},\alpha)$. By (i), the upper bound can only be attained at the two endpoints:
$\bar{l}(n_{h},n_{v},0)=$ $n_{h}$
$\bar{l}(n_{h},n_{v},1)=$ $\frac{9}{8}n_{h}-\frac{1}{4n_{h}}+\frac{1}{2}<\frac{9}{8}n_{h}$
While $\alpha=0$ is not feasible for operation, we conclude $\bar{l}(n_{h},n_{v},\alpha)<\frac{9}{8}n_{h}$. According to the sorting demands with an average distribution, we can easily derive that the average distance for all three paths are greater than $n_{h}$, then we have: $\bar{l}(n_{h},n_{v},\alpha)<n_{h}$. The proof is completed.
$\square$
B.4 Proof of Proposition 1
We first obtain the expression of $\tilde{T}_{M}(n_{h},n_{v},n_{l})$:
$\tilde{T}_{M}(n_{h},n_{v},n_{l})=\frac{D}{\tau_{e}}\cdot\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2[a+b(n_{h}+n_{v})]}\cdot\frac{\frac{2n_{l}}{n_{h}+n_{v}}-{(\frac{n_{l}}{n_{h}+n_{v}})}^{2}}{\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$
We further denote:
$G_{n_{h},n_{v}}(\alpha)=\frac{2\alpha-\alpha^{2}}{\bar{l}(n_{h},n_{v},\alpha)}$
$\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}=\frac{2(1-\alpha)\cdot\bar{l}(\cdot)-\frac{\partial\bar{l}(\cdot)}{\partial\alpha}\cdot(2\alpha-\alpha^{2})}{\bar{l}^{2}(\cdot)}$
From Lemma 1 and Lemma 2, we have:
$\lim_{\alpha\xrightarrow{}0}\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}=\frac{2}{\lim_{\alpha\xrightarrow{}0}\bar{l}(\cdot)}\geq\frac{2}{\sqrt{n_{h}n_{v}}}>0$
$\lim_{\alpha\xrightarrow{}1}\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}=\frac{-\frac{\partial\bar{l}(\cdot)}{\partial\alpha}}{\lim_{\alpha\xrightarrow{}0}\bar{l}^{2}(\cdot)}<0$
The non-negativity of the second-order derivative of $G_{n_{v},n_{h}}(\alpha)$ can be easily proved by obtaining the expression and checking the bounds of each term in the numerator. Hence, $G_{n_{v},n_{h}}(\alpha)$ has one zero point. The expression of the derivative of $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ can be written as:
$\frac{\partial\tilde{T}_{M}(n_{h},n_{v},n_{l})}{\partial}=\frac{D}{\tau_{e}}\cdot\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2[a+b(n_{h}+n_{v})]}\cdot\frac{1}{n_{h}+n_{v}}\cdot\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}$
It is immediate to prove that $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ has one zero point within the range $n_{l}\in(0,n_{h}+n_{v}]$ and it initially increases then decreases.
$\square$
Appendix C: Penalty Successive Linear Programming
The classic penalty successive linear programming (PSLP) enjoys good robustness and convergence properties for large-scale problems (Bazaraa et al.,, 2013). Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem along with the penalty function, and utilizes the concept of trust region (updated at each iteration) to control the step size. In each iteration $k$, a direction-finding linear program is formulated based on first-order Taylor series approximations to the objective and constraint functions, in addition to appropriate trust region restrictions on the direction components. The subproblem in iteration $k$ are formed as follows:
[…]
$s.t.\quad$ $y_{i}\geq g_{i}(\omega_{k})+\nabla{g_{i}(\omega_{k})}^{T}d,\quad i=1,2,\cdots,6$ (C-1)
$-\Delta_{k}\leq d\leq\Delta_{k}$ (C-2)
$y_{i}\geq 0,\quad i=1,2,\cdots,6$ (C-3)
where $\omega=\{n_{h},n_{v},n_{w}^{H},n_{r}^{H},n_{w}^{L},n_{r}^{L}\}$ is the set of all decision variables. $\Delta_{k}$ is the bound of $d$ in iteration $k$. If we impose the constraint that the number of robots is less than the available VPs (considering the parking demand of robots during downtime, this setting is reasonable), constraints in layout design optimization model LDP are replaced with six inequalities in the LP-S:
$g_{1}(\omega)=n_{r}^{H}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{H}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (C-4)
$g_{2}(\omega)=n_{r}^{L}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{L}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (C-5)
$g_{3}(\omega)=-\frac{D\cdot n_{r}^{H}}{\tau_{e}\cdot E[l(n_{h},n_{v},\frac{n_{w}^{H}}{n_{h}+n_{v}})]}+{T}^{H}\leq 0$ (C-6)
$g_{4}(\omega)=-\frac{D\cdot n_{r}^{L}}{\tau_{e}\cdot E[l(n_{h},n_{v},\frac{n_{w}^{L}}{n_{h}+n_{v}})]}+{T}^{L}\leq 0$ (C-7)
$g_{5}(\omega)=-(n_{h}-1)(n_{v}-1)+N_{o}\leq 0$ (C-8)
$g_{6}(\omega)=n_{w}^{H}-n_{h}-n_{v}\leq 0$ (C-9)
To avoid waste of resources, the number of loading station is always equal to the peak number of workers, thus $n_{l}$ is replaced by $n_{w}^{H}$ in the constraints. The details of solving PLSP is shown in Algorithm C-1. It could be roughly divided into 2 steps in each iteration: (1) Generate LP-S and obtain the step size $d_{k}$, determine whether to stop; (2) Adjust step bounds for the next iteration.
Algorithm C-1 The PLSP algorithm to solve model LDP
0:    An initial feasible solution $\omega_{0}$onfidence intervals $0\leq\Delta_{LB}\leq\Delta_{1}$arameters $0\leq\rho_{0}\leq\rho_{1}\leq\rho_{2}\leq 1$, $\phi\leq 1$Large enough constant $\mu$Maximum iterations $N_{m}$
0:  The optimal value of layout decision variables, $\omega$
1:  Initialization: $k=1$, $R_{k}=0$, $iteration=1$
2:  while $iteration<N_{m}$ do
3:     repeat
[…]
5:        if $C_{EL_{k}}(\omega_{k})-C_{EL_{k}}(\omega_{k}+d_{k})=0$ or $\lceil\omega_{k}+d_{k}\rceil=\omega_{k}$ then
6:           Return $\omega_{k}$
7:           Stop
8:        else
9:           Calculate $R_{k}(\omega_{k},d_{k})$
10:        end if
11:        if $R_{k}<\rho_{0}$ then
12:           $\Delta_{k}=\alpha\Delta_{k}$
13:        end if
14:     until $R_{k}\geq\rho_{0}$
15:     $\omega_{k+1}=\lceil\omega_{k}+d_{k}\rceil$
16:     if $\rho_{0}\leq R_{k}<\rho_{1}$ then
17:        $\Delta_{k+1}=\phi\Delta_{k}$
18:     else if $\rho_{1}\leq R_{k}<\rho_{2}$ then
19:        $\Delta_{k+1}=\Delta_{k}$
20:     else if $R_{k}\geq\rho_{2}$ then
21:        $\Delta_{k+1}=\Delta_{k}/\phi$
22:     end if
23:     $\Delta_{k+1}=max\{\Delta_{k+1},\Delta_{LB}\}$
24:     $k=k+1$
25:  end while
In iteration $k$, the termination of algorithm could be determined by calculating the intermediate variables as follows:
$C_{E}(\omega_{k})=C_{d}(\omega_{k})+\mu\cdot\sum_{i=1}^{6}max\{0,g_{i}(\omega_{k})\}$ (C-10)
$C_{EL_{k}}(\omega_{k})=C_{d}(\omega_{k})+\nabla{C_{d}(\omega_{k})}^{T}\cdot d_{k}+\mu\cdot\sum_{i=1}^{6}max\{0,g_{i}(\omega_{k})+{\nabla g_{i}(\omega_{k})}^{T}\cdot d_{k}\}$ (C-11)
$R_{k}=\frac{C_{E}(\omega_{k})-C_{E}(\omega_{k}+d_{k})}{C_{EL_{k}}(\omega_{k})-C_{EL_{k}}(\omega_{k}+d_{k})}$ (C-12)
where $\mu$ is a large enough constant. Given that the variables in the original problem are integers, we introduced a rounding-up step for $\omega_{k}$ and the bounds in the search process. Additionally, due to the relatively limited solution space, the number of iterations in experiments remained below 50, and the total solving time was in the order of seconds. Therefore, specific details are not explicitly described in the main text.

### v2

@@
 In the contemporary logistics industry, automation plays a pivotal role in enhancing production efficiency and expanding industrial scale.
-Autonomous mobile robots, in particular, have become integral to the modernization efforts in warehouses.
-One noteworthy application in robotic warehousing is the robotic sorting system (RSS), distinguished by its characteristics such as cost-effectiveness, simplicity, scalability, and adaptable throughput control.
-While previous research has focused on analyzing the efficiency of RSS, it often assumed an ideal robot management system ignoring potential queuing delays by assuming constant travel times.
-This study relaxes this assumption and explores the quantitative relationship between RSS configuration parameters and system throughput.
-We introduce a novel robot traffic management method, named the rhythmic control for sorting scenario (RC-S), for RSS operations, equipped with an estimation formula establishing the relationship between system performance and configurations.
+In particular, autonomous mobile robots have become integral to modernization efforts in warehouses.
+One noteworthy application in robotic warehousing is the robotic sorting system (RSS), which is distinguished by its cost-effectiveness, simplicity, scalability, and adaptable throughput control.
+Previous research on RSS efficiency often assumed an ideal robot management system, ignoring potential traffic delays and assuming constant travel times.
+We introduce a novel robot traffic management method, named Rhythmic Control for Sorting Scenario (RC-S), for RSS operations, along with an estimation formula that establishes the quantitative relationship between system performance and configurations.
 Simulations validate that RC-S reduces average service time by 10.3% compared to the classical cooperative A* algorithm, while also improving throughput and runtime.
-Based on the performance analysis of RC-S, we further develop a layout optimization model for RSS, considering RSS configuration, desired throughput, and costs, to minimize expenses and determine the best layout.
-Numerical studies show that at lower throughput levels, facility costs dominate, while at higher throughput levels, labor costs prevail.
+Based on the performance analysis of RC-S, we develop a layout optimization model that considers system configurations, desired throughput, and costs to minimize expenses and determine the optimal layout.
+Numerical studies show that facility costs dominate at lower throughput levels, while labor costs prevail at higher throughput levels.
 Additionally, due to traffic efficiency limitations, RSS is well-suited for small-scale operations like end-of-supply-chain distribution centers.
@@
 In this paper, we focus on the sorting process.
-In a distribution center or warehouse where batching and/or zoning policies are implemented within the order picking process, sorting is required to split the batch and consolidate items per customer order or destination for shipping.
-Conventional sorting systems commonly employ sorting conveyors equipped with different sorter technologies.
-The sorters above move steadily, sequentially pass through each outbound station, and release the loaded parcels at suitable locations (Boysen et al., 2019a, ).
+Sorting involves categorizing and consolidating parcels according to their order information and shipping destinations.
+Conventional sorting systems commonly employ conveyor-based sorters , where the actuators move along with the conveyor belt, sequentially passing through each outbound station and releasing the loaded parcels at the appropriate locations (Boysen et al., 2019a, ).
 These systems are highly appreciated for its efficiency and stability, while suffering from the inflexibility and significant space occupation (Boysen et al.,, 2023).
@@
 A novel sorting system utilizing mobile robots attracts increasing attention in recent years, which is called the robotic sorting system (RSS) (Zou et al.,, 2021).
-Figure 1(a) presents a dual-layer implementation of RSS.
-The site is dotted with considerable number of outlets, each of which is associated with a specific sorting category, gathering all the packages of that category.
-The hardware for the outlet can be a bin, a cage cart, or a chute on the floor, which connects to the packing equipment on the lower level.
-The loading stations are located on the periphery of the site, where workers are responsible for loading parcels onto the robots.
-Sorting robots are the devices responsible for carrying out the sorting operations, serving as the conveyor belts in conventional sorting systems.
-A top-mounted tray or conveyor device allows the robot to load or unload shipments.
+Figure 1(a) presents a dual-layer solution of RSS.
+The site is arranged with numerous delivery ports, each associated with a specific sorting category, gathering all the parcels of that category.
+A delivery port can be a bin, a cage cart, or a chute in a dual-layer structure.
+Robots deliver parcels to delivery ports, serving as the combination of conveyor belts and actuators in conventional sorting systems.
+Loading stations are located on the periphery of the site, where workers load parcels onto the robots.
+A top-mounted tray or conveyor actuator enables the robot to load or unload parcels.
 Figure 1(b) shows a specific model of sorting robot developed by Geekplus, a global technology company specialized in smart logistics and robots.
-The operational cycle of a sorting robot is as follows:
+The process for a robot to execute a sorting task is as follows:
   1. 1.
@@
     The robot returns to the waiting zone behind a loading station and becomes idle.
-The independent and modular nature of robotic operations enables the system design process to become more flexible.
-During off-hours break, warehouse managers can easily modify the control software to change the positions of loading stations, correspondences between outlets and destinations, and the topology of the aisle network (Xu et al.,, 2022).
-Compared with conveyor belt equipment, RSS possesses the advantage of scalability and throughput flexibility, which are highly suited to the high-demand and dynamic nature of the current e-commerce environment (Azadeh et al.,, 2017).
-To promote this advantage, robotics companies introduce an innovative business model called robots-as-a-service, in which robotics companies offer customers the opportunity to rent robots for use in manufacturing.
-As a result, the customers no longer have to contend with the high costs typically associated with purchasing advanced automation tools outright.
-In RSS, warehouse manager can adjust the number of robots based on sorting demand fluctuations, thereby reducing operations costs.
-However, the flexibility of robot system also brings additional operational challenge, namely the multi-agent path finding (MAPF) problem, which aims to resolve the potential conflicts among the trajectories of robots.
+E-commerce order fluctuate during special selling seasons; therefore, warehouse throughput must be designed to accommodate these variations.
+Unlike conveyor-based sorting systems, the independent and modular nature of robotic operations contributes to the flexibility and scalability of RSS, making it well-suited to dynamic demands of e-commerce (Azadeh et al.,, 2017).
+During off-hours, warehouse managers can easily configure the control software to change the status of loading stations, the correspondences between outlets and destinations, and the aisle network topology (Xu et al.,, 2022).
+To leverage the flexibility, robotics companies introduce the innovative Robots-as-a-Service (RaaS) business model, allowing logistics clients to rent robots as needed.
+RaaS eliminates the high costs associated with purchasing and maintaining robots and peripheral equipment, enabling warehouse managers to adjust the number of rented robots based on demand fluctuations.
+Despite these advantages, a research gap remains in investigating the flexibility of RSS and minimizing system costs, as accurately modeling system efficiency is challenging.
+In robotic systems, robots rely on path planning and coordination algorithms to complete tasks.
+The problem of finding conflict-free trajectories for all robots is known as the multi-agent path finding (MAPF) problem.
 The MAPF problem has been proven to be NP-hard (Yu and LaValle,, 2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in the instability of RSS efficiency.
-Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in an RSS, particularly when dealing with a substantial parcel flow.
-Existing studies often overlook issues related to robot coordination, making the idealized assumption that there are no conflicts or deadlocks.
-Additionally, these studies tend to assume a fixed route travel time excluding the potential queuing delays.
-As a result, their outcomes may become distorted, especially in situations of high traffic load, as the actual travel time of the robots deviates from the free-flow travel time.
+Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in RSS, particularly when dealing with a substantial parcel flow.
+However, existing researches often overlook issues related to robot coordination, making the idealized assumption that there are no conflicts or deadlocks.
+Additionally, most of these studies tend to assume a fixed route travel time, excluding potential queuing delays on the road.
+As a result, their outcomes may become distorted, as the actual travel time of the robots deviates from the free-flow travel time.
+Furthermore, these studies do not adequately explore the flexibility of RSS, particularly the ability to adjust the number of robots.
 Given the existing gap, this paper aims to address critical challenges in designing an efficient RSS, focusing on three key areas.
 First, we seek to ensure the system operates both stably and efficiently in a setting where multiple robots collaborate, a non-trivial challenge given the complexities of robot interaction.
-Second, the development of a precise model to quantify congestion effects within the RSS is essential.
+Second, the development of a precise model to quantify system efficiency within the RSS is essential.
 Such a model will enable accurate throughput estimation, which is vital for understanding the relationship between the size of the facility and its sorting capacity.
-Third, we emphasize the need for management strategies that exploit the flexibility of RSS to adapt to fluctuating demands cost-effectively.
+Third, we emphasize the need for management strategies that leverage the flexibility of RSS to achieve cost-effective responses to fluctuating demands.
 These targeted areas of improvement aim to enhance the overall performance and scalability of RSS implementations.
-To address the aforementioned challenges and achieve the system optimum, we first explore the potential of centralized control strategy for sorting robots.
-To fully utilize the global information of the robot control system, taking into account robot operational patterns, and safety distance regulations, we adopt a network-level control strategy for the online coordination of autonomous vehicles and propose a collision-free spatio-temporal path planning method for RSS.
-Compared to existing MAPF methods, it does not require high communication frequency and tight latency between the central host and each robot.
-Additionally, we derive an efficient heuristic algorithm to enhance system scalability.
-Second, building upon this control strategy, under any given system configurations, we conduct a quantitative assessment of the average travel distance of robots and the maximum throughput of the RSS.
-We introduce an estimation formula that elucidates the relationship between system performance metrics and configuration settings.
-Finally, we propose an optimization model aimed at minimizing the initial investment and average operations costs under variable demand.
-The solution includes layout design and resource allocation recommendations.
-Experiments are conducted to provide several insights about the RSS, such as the cost structure of the RSS across varying unit price conditions.
+To address the aforementioned challenges, we first explore the potential of centralized control strategy for sorting robots.
+To fully utilize global information while considering robot operational patterns and safety distance regulations, we adapt a network-level control strategy for coordinating autonomous vehicles and propose a collision-free spatio-temporal path planning method for RSS.
+An efficient heuristic algorithm is proposed to enhance the scalability.
+Second, building upon this control strategy, we derive an estimation formula for throughput in RSS given system configurations, which serves as an efficiency constraint in the planning problem.
+Finally, we propose an optimization model aimed at minimizing the initial investment and average operations costs under fluctuating demands.
+The solution includes layout design and resource allocation recommendations, specifically on how to adjust the number of robots and workers.
+Experiments provide insights into the cost structure of RSS under varying unit price conditions.
 To the best of our knowledge, this paper is among the first to incorporate the traffic issues of robots during operations into the configuration design of the RSS system.
-Our intention is to bridge operational and strategic planning in multi-robot coordination contexts, and hence contribute a comprehensive theoretical framework to support decision-making in such complex systems.
+Our intention is to bridge operational and strategic planning in multi-robot coordination contexts, thereby contributing a comprehensive theoretical framework to support decision-making in such complex and flexible systems.
 The remainder of this paper is organized as follows.
@@
 Based on the proposed framework, section 5 presents a system throughput estimation formula, which provides quantitative measurement of layout configuration’s impact to system efficiency.
-In section 6, we demonstrate the superiority of our proposed traffic management framework compared with benchmarks, as well as validate the throughput estimation formula, through comprehensive experiments.
+In section 6, we demonstrate the superiority of our proposed traffic management framework compared with benchmarks, as well as validate the accuracy of throughput estimation formula.
+To emphasize the impact of traffic issues on RSS, experiments are conducted to show that queueing network model has estimation biases for throughput in certain scenarios.
 Moving on to section 7, we propose a layout optimization model to minimize the total costs and introduce an efficient solution approach.
@@
 2 Literature Review
-The cost savings in labor through automation are particularly crucial in the picking process, and higher sorting depth enables parcel companies to offer convenient and low-cost same-day delivery services (Dekhne et al.,, 2019).
+The cost savings in labor through automation are particularly crucial in the picking process, and increased sorting efficiency enables parcel companies to offer convenient and low-cost same-day delivery services (Dekhne et al.,, 2019).
 However, the operations research community has scarcely explored the realm of robotized sorting systems, which represent the latest advance of warehouse automation.
-Among the existing literature, studies mainly focus on small parcel sorting systems in distribution centers.
-Boysen et al., (2023) conducted research on order sorting system, which involves additional tasks including Piece-to-Order assignment and Order-to-Collection Point assignment.
-Therefore, the primary focus of the referred study is not solely on scheduling and control of robots, and the travel time of robots to a specific outlet is assumed to be constant.Shi et al., (2021)’s study, on the other hand, investigates a human-robot hybrid sorting system, in which the scale of robot sorting is fixed, while the manual capacity is adjustable according to the demand.
-Regrettably, in this context, the considerable flexibility inherent in the robotic system has not been fully harnessed.
-In addition to the two aforementioned references, the remaining literature on robotic sortation, including this study, reorients the focus back to the original RSS itself.
+Among the existing literature, studies mainly focus on the efficiency of small parcel sorting systems in distribution centers.
 This body of work delves into the comprehensive assessment of transportation efficiency within the robot fleet under various system configurations.
@@
 Zi and Gao, 2020a ; Zi and Gao, 2020b assumed that each robot moves independently.
-Zou and Chen, (2020) introduced a semi-open queuing network (SOQN) model and subsequently improved its theoretical aspects in their follow-up research (Zou et al.,, 2021).
-The SOQN models in these two studies take into account the queuing behavior of robots at loading stations and outlets, partially addressing the issue of robot occupancy on the aisle.
+Zou and Chen, (2020) introduced a semi-open queuing network (SOQN) model and subsequently improved its theoretical aspects in their follow-up research (Zou et al.,, 2021), simplifying it to the closed queueing model (CQN) by assuming a sufficient high arrival rate of parcels.
+A queueing model takes into account the queuing behavior of robots at loading stations and outlets, partially addressing the issue of robot occupancy on the aisle.
 Furthermore, their second study compares how the minimum total costs of the system varied under different network topology and different unit cost of system components, given the target sorting efficiency.
 Xu et al., (2022) further investigated an RSS with a parcel-to-loading-station assignment mechanism, which could be regarded as a pre-sorting strategy.
-To simplify the model, the study removes the constraint on the number of robots, resulting in an open queuing network (OQN) model.
+This study removes the constraint on the number of robots, resulting in an open queuing network (OQN) model.
 Experimental results demonstrates that the introduction of pre-sorting effectively reduces the average travel distance of the robots but might lead to extra congestion upstream at the same time.
-Additionally, Liu et al., 2019b and Tan et al., (2021) focused on the task assignment problem with the objective of minimizing the sorting makespan.
+Liu et al., 2019b and Tan et al., (2021) focused on the task assignment problem with the objective of minimizing the sorting makespan.
 Both adopts a travel time model (closed-form travel time expression, defined by Azadeh et al., (2017)), and the second research formulates a mixed-integer programming model.
 Compared to queuing network models, travel time models exhibit significant distortion when the number of robots is high, as they allow multiple robots to simultaneously occupy a facility.
+Boysen et al., (2023) conducted research on order sorting system, which involves additional tasks including Piece-to-Order assignment and Order-to-Collection Point assignment.
+Therefore, the primary focus of the referred study is not solely on scheduling and control of robots, and the travel time of robots to a specific outlet is assumed to be constant.
+Shi et al., (2021)’s study, on the other hand, investigates a human-robot hybrid sorting system, in which the scale of robot sorting is fixed, while the manual capacity is adjustable according to the demand.
+Regrettably, in this context, the considerable flexibility inherent in the robotic system has not been fully harnessed.
 Table 1: Literature summary on RSS investigation
@@
 Throughout the existing research, a limited number of studies consider the floor space cost while others relax the constraints on aisle resources.
-Furthermore, finding collision-free paths for multiple vehicles has been proved to be an NP-hard problem (Surynek,, 2010), and some state-of-the-art algorithms (such as the CROTPN-based model in Sun et al.,, 2018) also struggle to completely eliminate the issue of road congestion resulting from an increasing number of robots.
-However, except the study by Shi et al., (2021), which employs a fitting method to predict the positive correlation between traffic flow and congestion, most of the studies neglect the mutual influence among multiple robots in aisle occupation for a compact system, leading to an overestimation of aisle capacity.
+Furthermore, finding collision-free paths for multiple vehicles has been proved to be an NP-hard problem (Surynek,, 2010), and some state-of-the-art algorithms also struggle to completely eliminate the issue of road congestion resulting from an increasing number of robots.
+However, except the study by Shi et al., (2021), which employs a fitting method to predict the positive correlation between traffic flow and congestion, most of the studies neglect the mutual influence among multiple robots in aisle occupation for a compact system, leading to an misestimate of traffic capacity.
 Table 1 presents a comprehensive overview of the literature reviewed on the RSS, highlighting its relevance to our paper.
-Next, we review the performance analysis and layout design of sorting systems based on conveyors.
-The study by Russell and Meller, 2003a comprehensively investigates the layout design problem in an order fulfillment system.
-Their descriptive model analyzes the impact of the number of workers and workstations on the system’s capacity, considering the costs of labor and equipment, as well as providing appropriate system configuration based on target throughput requirement.
-The study also validates their analytical results through simulation experiments and concluded the advantages of conveyor belt sorting compared to manual sorting.
-Bartholdi and Gue, (2000) studied the scenario of a cross-docking terminal, where freight needs to undergo sorting on a dragline conveyor system from strip doors to stack doors.
-They accounted for three types of congestion in the system and employed a simulated annealing method to examine sorting efficiency and labor costs across various layouts.
-Besides, the postal service industry faces greater challenges, including higher volume of shipments and tighter time window.
-Fedtke and Boysen, (2017) examined how design alternatives within the distribution center affect the performance of sorting system.
-The authors devised an integer programming model to optimize the minimum service distance under various configurations.
-Additionally, they developed heuristic algorithms for solutions.
-Similar methodologies have been applied in several other researches (Briskorn et al.,, 2017; Zenker and Boysen,, 2018).
-Readers are recommended to refer to Boysen et al., 2019a for an in-depth understanding of sortation conveyor systems.
-In addition to sortation systems, there are a sufficient number of studies on RMFS system in modern order picking process, which also consider the multi-robot management problem.
-An important operational metric of order picking is the proportion of effective working time for workers, and thus, related studies often utilize queuing network models to characterize the uncertainty in RMFS.
-Lamballais et al., (2017) proposed a SOQN model and provided insights into how performance is affected by factors such as the length-to-width ratio, workstation locations, number of robots.
-This modeling approach effectively captures the system bottleneck associated with robot blocking at workstations, which is validated through simulation.
-This modeling approach has proven instrumental in addressing challenges related to battery recovery, robot assignment, inventory allocation, and multi-deep layout (Zou et al.,, 2018; Roy et al.,, 2019; Tessensohn et al.,, 2020; Yang et al.,, 2022).
-Certain studies have relaxed the constraint on the number of robots, depending on the focus of their research (open queuing network: Yuan and Gong,, 2017, closed queuing network: Nigam et al.,, 2014; Roy et al.,, 2019).
-However, some researchers have recognized the limitations of queuing models in capturing the impact of aisle design on system throughput, and have chosen to address the problem from the perspective of MAPF problem in their studies.
-Wagner and Choset, (2011) proposed a novel M* algorithm that allowed for the planning of a larger number of robot trajectories compared to the original A* algorithm.
-However, they faced challenges due to significant memory requirements and time complexity, and could not guarantee a solution.
-Some researchers designed priority-based algorithms and integrated task assignment to reduce waiting time (Nguyen et al.,, 2019; Liu et al., 2019a, ), sacrificing algorithm optimality in exchange for stability and computational efficiency.
+Some researchers recognized the limitations of queuing models in capturing the impact of aisle layout design on multi-robot systems and chose to address the traffic issues by modeling them as MAPF problems.
+Wagner and Choset, (2011) proposed a novel M* algorithm that allowed for the planning of a larger number of robot trajectories.
+However, they faced challenges due to significant memory requirements and time complexity, and could not guarantee a feasible solution.
+Some researchers designed priority-based algorithms and integrated task assignment to reduce waiting time (Nguyen et al.,, 2019; Liu et al., 2019a, ), sacrificing solution optimality in exchange for stability and computational efficiency.
 Li et al., (2021) introduced an RHCR algorithm framework to tackle the windowed MAPF problem, providing an efficient solution for managing newly arrived orders in warehouse operations.
-The algorithm demonstrates scalability up to 1,000 agents and yields high-throughput solutions.
-However, despite their effectiveness, MAPF-based studies have encountered challenges in deriving a closed-form expression for throughput estimation, a critical element for robust system configurations.
-In summary, there are few studies on RMFS that comprehensively address both operational robot management problems and strategic warehouse design, with considering their interplay.
-We aim for the findings and framework of this study not only to be applicable to the RSS system, but also to be extendable to these multi-robot systems in future studies.
+However, despite their effectiveness, MAPF-based studies have faced challenges in deriving a closed-form expression for throughput estimation, which is crucial for creating reliable and efficient system configurations.
+Consequently, few studies on multi-robot systems comprehensively address both operational robot management and strategic layout design while considering their interplay.
+In summary, existing research on RSS systems lacks sufficient focus on robot traffic management and overall system costs.
+This study addresses these gaps by first developing a dedicated robot traffic management method for RSS at the operational level.
+Building on this method, we establish an optimal layout and configuration design model aimed at minimizing overall system costs in the long run, ensuring that the results are practically applicable in real-world warehouses.
 3 Problem Description
-Figure 2: Layout of RSS
 In an RSS, three zones are deployed from center to periphery in the field (Wang et al.,, 2021): 1) sorting zone; 2) loading zone; 3) waiting zone, as shown in Figure 2.
@@
 The decision of site planning mainly focuses on the design of sorting zone, particularly the configuration of the aisle network.
+Figure 2: Layout of RSS
 Without loss of generality, we make the following assumptions in this study:
@@
 That is, the loading time is negligible, and this parameter is considered implicitly in the loading rate of worker.
-  • The arrival rate of parcels destinating for each outlet at each loading station is the same [^2] When the sorting demand is unevenly distributed, this assumption can be achieved by increasing or decreasing the number of outlets corresponding to each destination.;
-  • The time for robots dropping parcels into outlets is negligible;
-  • Each outlet allows at most four robots in different directions releasing parcels simultaneously, and the drop-off operation by robots around one outlet does not affect each other;
+  • The arrival rate of parcels destinating for each outlet at each loading station is the same (When the sorting demand is unevenly distributed, this assumption can be achieved by increasing or decreasing the number of outlets corresponding to each destination.
+Additionally, outlets with extremely low and high arrival rates can be placed together to balance the flow);
+  • The time for robots dropping parcels into outlets is negligible (To satisfy this assumption, robots need to release goods in advance based on its speed before reaching the target outlet, similar to the behavior of conveyor-based sorters.
+If it is difficult to achieve due to technical and cost issues, a 2-second drop-off time per delivery must be introduced (Zou et al.,, 2021));
+  • The drop-off operation by robots around one outlet does not affect each other, i.e., each outlet allows at most four robots releasing parcels simultaneously;
   • Each robot can only carry one parcel during one delivery process;
-  • Workers are evenly distributed in the loading zone;
-  • The transportation of parcels after being collected in the outlet is not considered in this research, which may affect the operations stage for RSSs with different physical structure; samples of single-tier and two-tier design are provided by Zou et al., (2021);
+  • Loading stations are evenly distributed in the loading zone;
+  • The transportation of parcels after being collected in the outlet is not considered in this research.
   • The battery charging needs of robots are not considered.
 (a) Sample gird layout of a compact $4+4$ aisle network
-(b) Topology of sample network
+(b) Graph representation of sample layout
 Figure 3: Illustration of aisle network in sorting zone
 To better evaluate the performance of this system with high loads and complicated sorting demands, the microscopic traffic management of robots in the sorting zone needs careful investigation.
-Notably, we use the term ”aisle” instead of ”road” in the following content to clarify that each link in the network is single-lane.
-A sample grid network with compact aisle layout is depicted in Figure 3(a).
+Notably, we use the term “aisle” instead of “road” in the following content to clarify that each road in the network is single-lane.
+A sample grid-based map in RSS is depicted in Figure 3(a).
 It consists of $n_{v}$ vertical aisles and $n_{h}$ horizontal aisles.
-Unit cells are evenly distributed along each aisle and serve for robot localization (in some robotic warehouses, distinct unit cells can be identified by QR codes).
-The unit cell is a square with a side length of $D$.
-We assume the length of a robot is not greater than $D$.
-It allows robot to spin turn in a grid.
-Outlets collect the parcels dropped by robots.
-Outlets are evenly distributed over the sorting zone and divided by aisles, and each outlet occupies one grid.
-This layout ensures that robots move along aisles can directly drop parcels into outlets without additional turning, while also achieving sufficient density.
-Figure 3(b) shows the topology of sample layout in Figure 3(a).
-Conflict nodes (blue dot) connect the vertical and horizontal aisles, and unloading nodes (red hexagon) are the points where robots release parcels.
-To improve traffic efficiency, all aisles are one-way and the directions of any two adjacent aisles are opposite, which is a common strategy in real.
-To ensure the connectivity of the map, each loading station in loading zone has independent entrance and exit, and the number of horizontal and vertical aisles should be even.
-Hence, the number of loading stations is no more than the number of aisles.
-Figure 4: Common types of conflicts: (a) a swapping conflict, (b) a vertex conflict, (c) a following conflict, and (d) a cycle conflict
-The sorting robots share the same kinematic parameters, featuring a maximum velocity of $v_{max}$ and maximum acceleration/deceleration of $c_{max}$.
-Additionally, these robots possess the capability of in-place rotation, characterized by an average angular velocity denoted as $\omega_{r}$, with no consideration for turning acceleration or deceleration.
-To ensure collision-free movement of robots, the robot management system needs to avoid conflicts between different robots during path finding.
-Common conflicts, as illustrated in Figure 4, include the following four types (Stern et al.,, 2019): (a) Swapping conflict occurs when two robots are planned to swap locations at the same time; (b) Vertex conflict occurs when more than two robots are planned to occupy the same vertex; (c) Following conflict occurs when one robot is planned to occupy a vertex that was occupied by another robot in the previous timestep; (d) Cycle conflict occurs among multiple robots, with each robot having a following conflict with another robot.
-Among these conflicts, both swapping conflicts and cycle conflicts necessitate a re-planning of paths.
+Each cell in the map has a side length of $D$.
+Robots locate themselves using QR codes at the center of each cell and can only move between adjacent cells.
+We assume that the length of sorting robots is less than $D$, allowing them to rotate within a cell.
+We use the term “outlet” to refer to the special cells designated for collecting parcels delivered by robots.
+Outlets are evenly distributed throughout the sorting zone, divided by aisles.
+This layout allows robots to deliver parcels to outlets on both sides of the aisle and ensures that the outlets are arranged with sufficient density.
+Each loading station in the loading zone is connected to a pair of dedicated entrance and exit, which are adjacent to each other.
+Therefore, the number of loading stations cannot exceed the the number of aisles.
+Figure 3(b) shows the graph representation of the sample layout in Figure 3(a).
+In the graph, each cell that makes up the aisles is modeled as either a conflict node (blue dot), unloading node (red hexagon), or entrance/exit node (green/yellow square), with edges of length $D$ connecting adjacent nodes.
+Conflict nodes are at the intersections of aisles, unloading nodes are where robots deliver parcels, and entrance/exit nodes are at aisle ends.
+Figure 4: Common types of conflicts: (a) a swapping conflict, (b) a vertex conflict, (c) a following conflict, and (d) a deadlock conflict
+The sorting robots share the same kinematic parameters, with a maximum velocity of $v_{max}$ and maximum acceleration/deceleration of $c_{max}$.
+Robots possess the capability of in-place rotation, with angular velocity denoted as $\omega_{r}$.
+To ensure collision-free movement of robots, the robot management system needs to avoid conflicts in path finding.
+Common conflicts, as illustrated in Figure 4, include the following four types (Stern et al.,, 2019): (a) Swapping conflict occurs when two robots are planned to swap locations at the same time.
+It does not exist in a one-way network; (b) Vertex conflict occurs when more than two robots are planned to occupy the same vertex; (c) Following conflict occurs when one robot is planned to occupy a vertex that was occupied by another robot in the previous timestep.
+The latter is stationary or moving at a lower speed; (d) Deadlock conflict occurs among multiple robots, with each robot having a following conflict with another robot.
+Among these conflicts, both swapping conflicts and deadlock conflicts necessitate a re-planning of paths.
 The occurrence of either can result in significant losses in system efficiency.
+Figure 5: Overview of research methodology
 Planing problems related to RSS can be divided into two stages: the site planning stage and operations stage.
-In the site planning stage, decision maker concentrates on the division of zones for different functions in the warehouse.
+In the site planning stage, warehouse manager concentrates on the division of zones for different functions in the warehouse.
 The area of an RSS should be large enough to fulfill the designated throughput level in peak season, and the number of outlets should be no fewer than the number of individual sorting destinations.
-The decision variables of site planning stage are the number of horizontal and vertical aisles $n_{h},n_{v}$ and the number of loading stations $n_{l}$; in the operations stage, decision maker should adaptively adjust the number of workers and robots to minimize operations costs, given real-time sorting demands.
-As a warehouse is usually rented for years, it is advisable to keep it unchanged during this stage.
-The decision variables of operations stage are the number of workers and robots in peak season $n_{w}^{H},n_{r}^{H}$ and in off-peak season $n_{w}^{L},n_{r}^{L}$, where $L$ and $H$ in superscript indicate that the sorting demand is at low and high levels, respectively; and $w$ and $r$ in subscript represent worker and robot, respectively.
+The decision variables of site planning stage are the number of horizontal and vertical aisles $n_{h},n_{v}$ and the number of loading stations $n_{l}$; in the operations stage, warehouse manager should adaptively adjust the number of workers and robots to minimize operations costs, given expected sorting demands.
+Since a warehouse is usually rented for years, it is advisable to keep the zone division unchanged during its operation.
+The decision variables of operations stage are the number of workers and robots in operating period $\sigma$, noted as $n_{w}^{\sigma},n_{r}^{\sigma}$, respectively.
 The goal of RSS planning problem is to minimize the total costs, which consists of two parts: the facility costs $C_{f}$, and the operations costs $C_{o}$.
-The size of the site holds a trade-off between $C_{f}$ and $C_{o}$: With increased demand, a small sorting zone may lead to congestion, thereby increasing the service time for robots and raising the rental cost of robots in operations costs; whereas, expanding the size of RSS would escalate the facility costs.
-In sections 4-6, we will respectively present the traffic management of robots, conduct system performance analysis under the proposed traffic management framework, and validate the proposed algorithm and performance estimation formula.
-In section 7, a layout design optimization model is formulated to minimize the total costs of RSS with given demand.
+The size of the site holds a trade-off between $C_{f}$ and $C_{o}$: A small network leads to lower capacity, thereby increasing the waiting time of robots and raising the rental cost of robots in operations costs; whereas, an overly large network leads to increased site costs and longer travel distance.
+An overview of research methodology is shown in Figure 5.
+In sections 4-6, we present the traffic management of robots, conduct system efficiency analysis, and validate the performance and accuracy of estimation formula, respectively.
+In section 7, a layout design optimization model is formulated to minimize the total costs of RSS given the demand.
 4 Management of High-density Robot Traffic in RSS
-Managing a large number of robots within a dense network poses significant challenges, especially when striving to balance system throughput performance with the computational load of algorithms under unpredictable traffic demands.
-In order to address multi-robot path finding and eliminate traffic gridlock within an admissible computational duration, this study adopts an innovative autonomous vehicle management scheme, called Rhythmic Control (RC) (Lin et al.,, 2021).
+Managing a large number of robots within a dense network poses significant challenges, especially when balancing system throughput performance with the computational load of algorithms under unpredictable traffic conditions.
+To address multi-robot path finding and eliminate traffic gridlock within an admissible computational duration, this study adopts an innovative autonomous vehicle management scheme, called Rhythmic Control (RC) (Lin et al.,, 2021).
 It enables the planning of uninterrupted scheduling for robot fleets.
-In this section, we start with a motivation example to illustrate why introducing rhythm into robot management is necessary.
-Subsequently, we summarize the key concepts inherited from the original RC.
-By leveraging the node-reservation idea in classical MAPF solver, we propose a new centralized control framework for sorting scenario, i.e., the RC-S scheme.
-RC-S presents high orderliness in practical performance, serving as the foundation for theoretical analysis of system efficiency, which will be discussed in detail in section 5.
-4.1 Motivation example
-In conveyor-based sorting systems, there is always a mainline through which goods move.
-When passing through the target outlets, the goods are pushed out by barriers to complete the sorting process.
-We aim to combine the stability of conveyor-based sorting systems with the flexibility and cost-effectiveness of RSS.
-Consider a conveyor system comprising two distinct types of conveyor segments: Type A, with a set length and flow speed, and Type B, featuring a turntable capable of altering directions as depicted in Figure 5(a).
-It is assumed that the transit time for goods on segment A is odd, whereas for conveyor B, this time is essentially negligible.
-Positioned along the mainline’s direction, the indices of segment B are sequentially labeled $1,2,3,$ and so forth.
-The orientation of each B segment coincides with the mainline when its index shares the same parity as the current time; if not, it shifts to align with the branch line.
-Consequently, by initiating the loading of goods onto the main line exclusively at even-numbered intervals, the goods will seamlessly proceed along the mainline.
-To divert them from the mainline, we could activate a barrier on segment B, blocking the goods for one unit of time when they reach it.
-By controlling the state of all the barriers, we can divert the goods to different branch lines.
-The production rhythm of such line is two units of time.
-(a) Mainline-branch structure
-(b) Network of conveyor segments
-Figure 5: Conveyor system with two types of conveyor segment
-The rhythmic manner remains applicable when applying the aforementioned control method to a four-line network in Figure 5(b), where adjacent lines always have opposite directions.
-Each line maintains consistent throughput when all the barriers are inactive.
-Activating a barrier at the appropriate time allows goods to switch to another line.
-However, this may conflict with the original rhythm of goods on the new line.
-We need a scheduling and reservation mechanism to avoid such conflicts.
+In this section, we begin with the key concepts inherited from the original RC: the virtual platoon and the cycle.
+By leveraging the reservation mechanism, we propose a new centralized framework called Rhythmic Control in Sorting Scenario (RC-S).
+RC-S presents high orderliness and serves as the foundation for theoretical analysis of system efficiency, which will be discussed in detail in section 5.
 4.1 Virtual platoon and cycle
-If we use robots to replace the functionality of conveyors in the motivation example, then the problem becomes a MAPF issue.
 Lin et al., (2021) investigated a method of incorporating rhythm into traffic management at the macro-traffic level and introduced the concept of virtual platoon (VP).
-VPs represent the spatiotemporal slot that is generated in a rhythmic manner and is available for being occupied by vehicles.
-As shown in Figure 6(a), VPs move along a designated spatiotemporal trajectory and maintain a constant distance between each other, resembling an endless series of train carriages connected by invisible ropes of equal length in each road.
-We discretize time based on the generation time of VPs at the entrance, and we refer to each segment as a cycle.
-Each entrance generates VPs at a predetermined pace, noted as the cycle length, which is equal to the time gap between the arrival time of two consecutive VPs at each node in the network.
+VPs represent the spatiotemporal slot that is generated in a rhythmic manner and keep uniform linear motion.
+As shown in Figure 6(a), VPs move along a designated spatiotemporal trajectory and maintain a constant safe distance between each other.
+Figure 6(b) shows the spatiotemporal trajectory of VPs in a four-lane network, without any collision or deceleration.
+We discretize time based on the generation time of VPs at the entrance, and each time interval is noted as a cycle.
+The duration of each time interval is the cycle length.
 By staggering the entry times of VPs at different entrances within the same cycle, VPs pass one intersection alternately with relatively short headway and the intersection capacity could be maximized.
-Figure 6(b) shows the spatiotemporal trajectory of VPs in a four-lane network, without any collision or deceleration.
-Whenever a vehicle needs to pass through the network, the system reserves one unit of capacity from the corresponding VPs, and then releases it with an appropriate start cycle.
-Then that vehicle enters the network, following the movement of certain VPs within each route segments to reach its destination, and exits the network.
+Whenever a vehicle needs to pass through a lane, it is mandated to follow the movement of one VP in that lane until it reaches its destination and exits the lane.
+Before entering the network, the vehicle must wait for an unoccupied VP at the entrance.
+It first sends its destination to the control system, which then plans a conflict-free route and reserves one unit of capacity from the corresponding VPs in different segments of the route.
+The capacity is released when the vehicle leaves each VP.
 Through the reservation mechanism, RC ensures that the number of vehicles within each VP does not exceed its predetermined capacity.
-Therefore, the trajectories are conflict-free without any blockage.
+Consequently, the trajectories of vehicles are conflict-free.
 (a) Virtual platoons in a four-lane network
@@
 Figure 6: Concept of Rhythmic Control
-RC is essentially a traffic control method, akin to a combination of globally coordinated traffic signal control and road network perimeter control.
+RC is essentially a traffic control method, similar to a combination of globally coordinated traffic signal control and road network perimeter control.
 Within this framework, it is essential to clarify the movement patterns of each entity for integration into a sorting-specific robot management system.
-For detailed technical information on the RC strategy, we refer readers to Lin et al., (2021).
+For detailed information on the RC framework, we refer readers to Lin et al., (2021).
 The following points provide a foundational overview critical to our study on the implementation of the RC-S scheme:
@@
 4.2 Concept of RC-S
-In order to embody RC in the RSS, we regulate the behavior of robots, and propose the scheme of rhythmic control for sorting scenario, abbreviated as RC-S.
-RC-S can be regarded as a centralized MAPF algorithm.
+In order to embody RC in the RSS, we regulate the behavior of robots, and propose the RC-S scheme, which can be regarded as a centralized MAPF solver.
 To focus on the behavior of each individual robot, we discretize the map into grids, defining each grid where robots can pass through as a node.
-Given $(n_{v},n_{h})$, the network in sorting zone is represented by a directed graph $\mathcal{G}=(\mathcal{V}_{c},\mathcal{V}_{u},\mathcal{A})$, where $\mathcal{V}_{c}$, $\mathcal{V}_{u}$ and $\mathcal{A}$ stand for the set of conflict nodes, unloading nodes and links, respectively.
-To reduce traffic conflicts and deadlocks, we adopt a one-way network, where aisles from different directions are arranged in an interleaved pattern, as shown in Figure 7.
-Furthermore, the behaviors of robots are limited to three actions: moving straight, dropping off parcel without stopping, and left turning.
-The joint movement inside a virtual platoon is not practical for robots.
-To avoid collision within a platoon, the capacity of each VP within a cycle is set to one robot.
+Given $(n_{v},n_{h})$, the network in sorting zone is represented by a directed graph $\mathcal{G}=(\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{e},\mathcal{E})$, where $\mathcal{V}_{c}$, $\mathcal{V}_{u}$, $\mathcal{V}_{e}$ and $\mathcal{E}$ stand for the set of conflict nodes, unloading nodes, entrance/exit nodes and edges, respectively.
+To reduce intersections, we adopt a one-way network, where aisles from different directions are arranged in an interleaved pattern, as shown in Figure 7.
+Based on the network structure, each VP can accommodate at most one AGV.
 Let $\tau_{e}$ and $\tau_{c}$ denote the fixed travel time of VPs on each unit cell in the grid map and the cycle length of rhythmic control, respectively.
-Under RC-S, robot enters the network through one aisle with fixed interval $\tau_{c}$, thus the upper bound of robot flow on each grid is $1/\tau_{c}$.
-In each cycle, the distribution of VPs could be divided into four phases (Figure 7(a)-(d)), and the duration of each phase is $\tau_{e}$.
-We have:
+Under RC-S, the behaviors of robots are limited to three actions: moving straight, dropping off parcel, and turning.
+Robots enter the network with fixed interval $\tau_{c}$, thus the upper bound of robot flow on each grid is $1/\tau_{c}$.
+Figure 7: Illustration of RC-S. (a)-(d) are the four phases in one cycle.
+In each cycle, the distribution of VPs could be divided into four phases (Figure 7(a)-(d)), and the duration of each phase is $\tau_{e}$. We have:
 $\tau_{c}=4\tau_{e}$ (1)
-Figure 7: Illustration of RC-S. (a)-(d) are the four phases in one cycle.
-$\tau_{e}$ and $\tau_{c}$ are determined by the maximum velocity of robot $v_{max}$ and the maximum loading rate of each loading station $r_{ls}$, respectively:
+$\tau_{e}$ and $\tau_{c}$ are determined by the maximum velocity of robot $v_{max}$ and the maximum loading rate of each loading station $r_{l}$, respectively:
 $\tau_{e}\geq\frac{D}{v_{max}}$ (2)
-$\tau_{c}\geq\frac{1}{r_{ls}}$ (3)
+$\tau_{c}\geq\frac{1}{r_{l}}$ (3)
 where $D$ is the length of each grid.
@@
 Figure 8: Illustration of turning and transfer process for one robot (red dot)
-In addition, auxiliary road for vehicle turning in the original RC is removed, since the robot is capable to spin at the intersection grid for turning.
-The process of turning and transfer between VPs are demonstrated in Figure 8, which takes six consecutive phases.
-As shown in the figure, the robot to turn first decelerates to zero, stops following its original VP, rotates 90 degrees at the intersection, joins a new VP, and accelerate to the speed of VP.
-Turning action entails four phases for both the original and subsequent VPs, totaling eight phases that are unavailable to other robots.
-However, the actual motion spans merely six consecutive phases, thus leading to a reduction in traffic capacity.
+RSS does not require auxiliary turning lanes since robots are capable to rotate within the intersection grid.
+The process of turning and transfer between VPs for robots are demonstrated in Figure 8.
+As shown in the figure, the robot to turn first decelerates to zero, stops following its original VP, rotates 90 degrees at the intersection, joins a new VP and follows the movement of it.
+Turning action entails one cycle for both the original and subsequent VPs, leading to a reduction in traffic capacity.
 During this process, the parameters of RC-S needs to satisfy the following conditions:
@@
 Constraint (5) implies that the robot can catch up with the VP in a turn, where $d(\tau_{e},v_{VP},v_{max},c_{max})$ denotes the maximum travel distance a robot can cover in acceleration/deceleration process.
-The expression is derived in detail in Appendix B; Constraint (6) imposes limits on the rotation speed.
-In summary, incorporating expressions (1)-(6) and the kinematic parameters of the robots, we could determine the parameters for RC-S.
+The expression is derived in detail in Appendix B; Constraint (6) imposes limits on the rotation speed $\omega_{r}$, ensuring that the robot can complete a 90-degree turn within one phase.
+In summary, using expressions (1)-(6) and the kinematic parameters of the robots, we could determine the parameters for RC-S.
 The trajectory of robot in one delivery task should start from one loading station, pass the target outlet corresponding to the destination of parcel, and end at another loading station.
@@
 A feasible path in RC-S is a sequence of VPs and their occupying cycles.
-The trajectories traveled by these VPs during the occupied cycles are connected end-to-end, linking two loading stations at the periphery of the network.
-When a robot performs a sortation task, it moves from one loading station $i$ to another loading station $j$, $i,j\in\mathcal{S}$, where $\mathcal{S}$ is the set of all loading stations.
+The trajectories traveled by these VPs during the occupied cycles are connected end-to-end, linking two active loading stations at the periphery of the network.
+Definition 1 defines the candidate spatio-temporal paths for robots under RC-S, where an active loading station indicates that a worker is currently engaged in tasks.
+When a robot performs a sorting task, it moves from one loading station $i$ to another loading station $j$, $i,j\in\mathcal{L}$, where $\mathcal{L}$ is the set of all loading stations.
 Let $\mathcal{O}$ represent the set of outlets.
-Then, we denote the set of all feasible path as $\mathcal{R}=\cup_{i\in\mathcal{S},j\in\mathcal{S},k\in\mathcal{O}}\mathcal{R}(i,k,j)$, where $\mathcal{R}(i,k,j)$ is the subset containing the paths start from loading station $i$, pass by outlet $k$ and end at loading station $j$.
-To simplify the notation, we use $\mathcal{R}_{i,k}$ to represent the set of all feasible paths originating from station $i$ and passing by outlet $k$, and ^$\hat{\mathcal{R}}_{i}$ to represent the set of all feasible paths heading for station $i$, that is, $\mathcal{R}_{i,k}=\cup_{j\in\mathcal{S}}\mathcal{R}(i,k,j)$ and ^$\hat{\mathcal{R}}_{i}=\cup_{j\in\mathcal{S},k\in\mathcal{O}}\mathcal{R}(j,k,i)$.
+Then, we denote the set of all feasible path as $\mathcal{R}=\cup_{i\in\mathcal{L},j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(i,k,j)$, where $\mathcal{R}(i,k,j)$ is the subset containing the paths start from loading station $i$, pass by outlet $k$ and end at loading station $j$.
+To simplify the notation, we use $\mathcal{R}_{i,k}$ to represent the set of all feasible paths originating from station $i$ and passing by outlet $k$, and ^$\hat{\mathcal{R}}_{i}$ to represent the set of all feasible paths heading for station $i$, that is, $\mathcal{R}_{i,k}=\cup_{j\in\mathcal{L}}\mathcal{R}(i,k,j)$ and ^$\hat{\mathcal{R}}_{i}=\cup_{j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(j,k,i)$.
 In each cycle, the feasible path assignment problem (FPA) for all the loaded robots is formulated as follows.
-$\underset{x_{i,k}^{r},\hat{x}_{i,k}}{\min}\ \sum_{i\in\mathcal{S}}\sum_{k\in\mathcal{O}}(\sum_{r\in\mathcal{R}_{i,k}}c_{i,k}^{r}x_{i,k}^{r}+\hat{c}_{i,k}\hat{x}_{i,k})$
-$s.t.$ $\sum_{i\in\mathcal{S}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}\delta_{i,k}^{r,\nu,e}x_{i,k}^{r}\leq N_{\nu}^{e}$ $\forall\nu\in\mathcal{V}_{c}\cup\mathcal{V}_{u},e\in\mathcal{E}$ (7)
-$\hat{x}_{i,k}+\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=d_{i,k}$ $\forall i\in\mathcal{S},k\in\mathcal{O}$ (8)
-$\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=\sum_{j\in\mathcal{S}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{j,k}\cap\hat{\mathcal{R}}_{i}}x_{j,k}^{r}$ $\forall i\in\mathcal{S}$ (9)
-$x_{i,k}^{r},\hat{x}_{i,k}\in\{0,1\}$ $\forall i\in\mathcal{S},k\in\mathcal{O},r\in\mathcal{R}_{i,k}$ (10)
+$\underset{x_{i,k}^{r},\hat{x}_{i,k}}{\min}\ \sum_{i\in\mathcal{L}}\sum_{k\in\mathcal{O}}(\sum_{r\in\mathcal{R}_{i,k}}c_{i,k}^{r}x_{i,k}^{r}+\hat{c}_{i,k}\hat{x}_{i,k})$
+$s.t.\quad$ $\sum_{i\in\mathcal{L}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}\delta_{i,k}^{r,\nu,l}x_{i,k}^{r}\leq N_{\nu}^{l}$ $\forall\nu\in\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{e},l\in\mathcal{C}$ (7)
+$\hat{x}_{i,k}+\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=d_{i,k}$ $\forall i\in\mathcal{L},k\in\mathcal{O}$ (8)
+$\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=\sum_{j\in\mathcal{L}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{j,k}\cap\hat{\mathcal{R}}_{i}}x_{j,k}^{r}$ $\forall i\in\mathcal{L}$ (9)
+$x_{i,k}^{r},\hat{x}_{i,k}\in\{0,1\}$ $\forall i\in\mathcal{L},k\in\mathcal{O},r\in\mathcal{R}_{i,k}$ (10)
 There are two sets of variables: $x_{i,k}^{r}$ equals one if the feasible path $r\in\mathcal{R}_{i,k}$ is reserved in current cycle by a robot, and zero otherwise; $\hat{x}_{i,k}$ equals one if the robot at loading station $i$ is heading for outlet $k$ and is kept waiting until the next cycle, and zero otherwise.
@@
 The objective function is to minimize the total costs in this cycle.
-Constraint (7) indicates that in each cycle, at most one robot can be present at node $\nu$, where $N_{\nu}^{e}$ is the pre-set remaining capacity in node $\nu$ in cycle $e$, $N_{\nu}^{e}\leq 1$.
-Parameter $\delta_{i,k}^{r,\nu,e}$ equals one if feasible path $r$ will occupy node $\nu$ in cycle $e$ and zero otherwise.
+Constraint (7) indicates that in each cycle, at most one robot can be present at node $\nu$, where $N_{\nu}^{l}$ is the updated remaining capacity in node $\nu$ in cycle $l$, $N_{\nu}^{l}\leq 1$.
+Parameter $\delta_{i,k}^{r,\nu,l}$ equals one if feasible path $r$ will occupy node $\nu$ in cycle $l$ and zero otherwise.
 Constraint (8) describes the path allocation results for a robot at loading station $i$, where $d_{i,k}$ equals one if the robot is heading for outlet $k$, and zero otherwise.
@@
 The entry cycle of a specific feasible path is the earliest cycle at which all VPs in the path are unoccupied.
+Algorithm 1 A heuristic method to implement RC-S
+0:  Load the reservation table that records the status of VPs on the map within $N_{p}$ cycles. Set the maximum search range in each cycle as $N_{t}$. Calculate feasible path set $\mathcal{R}$.
+0:  At the beginning of cycle $t$, add the first robot of each queue at entrance to the candidate list.
+0:  If the candidate list is empty, go to step 5; else, find the robot with the longest waiting time in the candidate list, obtain its origin $i$ and target outlet $k$.
+Calculate the entry cycle of each feasible path in $\mathcal{R}_{i,k}$, and choose the feasible path with the smallest entry cycle.
+0:  If the entry cycle is less than $N_{t}$, reserve the VPs according to the feasible path, and assign the path to the robot; else, add the waiting time of the robot by $\tau_{c}$.
+0:  Remove the robot from the candidate list and go to step 2.
+0:  Record the newly generated VPs in the reservation table, labeled as unoccupied. Remove the VPs which have left the network. Output the assignment of feasible path.
 So far, we have identified the essential components required to integrate heuristics into the RC-S framework: recording the status of VPs ever existed, and for each delivery task, finding a feasible path with an entry cycle, and reserving the VPs in the space-time map.
@@
 The optimal strategy of RC-S requires selecting a combination of feasible path and entry time for each robot, and checking the occupancy status of VPs.
-Algorithm 1 A heuristic method to implement RC-S
-0:  Set the maximum working cycle $N_{c}$.
-Use a reservation table to record the status of VPs on the map within $N_{p}$ cycles.
-Set the maximum search range in each cycle as $N_{t}$.
-Calculate feasible path set $\mathcal{R}$.
-0:  Set $t=1$;
-0:  At the beginning of cycle $t$, add the first robot of each queue at entrance to the candidate list.
-0:  If the candidate list is empty, go to step 6; else, find the robot with the longest waiting time in the candidate list, obtain its origin $i$ and target outlet $k$.
-Calculate the entry cycle of each feasible path in $\mathcal{R}(i,j,k)$, and choose the feasible path with the smallest entry cycle.
-0:  If the entry cycle is less than $N_{t}$, reserve the VPs according to the feasible path; else, add the waiting time of the robot by $\tau_{c}$.
-0:  Remove the robot from the candidate list and go to step 3.
-0:  $t=t+1$. If $t>N_{c}$, end; else, add new VPs at the entrance of each aisle, labeled as unoccupied in the reservation table. Remove the VPs which have left the network. go to step 2.
 Algorithm 1 has a complexity of $O(N_{t}\cdot\lvert R\rvert)$ in each cycle.
@@
 5 Performance Analysis of RC-S
-In a multi-robot system, the mutual influence among robots often leads to suboptimal utilization of roadway resources.
-In an ideal scenario, without any robot making turns and merely traveling from a loading station to a designated station, the flow of robots would be uniformly distributed across each link, maximizing network capacity.
-However, under such circumstances, it becomes unfeasible for a single loading station to serve multiple outlets—a common requirement in sortation scenarios, especially those without pre-sorting.
-Therefore, we depict the network capacity reduction mechanism by analyzing the characteristics of RC-S, and investigate how the number of operational loading stations and the total number of robots affect the traffic.
+This section elaborates on the network capacity by analyzing the characteristics of RC-S and investigates how the system configurations affect system throughput.
+A throughput estimation formula is proposed, serving as a constraint for the planning model in section 7.
 Two important metrics in traffic analysis (Daganzo,, 2010) are the total vehicular distance traveled per second of operation $\bar{m}(veh\cdot m/s)$, and the average travel distance $\bar{l}$.
@@
 $\tilde{T_{O}}=\frac{\bar{m}}{\bar{l}}$ (11)
-Under the logic of RC-S, the first metric $\bar{m}$ is numerically consistent with the product of the number of occupied VPs, denoted by $n^{occupied}_{VP}$, and the speed of each VP, denoted by $v_{VP}$:
+Under the mechanism of RC-S, the first metric $\bar{m}$ is numerically consistent with the product of the number of occupied VPs, denoted by $n^{occupied}_{VP}$, and the speed of each VP, denoted by $v_{VP}$:
 $\bar{m}=n^{occupied}_{VP}\cdot v_{VP}$ (12)
-The speed of each VP could be obtained by expression (4)-(6) with the parameters of RC-S, while the number of occupied VPs is influenced by two factors: workforce factor $\kappa$, representing the impact of the number of workers on the quantity of VPs that can be occupied; attenuation factor $\beta$, representing the utilization of VPs by the feasible paths generated by RC-S.
+The speed of each VP could be obtained by expression (4)-(5) with the parameters of RC-S.
+For the number of occupied VPs, an intuitive idea is that it is related to the number of workers engaged in loading parcels in the loading zone and the spatial utilization under RC-S.
+We introduce two factors: workforce factor $\kappa$, representing the impact of the number of workers on the quantity of VPs that can be occupied; attenuation factor $\beta$, representing the utilization of VPs by the feasible paths generated by RC-S.
 In addition, the number of available robots $n_{r}$ also limits the upper bound of occupied VPs.
@@
 where $n_{VP}$ denotes the total number of VPs in the network, which is constant given the scale of the network.
-According to the phase analysis in section 4.3, the RC-S scheme stipulates that half of the aisle cells are occupied by VPs at any given time.
+According to the phase analysis in section 4.2, the RC-S scheme stipulates that the number of VPs is equal to the number of intersection nodes at any given time.
 Therefore, we obtain:
-$n_{VP}=\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2}$ (14)
+$n_{VP}=n_{h}\cdot n_{v}$ (14)
 Now we investigate the forms of the two factors.
 Workforce factor
-In the operational context, the relationship between a VP and its respective entrance mirrors the relationship between a conveyor belt and its driving electric motor.
-A VP becomes eligible for occupation only when its associated loading station, situated at the aisle’s entrance, is operational—indicative of the corresponding worker being actively engaged in tasks.
-The quantity of such active VPs is a critical determinant of the maximum capacity for robot throughput.
-Take, for instance, a scenario where 8 workers are stationed within an RSS featuring a $6+6$ aisle network.
-To ensure even distribution of traffic flow and mitigate congestion at the aisle network’s periphery, these workers are strategically positioned along the four edges, as depicted in Figure 9.
-For ease of discussion, the notation $\alpha$ will be employed to denote the ratio of active workers to the network’s maximum operational capacity in discussions that follow.
+This factor represents the impact of the number of workers on the quantity of VPs that can be occupied.
+For each aisle, the newly generated VPs at the entrance could be occupied by loaded robots when its corresponding loading station is active and worker is engaged in tasks.
+The quantity of such VPs is a critical determinant of the maximum throughput.
+Consider, for instance, a scenario where eight workers are stationed in an RSS with a $6+6$ aisle network.
+To ensure the balance of traffic flow and mitigate congestion at the aisle network’s periphery, workers are are evenly divided into four groups and positioned along the four edges of the site, as depicted in Figure 9.
+Four horizontal aisles and four vertical aisles are served by workers.
+We merge the area covered by these aisles and obtain the blue polygon.
+Figure 9: The impact of number of workers on the VPs that can be occupied
+For ease of discussion, the notation $\alpha$ will be employed to denote the ratio of workers to the network’s maximum worker capacity.
 $\alpha=\frac{n_{w}}{n_{h}+n_{v}}$ (15)
-Figure 9: The impact of number of workers on the VPs that can be occupied
-The top 4 horizontal aisles and left 4 vertical aisles are served by workers, and thus robots could be loaded with parcels at the corresponding entrances.
-We merge the areas covered by these aisles and obtain the blue polygon.
-The ratio of area, namely the workforce factor $\kappa$, could be obtained by:
+The proportion of the blue region, noted as the workforce factor $\kappa$, could be obtained by:
 $\kappa=1-{(1-\alpha)}^{2}$ (16)
+We propose the following proposition regarding the relationship between the workforce factor and the number of occupied VPs:
+PROPOSITION 1.
+If either (i) the candidate feasible paths in RC-S include no more than two turns, or (ii) the average travel distance of a task is less than $(2-\alpha)min\{n_{v},n_{h}\}$, then an upper bound of the proportion of occupied VPs is $\kappa$;
+The proof is presented in Online Appendix B. Proposition 1 provides with support for estimating the maximized network throughput. Next, we analyze the utilization rate of VPs using the RC-S scheme.
 Attenuation factor
-We now examine the path reservation strategy within the RC-S framework.
+This factor represents the utilization of VPs by the feasible paths generated by RC-S.
+We begin with examine the path reservation strategy within the framework of RC-S .
 In a first-come-first-served basis, the system allocates to each newly requested sorting task an available robot, assigning it the currently shortest viable route.
 This robot then proceeds through the network along a predefined spatio-temporal path, ensuring a collision- and queue-free delivery within the aisle network.
-However, this spatio-temporal reservation method exhibits a degree of ”inflexibility” because it obligates the robot to adhere strictly to its assigned spatio-temporal trajectory, precluding any alterations like deceleration or temporary halts.
-Such inflexibility can lead to underutilization of traffic capacity, as VPs that are momentarily unoccupied cannot always be sequentially linked to construct a continuous path if they do not intersect at the same point in time, thereby diminishing the network’s traffic capacity.
-To quantify the impact of this reduction, we introduce an attenuation factor $\beta$, representing the proportion of network throughput decrease relative to its potential maximum.
-Figure 10: The ”pile up” process of spatiotemporal occupation under RC.
-Figure 11: The loss of capacity in different network.
+However, this spatio-temporal reservation mechanism obligates the robot to adhere strictly to its assigned spatio-temporal trajectory, precluding any alterations like deceleration or temporary halts.
+Such inflexibility can lead to underutilization of traffic capacity, as VPs that are momentarily unoccupied cannot always be sequentially linked to construct a continuous path if they do not intersect at the same point in time.
+To quantify the impact of this reduction, we introduce an attenuation factor $\beta$, representing the spatial utilization.
+Figure 10: The ”pile up” process of spatiotemporal occupation under RC
+Figure 11: The loss of capacity in different network
 The attenuation factor $\beta$ rises with the increase of $(n_{h}+n_{v})$.
@@
 To this end, we derive the final expression of number of occupied VPs:
-$n^{occupied}_{VP}(\alpha,n_{r})=\min\{\frac{1-{(1-\alpha)}^{2}}{a+b(n_{h}+n_{v})}\cdot\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2},\ n_{r}\}$ (18)
+$n^{occupied}_{VP}(\alpha,n_{r})=\min\{\frac{1-{(1-\alpha)}^{2}}{a+b(n_{h}+n_{v})}\cdot n_{h}n_{v},\ n_{r}\}$ (18)
 LEMMA 1.
@@
 Average travel distance
-We now derive the average travel distance of a sorting task, as the remaining part of the throughput estimation formula.
+We now derive the average travel distance $\bar{l}$ of a sorting task, as the remaining part of the throughput estimation formula.
 Based on the mechanism of RC-S, an appropriate selection of paths is based on the principles of evenly distributed sorting demand, workload balancing of workers, and minimizing travel distance.
@@
 LEMMA 2.
-[…]
+For uniformly distributed sorting demands, if the length-to-width ratio is less than 2 and $min\{n_{h},n_{v}\}\geq 4$, (i) as $\alpha$ increases, $\bar{l}(n_{h},n_{v},\alpha)$ decreases initially and then increases; (ii) when the site size is fixed ($n_{h}\cdot n_{v}=\textit{constant}$), a square-shaped site has a smaller average travel distance; (iii) when the site is square-shaped, the range of average travel distance could be obtained by $2D\cdot n_{h}<\bar{l}(n_{h},n_{v},\alpha)<2D\cdot\frac{9}{8}n_{h}$.
 Proofs for Lemma 2 is provided in Appendix B.
@@
 $\frac{\partial\tilde{T}_{M}(\cdot)}{\partial n_{l}}$ has one zero point within the valid range of $n_{l}$.
-Denote the value as $n_{l}^{(0)}$.
-Then $\tilde{T}_{M}(\cdot)$ is monotonically increasing in the interval $(0,n_{l}^{(0)}]$ and monotonically decreasing in the interval $(n_{l}^{(0)},n_{h}+n_{v}]$.
+Denote the value as $n_{l}^{(0)}$, then $\tilde{T}_{M}(\cdot)$ is monotonically increasing in the interval $(0,n_{l}^{(0)}]$ and monotonically decreasing in the interval $(n_{l}^{(0)},n_{h}+n_{v}]$.
 The proof of Proposition 2 can be derived by incorporating Lemma 1 and Lemma 2, which are depicted in detail in Appendix B.
-Proposition 1 reveals that under the RC-S robot control scheme, the system’s maximum operational efficiency has a critical point, and this critical point is not achieved when all loading stations are open.
-This may seem counter-intuitive, but remember that we have assumptions of workload balancing and uniformly distributed sorting demands.
-When any short-sighted dispatch strategy is employed at the loading stations, robots dispatched from the corners will negatively impact the overall throughput of the entire system.
+Proposition 2 reveals that under the RC-S robot control scheme, the maximum throughput has a critical point, and this critical point is not achieved when all loading stations are activated, due to workload balancing and uniformly distributed demands.
+When the proportion of active loading stations approaches 1, the growth rate of $n^{occupied}_{VP}$ will be slower than the growth rate of the service path length, resulting in a decrease in throughput.
+We will verify the formula through experiments in the next section.
 6 Numerical Validation
@@
 In each experiment, we set a warm-up period of 10 minutes and then record the system performance for 50 minutes of continuous operations.
+Table 3: Results on 12+12 aisle network
+$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time (ms)
+RC-S CA* RC-S CA* RC-S CA*
+40 8.83 4.72 11.73 14.17 5.67 5.99
+80 15.12 9.03 12.92 14.57 6.80 16.85
+120 17.45 12.40 13.89 15.24 8.50 35.37
+160 18.06 14.32 14.26 15.79 9.31 55.31
+200 18.33 14.94 14.51 16.03 9.63 63.36
+Table 4: Results on 20+20 aisle network
+$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time (ms)
+RC-S CA* RC-S CA* RC-S CA*
+50 7.68 3.94 18.85 21.75 15.53 11.61
+100 14.91 7.79 19.33 21.95 15.79 29.94
+200 26.82 15.07 20.77 22.54 19.76 99.61
+300 32.32 22.58 21.48 23.33 29.27 226.76
+400 33.66 24.37 23.41 23.92 37.40 363.45
 Tables 3 and 4 report the system throughput, average service time and average run-time.
-The comparative results indicate that the runtime of RC-S is consistently lower than that of CA*, and in larger-scale cases, it is even less than $1/10$ of the latter.
+The comparative results indicate that the run-time of RC-S is consistently lower than that of CA*, and in larger-scale cases, it is even less than $1/10$ of the latter.
 Moreover, RC-S always assigns shorter service paths to robots, resulting in an average service time reduction of 10.3% compared to CA*.
@@
 This observation underscores the suitability of the RC-S algorithm for environments like RSS, characterized by a compact network and numerous outlets as obstacles.
-Table 3: Results on 12+12 aisles directed map
-$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time (ms)
-RC-S CA* RC-S CA* RC-S CA*
-40 8.81 4.72 11.70 14.17 5.26 5.99
-80 14.94 9.03 12.90 14.57 6.32 16.85
-120 16.99 12.40 13.81 15.24 7.52 35.37
-160 17.58 14.32 14.15 15.79 8.07 55.31
-200 17.81 14.94 14.32 16.03 8.35 63.36
-Table 4: Results on 20+20 aisles directed map
-$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time (ms)
-RC-S CA* RC-S CA* RC-S CA*
-50 7.41 3.94 19.34 21.75 14.38 11.61
-100 14.12 7.79 19.92 21.95 14.96 29.94
-200 21.04 15.07 20.99 22.54 19.32 99.61
-300 22.34 21.18 21.48 23.33 21.30 226.76
-400 22.91 24.37 21.97 23.92 22.08 363.45
+The difference in algorithm performance is primarily due to the following improvements: (1) RC-S optimizes by selecting from candidate spatio-temporal paths, while CA* and similar cell-based search algorithms consider variable waiting times at each cell and allow robots to return to previously visited cells; (2) RC-S limits the total number and maximum density of robots and is free of conflicts, while high robot density around outlets in benchmark algorithm leads to significant resource expenditure on conflict resolution.
 6.2 Throughput estimation formula validation
-This section validates the throughput estimation formula proposed in section 5.
+This section validates the throughput estimation formula proposed in section 5 across different active ratios of loading stations.
 We consider 5 scenarios with different network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively.
@@
 This indicates that Algorithm 1, in practical execution, cannot effectively balance the distribution of robots in the loading stations in the extreme situation.
-Consequently, it incurs a certain efficiency loss, resulting in a decrease in the throughput, as demonstrated in Proposition 1.
+6.3 Efficiency underestimating in queueing models
+Queueing models have been widely used in analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different work pieces (Buzacott and Shanthikumar,, 1993).
+In robotic systems, robots function similarly to pallets in traditional manufacturing systems, transferring work pieces from one machine to another for processing.
+However, we find that queueing models could not capture the efficiency of robots accurately when the number of robots has not reached the network’s capacity limit.
+We apply the CQN model established in Zou et al., (2021), where a two-tier RSS model with closest Robot-to-Loading-Station assignment rule is modified to meet the assumptions in section 3.
+Eight scenarios of large-scale network with few robots are tested, using the same experimental setup as in section 6.2.
+The results are shown in Table 5.
+The CQN model significantly underestimates system throughput in scenarios with few robots.
+It fails to capture variations in traffic congestion across different flow rates and overestimates road conflicts under low-demand conditions, thereby underestimating individual robot efficiency.
+Although the CQN model provides accurate estimates when the network approaches capacity, it results in unnecessary expenses for surplus robots during off-peak periods.
+Table 5: Error in throughput estimation
+$n_{h}+n_{v}$ $20+20$ $24+24$
+$n_{r}$ 50 100 200 300 50 100 200 300
+Throughput (×$\times 10^{3}$/h) Simulation 8.43 16.25 28.60 32.90 7.27 14.30 27.01 36.49
+CQN 5.98 11.87 22.98 31.47 4.96 9.88 19.52 28.61
+RC-S 7.83 15.66 31.32 33.86 6.55 13.10 26.19 39.08
+Relative error (%) CQN 29.03 26.95 19.65 4.34 31.82 30.92 27.73 21.61
+RC-S 7.12 3.65 -9.52 -2.92 9.92 8.41 3.04 -7.10
 7 Layout Design Optimization
-The layout design problem aims at seeking for a balance between facility costs and long-term operating costs. The facility costs $C_{f}$ and operations costs $C_{o}$ are calculated as follows.
+The layout design problem aims at seeking for a balance between facility costs and long-term operating costs.
+A larger facility will achieve higher maximum throughput; however, it will also increase construction and equipment costs.
+Considering transportation costs and supply stability, we analyze the following scenario: The warehouse signs a long-term RaaS contract with a robotics company, with each year divided into several operating periods denoted by $\sigma$.
+The proportion of each period’s duration within the year is represented by $\theta^{\sigma}$, and the planned throughput level during each period is $T^{\sigma}$.
+The facility costs $C_{f}$ and operations costs $C_{o}$ are calculated as follows.
 $C_{f}=P_{s}\cdot[2D\cdot(n_{h}-1)+W_{w}+W_{l}]\cdot[2D\cdot(n_{v}-1)+W_{w}+W_{l}]+P_{l}\cdot n_{l}$ (30)
-$C_{o}=P_{w}\cdot[\theta\cdot n_{w}^{H}+(1-\theta)\cdot n_{w}^{L}]+P_{r}\cdot[\theta\cdot n_{r}^{H}+(1-\theta)\cdot n_{r}^{L}]$ (31)
-Parameter $P_{s}$ and $P_{l}$ in equation (30) denote the discounted site rental cost per square meter and equipment cost per loading station, while $M_{w}$ and $M_{r}$ in equation (31) denote the discounted unit labor cost and unit rental cost of robot.
-We assume that there are two typical values of sorting demand, namely the average off-peak season demand ${T}^{L}$ and the average peak season demand ${T}^{H}$, with $\theta$ denoting the ratio of peak seasons.
-The objective of the problem is as follows:
-$C_{d}({T}^{L},{T}^{H},\theta,N_{o})=C_{f}+C_{o}$ (32)
+$C_{o}=\sum_{\sigma\in\mathcal{S}}\theta^{\sigma}(P_{w}n_{w}^{\sigma}+P_{r}n_{r}^{\sigma})$ (31)
+Parameter $P_{s}$ and $P_{l}$ in equation (30) denote the discounted site rental cost per square meter and equipment cost per loading station, while $P_{w}$ and $P_{r}$ in equation (31) denote the discounted unit labor cost and unit rental cost of robot.
+The objective of the problem is modeled as follows:
+$C_{d}(T^{\sigma},\theta^{\sigma},N_{o})=C_{f}+C_{o}$ (32)
 where parameter $N_{o}$ denotes the minimum number of outlets that should be covered.
@@
 We formulate the layout design problem (LDP) as following integer programming model.
-$\underset{n_{h},n_{v},n_{l},n_{w}^{H},n_{r}^{H},n_{w}^{L},n_{r}^{L}}{\min}\ C_{d}$
+$\underset{n_{h},n_{v},n_{l},n_{w}^{\sigma},n_{r}^{\sigma}}{\min}\ C_{d}$
 […]
-$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{H},n_{r}^{H})\geq{T}^{H}$ (33)
-$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{L},n_{r}^{L})\geq{T}^{L}$ (34)
+$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{\sigma},n_{r}^{\sigma})\geq T^{\sigma}\qquad\forall\sigma\in\mathcal{S}$ (33)
 $(n_{h}-1)(n_{v}-1)\geq N_{o}$ (34)
 $n_{l}\leq n_{h}+n_{v}$ (35)
-$n_{w}^{H}\leq n_{l}$ (37)
-$n_{w}^{L}\leq n_{l}$ (38)
-$n_{h}=2k_{h}\quad\forall k_{h}\in\mathcal{Z}_{+}$ (39)
-$n_{v}=2k_{v}\quad\forall k_{v}\in\mathcal{Z}_{+}$ (40)
-$n_{h},n_{v},n_{w}^{H},n_{r}^{H},n_{w}^{L},n_{r}^{L}\in\mathcal{Z}_{+}^{(6)}$ (41)
-where constraints (33) and (34) ensure that the solution meets predicted sorting requirements.
+$n_{w}^{\sigma}\leq n_{l}\qquad\forall\sigma\in\mathcal{S}$ (36)
+$n_{h}=2k_{h}\ ,\quad k_{h}\in\mathcal{Z}_{+}$ (37)
+$n_{v}=2k_{v}\ ,\quad k_{v}\in\mathcal{Z}_{+}$ (38)
+$n_{w}^{\sigma},n_{r}^{\sigma}\in\mathcal{Z}_{+}^{2\left|\mathcal{S}\right|}$ (39)
+Constraint (33) ensures that the system throughput meets predicted sorting requirements in different periods.
 Constraint (34) restricts the lower bound of the site size to accommodate an adequate number of outlets.
 Constraint (35) represents the maximum number of loading stations accommodated in the aisle network under RC-S.
-Constraints (37) and (38) limit the maximum number of workers to be stationed.
-By the design of the system in section 3, we add constraints (39)-(40) to ensure the number of aisles even.
-The remaining constraints include the expressions to calculate the cost of each component and the estimation formula of system efficiency.
+Constraint (36) limits the maximum number of workers to be stationed.
+Constraints (37)-(38) ensure the number of aisles even.
+The remaining constraint (39) requires the variables to be positive integers.
 It is difficult to solve LDP because of the non-linearity of the inequality constraints, especially for large-scale problems.
@@
 To ensure the constraints continuous and differentiable, we introduce additional constraints requiring that the number of robots is less than the number of available VPs, shown as follows.
-$n_{r}^{H}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{H}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (42)
-$n_{r}^{L}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{L}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (43)
+$n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\quad\forall\sigma\in\mathcal{S}$ (40)
 At the end of the algorithm, the optimal solution of the relaxed problem is converted into a feasible solution for the original problem by rounding up.
@@
 The data regarding the equipment cost of loading stations and the rental cost of robots are obtained from Geekplus company.
-The ratio of peak month $\theta$ is set to 1/6, which means that every quarter has a two-week-long promotion campaign in average.
-Meanwhile, we set ${T}^{L}=0.8\cdot{T}^{H}$ to ensure that there is no enormous space waste during the off-peak period.
-The default parameter of experiments is shown in table 5 and the configurations of RC-S keeps the same with table 2.
+We assume that the warehouse only expands its sorting capacity during a few major shopping events, meaning there are two typical levels of sorting demand, namely the average off-peak season demand ${T}^{L}$ and the average peak season demand ${T}^{H}$, with ${T}^{L}=0.8\cdot{T}^{H}$.
+The ratios of the two periods are set as $(\theta^{L},\theta^{H})=(5/6,1/6)$, which means that each quarter typically includes a two-week-long shopping event.
+The default values of parameters is shown in table 6 and the configurations of RC-S keeps the same with table 2.
 Our algorithms are coded in Python 3.11 with the solver Gurobi Optimizer 9.0, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM.
 All processes are run single-threaded.
-Table 5: Default parameter in sensitivity analysis
+Table 6: Default values of parameters in section 8
 $M_{s}(CNY/m^{2}\cdot mo)$ $M_{l}(CNY/mo)$ $M_{w}(CNY/mo)$ $M_{r}(CNY/mo)$ $W_{w}(m)$ $W_{l}(m)$
@@
 We begin by investigating the impact of site rental cost under two different scenarios: one with $N_{o}=100$ (scenario 1) and another with $N_{o}=400$ (scenario 2).
-The monthly rental cost of site floor space per square meter varies from $10\ CHY$ to $30\ CHY$.
+The monthly rental cost of site floor space per square meter varies from $10\ CNY$ to $30\ CNY$.
 The results are presented in Figure 16(a) and Figure 16(b).
@@
   • Comparing the data for $T^{H}=15,000$ and $T^{H}=30,000$ in Table 7, the total costs for the latter are nearly three times those for the former.
-This indicates that due to the simultaneous growth of $C_{s}$ and $C_{o}$, the total costs growth rate of the RSS system exceeds the throughput growth rate.
+This indicates that due to the simultaneous growth of $C_{f}$ and $C_{o}$, the total costs growth rate of the RSS system exceeds the throughput growth rate.
 From the results of study in site rental cost, the derived managerial insights are summarized as follows.
@@
 Instead of running RSS at full capacity, it is more cost-effective to appropriately expand the site size and reduce the proportion of activated loading stations.
-8.2 Labor cost
-In a similar manner as our previous experiments, we now explore the influence of different labor costs on the total costs in two scenarios with $N_{o}=100$ and $N_{o}=400$.
-The monthly labor cost of each worker varies from $5,000\ CHY$ to $12,000\ CHY$.
-The proportions of labor costs and robot rental costs are both visualized in Figure 17(a) and Figure 17(b).
-Our observations are concluded as follows:
 (a) $N_{o}=100$ (scenario 1)
@@
 Figure 17: Comparison of total costs and labor costs proportion on different unit labor cost
+8.2 Labor cost
+In a similar manner as our previous experiments, we now explore the influence of different labor costs on the total costs in two scenarios with $N_{o}=100$ and $N_{o}=400$.
+The monthly labor cost of each worker varies from $5,000\ CNY$ to $12,000\ CNY$.
+The proportions of labor costs and robot rental costs are both visualized in Figure 17(a) and Figure 17(b).
+Our observations are concluded as follows:
   • As $T^{H}$ increases, the proportion of labor costs gradually rises and eventually stabilizes.
@@
 Today’s commonly used conveyor sorters, equipped with high-power motors, necessitate substantial initial investments but experience only slight increments in total costs as sorting demands grow (Zou et al.,, 2021).
-Furthermore, the optimal system configurations of a conveyor sorter are not easily affected by fluctuations in the unit price of cost components (Russell and Meller, 2003b, ), which makes it suitable for long-term operation in large-scale sorting scenarios.
+Furthermore, the optimal system configurations of a conveyor sorter are not easily affected by fluctuations in the unit price of cost components (Russell and Meller,, 2003), which makes it suitable for long-term operation in large-scale sorting scenarios.
 Therefore, we conclude the following insight regarding the application of RSS:
@@
 In the operations stage, efficient robot scheduling emerges as a crucial element for unlocking the business value of RSS.
-We propose an innovate RC-S scheme, which serves as an framework for managing a considerable number of robots simultaneously and efficiently.
+We propose an innovative RC-S scheme, which serves as a framework for managing a considerable number of robots simultaneously and efficiently.
 We provide a detailed description of the composition of phases and the method for setting parameters in RC-S.
@@
 Subsequently, a theoretical analysis of the efficiency of RC-S is conducted, exploring the impact of various system configurations on throughput.
-In the validation section, we first compare our proposed traffic management framework with other common framework based on A*.
-The simulation results indicate that our control method achieves a high level of performance and computational efficiency.
+In the validation section, we first compare our proposed traffic management framework with classical cooperative A* algorithm.
+The simulation results indicate that our control method achieves a higher level of performance and computational efficiency.
 Furthermore, we validate the throughput estimation formula.
-Results show that the analytical model fits the real traffic of robots well.
-In the site planning stage, warehouse managers need to make decisions that involve a trade-off between the initial site investment $C_{s}$, and the discounted expected operations costs $C_{o}$.
+Results show that queueing models can experience significant distortion in some scenarios, while our model fits the real traffic flows of robots more precisely.
+In the site planning stage, warehouse managers need to make decisions that involve a trade-off between the initial investment of facility $C_{f}$, and the discounted operations costs $C_{o}$.
 We analyze the cost composition of the RSS, and propose a layout design optimization model LDP that minimizes the total system costs.
-Through a detailed sensitivity analysis examining the cost dynamics of site rental and labor, our findings demonstrate a distinct cost distribution based on throughput levels.
+Through a detailed sensitivity analysis examining the cost dynamics of site rental and labor, we investigate cost proportions under distinct throughput levels.
 Specifically, at lower throughput, the rental cost for the site represents a considerable proportion of the total costs.
@@
 Secondly, extending the analysis to cover the entire process from order fulfillment to sorting in modern warehouses would provide valuable insights for enhancing system performance.
+Declaration of generative AI and AI-assisted technologies in the writing process
+During the preparation of this work the authors used ChatGPT in order to check for grammatical errors and improve readability.
+After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
 Appendix A: Main notations
@@
 RMFS robotic mobile fulfill system
+RaaS Robot-as-a-Service
 MAPF multi-agent path finding problem
@@
 Sets
-$\mathcal{G}$
-graph constituted of all nodes and links
-$\mathcal{V}_{c}$
-set of all conflict nodes
-$\mathcal{V}_{u}$
-set of all unloading nodes
-$\mathcal{A}$
-set of all links
-$\mathcal{S}$
-set of all loading stations
-$\mathcal{O}$
-set of all outlets
-$\mathcal{E}$
-set of all cycles
-$\mathcal{R}$
-set of all feasible path
-$\mathcal{R}(i,j,k)$
-set of feasible path connecting loading station $i$ and $j$, passing outlet $k$
-$\mathcal{R}_{i},k$
-set of feasible path starting at loading station $i$ and passing by outlet $k$
-^$\hat{\mathcal{R}}_{i}$
-set of feasible path ending at loading station $i$
-$\mathcal{I}_{x}$
-set of locations of loading stations on the top and bottom sides
-$\mathcal{I}_{y}$
-set of locations of loading stations on the left and right sides
+$\mathcal{G}$ graph constituted of all nodes and links
+$\mathcal{V}_{c}$ set of conflict nodes
+$\mathcal{V}_{u}$ set of unloading nodes
+$\mathcal{V}_{e}$ set of entrance/exit nodes
+$\mathcal{E}$ set of all edges
+$\mathcal{L}$ set of all loading stations
+$\mathcal{O}$ set of all outlets
+$\mathcal{C}$ set of all cycles
+$\mathcal{R}$ set of all feasible path
+$\mathcal{R}(i,j,k)$ set of feasible path connecting loading station $i$ and $j$, passing outlet $k$
+$\mathcal{R}_{i},k$ set of feasible path starting at loading station $i$ and passing by outlet $k$
+^$\hat{\mathcal{R}}_{i}$ set of feasible path ending at loading station $i$
+$\mathcal{I}_{x}$ set of locations of loading stations on the top and bottom sides
+$\mathcal{I}_{y}$ set of locations of loading stations on the left and right sides
+$\mathcal{S}$ set of operating periods
 Parameters
-$D$
-length of each grid in the sorting zone
-$W_{w}$
-width of waiting zone
-$W_{l}$
-width of loading zone
-$C_{f}$
-the facility costs of RSS
-$C_{o}$
-the operations cost of RSS
-$C_{d}$
-the total costs of RSS
-$P_{f}$
-cost of facility per square floor space
-$P_{l}$
-cost of equipment in one loading station
-$P_{w}$
-cost of labor per person per month
-$P_{r}$
-cost of robot per vehicle per month
-${T}_{H},{T}_{L}$
-target throughput level in peak/off-peak season
-$\theta$
-Ratio of peak seasons
-$N_{o}$
-minimum number of outlets required to meet the sorting category demands
-$\gamma$
-weight of operations costs in optimization model
-$\tau_{c}$
-time interval of VPs in RC-S
-$\tau_{e}$
-travel time of VPs on each link in RC-S
-$v_{VP}$
-speed of VPs
-$v_{max}$
-maximum speed of robots
-$c_{max}$
-maximum acceleration/deceleration of robots
-$r_{ls}$
-maximum loading rate of a loading station
-$c_{i,k}^{r}$
-travel time of feasible path $r\in R_{i,k}$
-$\hat{c}_{i,k}$
-penalty of delay of a sorting task from loading station $i$ to outlet $k$
-$\delta_{i,k}^{r,\nu,e}$
-incidence between feasible path $r\in R_{i,k}$, node $\nu$ and cycle $e$
-$N_{\nu}^{e}$
-remaining capacity in node $\nu$ in cycle $e$
-$d_{i,k}$
-sorting demand from loading station $i$ to outlet $k$
+$D$ length of each grid in the sorting zone
+$W_{w}$ width of waiting zone
+$W_{l}$ width of loading zone
+$C_{f}$ the facility costs of RSS
+$C_{o}$ the operations cost of RSS
+$C_{d}$ the total costs of RSS
+$P_{f}$ cost of facility per square floor space
+$P_{l}$ cost of equipment in one loading station
+$P_{w}$ cost of labor per person per month
+$P_{r}$ cost of robot per vehicle per month
+${T}^{\sigma}$ target throughput level in operating period $\sigma$
+$\theta^{\sigma}$ Ratio of operating period $\sigma$
+$N_{o}$ minimum number of outlets required to meet the sorting category demands
+$\gamma$ weight of operations costs in optimization model
+$\tau_{c}$ time interval of VPs in RC-S
+$\tau_{e}$ travel time of VPs on each link in RC-S
+$v_{VP}$ speed of VPs
+$v_{max}$ maximum speed of robots
+$c_{max}$ maximum acceleration/deceleration of robots
+$r_{l}$ maximum loading rate of a loading station
+$c_{i,k}^{r}$ travel time of feasible path $r\in R_{i,k}$
+$\hat{c}_{i,k}$ penalty of delay of a sorting task from loading station $i$ to outlet $k$
+$\delta_{i,k}^{r,\nu,l}$ incidence between feasible path $r\in R_{i,k}$, node $\nu$ and cycle $l$
+$N_{\nu}^{l}$ remaining capacity in node $\nu$ in cycle $l$
+$d_{i,k}$ sorting demand from loading station $i$ to outlet $k$
 Variables
-$x_{i,k}^{r}$
-decision on whether feasible path $r\in R_{i,k}$ is reserved in current cycle
-$\hat{x}_{i,k}$
-decision on whether the robot located at loading station $i$ should wait until the next cycle, with the target outlet $k$
-$n_{h},n_{v}$
-number of horizontal and vertical aisles in sorting zone
-$n_{l}$
-number of loading stations in loading zone
-$n_{w}^{H},n_{w}^{L}$
-number of workers in peak/off-peak season
-$n_{r}^{H},n_{r}^{L}$
-number of robots in peak/off-peak season
+$x_{i,k}^{r}$ decision on whether feasible path $r\in R_{i,k}$ is reserved in current cycle
+$\hat{x}_{i,k}$ decision on whether the robot located at loading station $i$ should wait until the next cycle, with the target outlet $k$
+$n_{h},n_{v}$ number of horizontal and vertical aisles in sorting zone
+$n_{l}$ number of loading stations in loading zone
+$n_{w}^{\sigma}$ number of workers in operating period $\sigma$
+$n_{r}^{\sigma}$ number of robots in operating period $\sigma$
 Appendix B: Proofs
-B.1 Derivation of Maximum Travel Distance in Constraint 5
+B.1 Proof of Proposition 1
+(i):
+By definition 1, a feasible path of RC-S connects two active loading stations.
+If a feasible path has no more than two turns, all its turning points are within the blue region; otherwise, its endpoint will fall into a non-active loading station.
+The movement trajectories of occupied VPs will not cross the boundaries of the blue region; therefore, their upper limit is $\kappa$.
+Condition (i) is sufficient.
+(ii):
+We derive the average road length that each loading station can be allocated, noted as $\tilde{d}_{ls}$:
+$\tilde{d}_{ls}(\alpha)$ $=$ $\frac{\big{(}1-{(1-\alpha)}^{2}\big{)}\cdot n_{h}n_{v}}{\frac{\alpha(n_{h}+n_{v})}{2}}$
+$=$ $\frac{(2-\alpha)\cdot 2n_{h}n_{v}}{n_{h}+n_{v}}\geq(2-\alpha)\cdot min\{n_{h},n_{v}\}$
+When the average travel distance of a task is less than $\tilde{d}_{ls}$, the proportion of the union of segments passed by the occupied VPs in the total set of segments is less than $\kappa$.
+Condition (ii) is sufficient.
+$\square$
+B.2 Derivation of Maximum Travel Distance in Constraint 5
 We first examine the acceleration process of the robot after a turn.
@@
 $\square$
-B.2 Derivation of Average Travel Distance in Section 5
+B.3 Derivation of Average Travel Distance in Section 5
 Let $I_{x}$ represents the set of x-coordinates of loading stations on the top and bottom sides, while $I_{y}$ represents the set of y-coordinates of loading stations on the left and right sides.
@@
 $=$ $2D\cdot(\frac{n_{h}+n_{v}}{2}+\frac{(1+\alpha)}{4}\cdot\frac{n_{h}n_{v}}{n_{h}+n_{v}})$
-B.3 Proof of Lemma 2
+B.4 Proof of Lemma 2
 We first prove (i). By equation (27), the expression of average travel distance of a sorting task is as follows:
@@
 $\square$
-B.4 Proof of Proposition 1
+B.5 Proof of Proposition 2
 We first obtain the expression of $\tilde{T}_{M}(n_{h},n_{v},n_{l})$:
-$\tilde{T}_{M}(n_{h},n_{v},n_{l})=\frac{D}{\tau_{e}}\cdot\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2[a+b(n_{h}+n_{v})]}\cdot\frac{\frac{2n_{l}}{n_{h}+n_{v}}-{(\frac{n_{l}}{n_{h}+n_{v}})}^{2}}{\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$
+$\tilde{T}_{M}(n_{h},n_{v},n_{l})=\frac{D}{\tau_{e}}\cdot\frac{n_{h}n_{v}}{a+b(n_{h}+n_{v})}\cdot\frac{\frac{2n_{l}}{n_{h}+n_{v}}-{(\frac{n_{l}}{n_{h}+n_{v}})}^{2}}{\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$
 We further denote:
@@
 The expression of the derivative of $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ can be written as:
-$\frac{\partial\tilde{T}_{M}(n_{h},n_{v},n_{l})}{\partial}=\frac{D}{\tau_{e}}\cdot\frac{n_{h}(n_{v}-1)+n_{v}(n_{h}-1)}{2[a+b(n_{h}+n_{v})]}\cdot\frac{1}{n_{h}+n_{v}}\cdot\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}$
+$\frac{\partial\tilde{T}_{M}(n_{h},n_{v},n_{l})}{\partial n_{l}}=\frac{D}{\tau_{e}}\cdot\frac{n_{h}n_{v}}{a+b(n_{h}+n_{v})}\cdot\frac{1}{n_{h}+n_{v}}\cdot\frac{\partial G_{n_{h},n_{v}}(\alpha)}{\partial\alpha}$
 It is immediate to prove that $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ has one zero point within the range $n_{l}\in(0,n_{h}+n_{v}]$ and it initially increases then decreases.
@@
 […]
-$s.t.\quad$ $y_{i}\geq g_{i}(\omega_{k})+\nabla{g_{i}(\omega_{k})}^{T}d,\quad i=1,2,\cdots,6$ (C-1)
+$s.t.\quad$ $y_{i}\geq g_{i}(\omega_{k})+\nabla{g_{i}(\omega_{k})}^{T}d,\quad i=1,2,\cdots,2\left|S\right|+2$ (C-1)
 $-\Delta_{k}\leq d\leq\Delta_{k}$ (C-2)
-$y_{i}\geq 0,\quad i=1,2,\cdots,6$ (C-3)
-where $\omega=\{n_{h},n_{v},n_{w}^{H},n_{r}^{H},n_{w}^{L},n_{r}^{L}\}$ is the set of all decision variables.
+$y_{i}\geq 0,\quad i=1,2,\cdots,2\left|S\right|+2$ (C-3)
+where $\omega=\{n_{h},n_{v},n_{w}^{\sigma},n_{r}^{\sigma}\}$ is the set of all decision variables.
 $\Delta_{k}$ is the bound of $d$ in iteration $k$.
-If we impose the constraint that the number of robots is less than the available VPs (considering the parking demand of robots during downtime, this setting is reasonable), constraints in layout design optimization model LDP are replaced with six inequalities in the LP-S:
-$g_{1}(\omega)=n_{r}^{H}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{H}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (C-4)
-$g_{2}(\omega)=n_{r}^{L}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{L}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0$ (C-5)
-$g_{3}(\omega)=-\frac{D\cdot n_{r}^{H}}{\tau_{e}\cdot E[l(n_{h},n_{v},\frac{n_{w}^{H}}{n_{h}+n_{v}})]}+{T}^{H}\leq 0$ (C-6)
-$g_{4}(\omega)=-\frac{D\cdot n_{r}^{L}}{\tau_{e}\cdot E[l(n_{h},n_{v},\frac{n_{w}^{L}}{n_{h}+n_{v}})]}+{T}^{L}\leq 0$ (C-7)
-$g_{5}(\omega)=-(n_{h}-1)(n_{v}-1)+N_{o}\leq 0$ (C-8)
-$g_{6}(\omega)=n_{w}^{H}-n_{h}-n_{v}\leq 0$ (C-9)
-To avoid waste of resources, the number of loading station is always equal to the peak number of workers, thus $n_{l}$ is replaced by $n_{w}^{H}$ in the constraints.
-The details of solving PLSP is shown in Algorithm C-1.
-It could be roughly divided into 2 steps in each iteration: (1) Generate LP-S and obtain the step size $d_{k}$, determine whether to stop; (2) Adjust step bounds for the next iteration.
+$g_{i}(\omega_{k})$ represents the left-hand side of standard form inequality constraints (C-4)-(C-7) derived from the original LDP, assuming that the number of robots is less than the available VPs (considering the parking demand of robots during downtime, this setting is reasonable):
+$-(n_{h}-1)(n_{v}-1)+N_{o}\leq 0$ (C-4)
+[…]
+$n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\qquad\forall\sigma\in\mathcal{S}$ (C-6)
+$-\frac{D\cdot n_{r}^{\sigma}}{\tau_{e}\cdot E[l(n_{h},n_{v},\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})]}+{T}^{\sigma}\leq 0\qquad\forall\sigma\in\mathcal{S}$ (C-7)
+[…]
 Algorithm C-1 The PLSP algorithm to solve model LDP
@@
 In iteration $k$, the termination of algorithm could be determined by calculating the intermediate variables as follows:
-$C_{E}(\omega_{k})=C_{d}(\omega_{k})+\mu\cdot\sum_{i=1}^{6}max\{0,g_{i}(\omega_{k})\}$ (C-10)
-$C_{EL_{k}}(\omega_{k})=C_{d}(\omega_{k})+\nabla{C_{d}(\omega_{k})}^{T}\cdot d_{k}+\mu\cdot\sum_{i=1}^{6}max\{0,g_{i}(\omega_{k})+{\nabla g_{i}(\omega_{k})}^{T}\cdot d_{k}\}$ (C-11)
-$R_{k}=\frac{C_{E}(\omega_{k})-C_{E}(\omega_{k}+d_{k})}{C_{EL_{k}}(\omega_{k})-C_{EL_{k}}(\omega_{k}+d_{k})}$ (C-12)
+$C_{E}(\omega_{k})=C_{d}(\omega_{k})+\mu\sum_{i}max\{0,g_{i}(\omega_{k})\}$ (C-8)
+$C_{EL_{k}}(\omega_{k})=C_{d}(\omega_{k})+\nabla{C_{d}(\omega_{k})}^{T}d_{k}+\mu\sum_{i}max\{0,g_{i}(\omega_{k})+{\nabla g_{i}(\omega_{k})}^{T}\cdot d_{k}\}$ (C-9)
+$R_{k}=\frac{C_{E}(\omega_{k})-C_{E}(\omega_{k}+d_{k})}{C_{EL_{k}}(\omega_{k})-C_{EL_{k}}(\omega_{k}+d_{k})}$ (C-10)
 where $\mu$ is a large enough constant.
@@
 Additionally, due to the relatively limited solution space, the number of iterations in experiments remained below 50, and the total solving time was in the order of seconds.
-Therefore, specific details are not explicitly described in the main text.
+To ensure the feasibility of the solution for the original problem, the numerical update process includes a rounding-up step.

### v3

@@
 Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in RSS, particularly when dealing with a substantial parcel flow.
-However, existing researches often overlook issues related to robot coordination, making the idealized assumption that there are no conflicts or deadlocks.
-Additionally, most of these studies tend to assume a fixed route travel time, excluding potential queuing delays on the road.
-As a result, their outcomes may become distorted, as the actual travel time of the robots deviates from the free-flow travel time.
+Zou et al., (2021) proposed several closed queuing network (CQN) models to quantitatively analyze the impact of congestion and designed an algorithm to estimate throughput.
+Validated through numerical experiments, these CQN models accurately reflect the performance of a real case from Deppon Express.
+However, the models primarily focus on the process of agents transferring between queues rather than on traffic flow and do not fully consider traffic management methods to mitigate conflicts.
+This aspect makes them less suited for analyzing the efficiency of MAPF solvers, such as the method proposed in this paper.
+On the other hand, other existing studies often overlooked issues related to robot coordination, relying on the idealized assumption that there are no conflicts or deadlocks.
+Consequently, their results may become distorted as the actual travel time of robots deviates from the free-flow travel time.
 Furthermore, these studies do not adequately explore the flexibility of RSS, particularly the ability to adjust the number of robots.
 Given the existing gap, this paper aims to address critical challenges in designing an efficient RSS, focusing on three key areas.
-First, we seek to ensure the system operates both stably and efficiently in a setting where multiple robots collaborate, a non-trivial challenge given the complexities of robot interaction.
-Second, the development of a precise model to quantify system efficiency within the RSS is essential.
-Such a model will enable accurate throughput estimation, which is vital for understanding the relationship between the size of the facility and its sorting capacity.
-Third, we emphasize the need for management strategies that leverage the flexibility of RSS to achieve cost-effective responses to fluctuating demands.
-These targeted areas of improvement aim to enhance the overall performance and scalability of RSS implementations.
+First, we propose a novel robot traffic management method to enhance system throughput, ensuring stable and efficient operations in multi-robot collaboration settings while overcoming the complexities of robot interactions.
+Second, we develop a precise model to quantify system efficiency in RSS, enabling accurate throughput estimation to clarify the relationship between system configurations and sorting capacity.
+Third, we highlight the need for management strategies that leverage RSS flexibility to provide cost-effective responses to fluctuating demands.
+These contributions aim to improve the performance and scalability of RSS implementations.
 To address the aforementioned challenges, we first explore the potential of centralized control strategy for sorting robots.
@@
 Figure 6: Concept of Rhythmic Control
-RC is essentially a traffic control method, similar to a combination of globally coordinated traffic signal control and road network perimeter control.
+RC is essentially a traffic control method, similar to a combination of globally coordinated traffic signal control and road network flow control.
 Within this framework, it is essential to clarify the movement patterns of each entity for integration into a sorting-specific robot management system.
@@
 In this section, we compare our proposed RC-S against cooperative A* (CA*) (Silver,, 2021) for traffic management in a warehouse, to better illustrate the superiority of RC-S in terms of AGV service distance and system throughput.
-CA* is based on a simple prioritized-planning scheme: each agent is assigned a unique priority, and based on the order of priorities, algorithm find the shortest path for each agent that avoids conflicts with agents of higher priority.
+CA* is a framework based on a simple prioritized-planning scheme: each agent is first assigned a unique priority, and based on the order of priorities, algorithm find the shortest path for each agent that avoids conflicts with agents of higher priority.
 We use SIPP (Phillips and Likhachev,, 2011), an efficient variant of location-time A*, as the lower-level solver of CA*.
@@
 Table 2: RC-S parameters in simulation
-$D(m)$ $\tau_{e}(s)$ $\tau_{c}(s)$ $v_{p}(m/s)$
+$D(m)$ $\tau_{e}(s)$ $\tau_{c}(s)$ $v_{VP}(m/s)$
 1 0.5 2 2
@@
 The simulations do not consider acceleration and deceleration phases.
-We conduct 10 rounds of replication for each test.
-In each experiment, we set a warm-up period of 10 minutes and then record the system performance for 50 minutes of continuous operations.
+We conduct 50 repeated experiments for each scenario.
+In each experiment, we set a warm-up period of 30 minutes and then record the system performance for 60 minutes of continuous operations.
+Our algorithms and simulations are coded in Python 3.11, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM.
+All processes are run single-threaded.
 Table 3: Results on 12+12 aisle network
-$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time (ms)
+$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time per cycle(ms)
 RC-S CA* RC-S CA* RC-S CA*
@@
 Table 4: Results on 20+20 aisle network
-$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time (ms)
+$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time per cycle (ms)
 RC-S CA* RC-S CA* RC-S CA*
@@
 Other experimental parameters are specified in section 6.1 in detail.
-We conduct 20 rounds of replication for each scenario.
-In each experiment, we set a warm-up period of one hour and then record the system performance for three hours of continuous operation.
 Figure 14: Analytical formula compared with simulation results.
@@
 6.3 Efficiency underestimating in queueing models
-Queueing models have been widely used in analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different work pieces (Buzacott and Shanthikumar,, 1993).
+Queueing models have been widely used in analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different work pieces (John A.
+Buzacott,, 1993).
 In robotic systems, robots function similarly to pallets in traditional manufacturing systems, transferring work pieces from one machine to another for processing.
@@
 Eight scenarios of large-scale network with few robots are tested, using the same experimental setup as in section 6.2.
-The results are shown in Table 5.
+The results are shown in Table 5, where Eq.(28) represents the proposed estimation formula.
 The CQN model significantly underestimates system throughput in scenarios with few robots.
-It fails to capture variations in traffic congestion across different flow rates and overestimates road conflicts under low-demand conditions, thereby underestimating individual robot efficiency.
+It lacks the capability to accurately depict traffic capacity under low flow conditions with RC-S control and overestimates the impact of conflicts under low-demand conditions, thereby underestimating individual robot efficiency.
 Although the CQN model provides accurate estimates when the network approaches capacity, it results in unnecessary expenses for surplus robots during off-peak periods.
-Table 5: Error in throughput estimation
+Table 5: Error in throughput estimation under RC-S control
 $n_{h}+n_{v}$ $20+20$ $24+24$
@@
 CQN 5.98 11.87 22.98 31.47 4.96 9.88 19.52 28.61
-RC-S 7.83 15.66 31.32 33.86 6.55 13.10 26.19 39.08
+Eq.(28) 7.83 15.66 31.32 33.86 6.55 13.10 26.19 39.08
 Relative error (%) CQN 29.03 26.95 19.65 4.34 31.82 30.92 27.73 21.61
-RC-S 7.12 3.65 -9.52 -2.92 9.92 8.41 3.04 -7.10
+Eq.(28) 7.12 3.65 -9.52 -2.92 9.92 8.41 3.04 -7.10
+we also conduct a experiments on a real case of Deppon Express described in Zou et al., (2021), where the network scale is $n_{h}+n_{v}=18+6$, with 108 outlets, 6 loading stations and 170 robots.
+The The aisle width $D$, maximum robot speed $v_{max}$, and worker loading rate $r_{l}$ were set according to the data provided by Deppon Express.
+We compared the performance of RC-S with the results from Zou et al., (2021), as shown in Table 6, where $T_{D}$, $T_{CQN}$ and $T_{RCS}$ represent the throughput of the Deppon system, the CQN model and the simulation under RC-S scheme, respectively.
+Combined with the experimental results in section 6.1, RC-S scheme significantly improves the operational efficiency of robots, thereby enhancing the overall system throughput.
+Meanwhile, the CQN model effectively captures the congestion in the RSS system implemented by Deppon Express, but underestimates the traffic capacity of the network when the RC-S scheme is applied.
+Table 6: Real Case Validation
+$D(m)$ $r_{w}(/s)$ $v_{max}(m/s)$ $\tau_{c}(s)$ $v_{VP}(m/s)$ $T_{D}(/h)$ $T_{CQN}(/h)$ $T_{RCS}(/h)$
+0.6 0.42 2 1.2 2 7,163 6,907 8,758
+6.4 Evaluating RC-S across diverse scenarios
+In real sorting centers, due to cost constraints and demand arrival rates, loading stations are often not spread across the entire perimeter of the RSS, resulting in only partial availability of network entrances and exits.
+The placement of outlets is also restricted by site factors, such as load-bearing columns or the presence of certain equipment.
+Additionally, demand arrival rates are heterogeneous and are not uniformly distributed across each outlet.
+In this section, we evaluate the performance of RC-S across various application scenarios and test the impact of different system configurations on efficiency by simulation.
+For each set of experiments, we evaluate the overall performance across 50 test runs.
+Each run includes a 30-minute warm-up period, followed by an observation period where average data is recorded over one hour.
+We first compared the impact of different numbers of active loading stations on average throughput.
+We consider five scenarios with varying network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively.
+The number of robots is set to be sufficient in each scenario, ensuring that it does not become a bottleneck for sorting process.
+Other RC-S parameters are specified in detail in Table 2.
+Figure 16 shows the trend of throughput across the five scenarios.
+It can be observed that large-scale networks not only have a higher throughput upper bound, but their slope—representing the average efficiency per loading station—is also generally higher.
+This is because, under RC-S, larger networks provide more space to distribute incoming traffic across different aisles, thereby reducing congestion at each entrance.
+Figure 16: The effect of different numbers of active loading stations
+(a) Division of outlet areas in the network
+(b) RC-S performance in different scenarios
+Figure 17: The effect of closing outlets in the network
+We next evaluate the effect of closing some outlets on system throughput.
+The network size is set to $n_{h}=n_{v}=20$, with other experimental parameters remaining constant.
+We consider four scenarios for closing outlets: (1) randomly closing outlets; (2) closing outlets in the corners; (3) closing outlets along the edges; (4) closing outlets in the center.
+To control variates, the number of closed outlets in each scenario is set to 1/9 of the total.
+Scenarios (2), (3), and (4) correspond to areas 1, 2, and 3 in Figure 17(a), respectively.
+Results from 50 repeated experiments for each scenario are shown in Figure 17(b).
+The green dashed line represents the average throughput when all outlets are open.
+It can be observed that randomly closing outlets has no significant impact on system throughput.
+Closing outlets in areas 1 and 2 results in a decrease in average throughput, while closing outlets in area 3 leads to a slight increase in average throughput.
+It is intuitive that closing outlets in a specific area causes an uneven distribution of demand, leading to imbalanced network traffic and increased traffic load in other areas of the network.
+However, from another perspective, it reduces robot turns within the closed area, thereby improving the utilization of spatio-temporal resources.
+Since most shortest paths pass through the central area, closing central outlets in scenario 4 results in higher throughput compared to the normal case.
+Figure 18: The effect of heterogeneous demand distribution
+In addition, we evaluate the effect of heterogeneous demand distribution on system efficiency.
+In real sorting centers, sorting destinations with higher demand arrival rates are often assigned multiple outlets to minimize differences in the visit frequency for each outlet; however, this cannot entirely eliminate demand heterogeneity.
+We assume there are two types of outlets in the RSS: one type has an average demand arrival rate twice that of the other, with both following a normal distribution.
+For outlets with extremely low arrival rates, the situation is similar to the previous experiments involving closed outlets, and their impact can be minimized by random assignment.
+The two types of outlets are spatially uniformly distributed throughout the network.
+Keeping other experimental parameters constant, we adjust the proportion of the two types of outlets, and the system performance are shown in Figure 18.
+The results indicate that with a fixed total demand and varying proportions of high-demand outlets, RC-S maintains efficiency, and the impact of demand heterogeneity on system throughput is negligible.
+This is achieved by leveraging demand arrival rates for different destinations and balancing the visit frequency to each area, ensuring a well-distributed traffic flow across the network.
+Table 7: RC-S performance in large-scale system
+$n_{h},n_{v}$ Throughput Avg. path Avg. time Max. time # Feasible paths
+($\cdot{10}^{4}$/h) length (m) per cycle (s) per cycle (s)
+22 3.871 27.832 0.0359 0.0942 118,136
+24 4.331 30.062 0.0480 0.1255 167,084
+26 4.699 32.517 0.0614 0.2393 229,888
+28 5.156 34.734 0.0836 0.2436 308,948
+30 5.517 37.261 0.1004 0.2820 406,856
+[…]
 7 Layout Design Optimization
@@
 The details of PSLP algorithm are shown in appendix C.
-8 Numerical Examples
+8 Numerical Examples for Layout Design
 In this section, we conduct a sensitivity analysis on the site rental cost per square meter $P_{s}$, and labor cost per man-month $P_{w}$, focusing on the trend of total costs under different scenarios as the sorting throughput level changes.
@@
 10 400 5,000 200 5 5
-8.1 Site rental cost
+8.1 Sensitivity Analysis on the site rental cost
 We begin by investigating the impact of site rental cost under two different scenarios: one with $N_{o}=100$ (scenario 1) and another with $N_{o}=400$ (scenario 2).
@@
 Figure 20: Comparison of total costs and labor costs proportion on different unit labor cost
-8.2 Labor cost
+8.2 Sensitivity Analysis on the labor cost
 In a similar manner as our previous experiments, we now explore the influence of different labor costs on the total costs in two scenarios with $N_{o}=100$ and $N_{o}=400$.

### v4

@@
 Previous research on RSS efficiency often assumed an ideal robot management system, ignoring potential traffic delays and assuming constant travel times.
-We introduce a novel robot traffic management method, named Rhythmic Control for Sorting Scenario (RC-S), for RSS operations, along with an estimation formula that establishes the quantitative relationship between system performance and configurations.
+To address this gap, we introduce a novel robot traffic management method, named Rhythmic Control for the Sorting Scenario (RC-S), for RSS operations, along with an analytical estimation formula that establishes the quantitative relationship between system performance and configurations.
 Simulations validate that RC-S reduces average service time by 10.3% compared to the classical cooperative A* algorithm, while also improving throughput and runtime.
@@
 Numerical studies show that facility costs dominate at lower throughput levels, while labor costs prevail at higher throughput levels.
-Additionally, due to traffic efficiency limitations, RSS is well-suited for small-scale operations like end-of-supply-chain distribution centers.
+Additionally, due to traffic efficiency limitations, an RSS is well-suited for small-scale operations like end-of-supply-chain distribution centers.
 Keywords: Logistics; Robotic sorting system; Robot management system; Performance evaluation; Layout design
@@
 Improved logistics and delivery services have fueled the rapid growth of e-commerce in the 21st century.
-The development of stable supply chains, same-day or next-day delivery options, and hassle-free return policies increase consumer confidence in online shopping (BusinessWire,, 2023).
-In order to handle the increasing volume of online orders, a new generation of warehouses specifically catering to individual customers has become a hot topic for logistics companies.
-This kind of warehouse efficiently meets the demand for small orders with tight delivery schedules through the implementation of automated equipments (Boysen et al., 2019b, ).
-For instance, robotic mobile fulfill system (RMFS), which employs a rack-moving mechanism, is widely used in intra-warehouse logistics, such as Amazon’s KIVA systems (Wurman et al.,, 2008).
-The main activities performed in warehouse include: (1) receiving, (2) transfer and put away, (3) order picking/selection, (4) accumulation/sorting, (5) cross-docking, and (6) shipping (De Koster et al.,, 2007).
+The development of stable supply chains, same-day or next-day delivery options, and hassle-free return policies increases consumer confidence in online shopping (ReginaLoBiondo).
+To handle the increasing volume of online orders, a new generation of warehouses specifically catering to individual customers has become a hot topic for logistics companies.
+This type of warehouse efficiently meets the demand for small orders with tight delivery schedules through the implementation of automated equipment (BOYSEN2019396).
+For instance, the robotic mobile fulfillment system (RMFS), which employs a rack-moving mechanism, is widely used in intra-warehouse logistics, such as Amazon’s KIVA systems (wurman2008coordinating).
+The main activities performed in a warehouse include: (1) receiving, (2) transfer and put-away, (3) order picking/selection, (4) accumulation/sorting, (5) cross-docking, and (6) shipping (de2007design).
 In this paper, we focus on the sorting process.
 Sorting involves categorizing and consolidating parcels according to their order information and shipping destinations.
-Conventional sorting systems commonly employ conveyor-based sorters , where the actuators move along with the conveyor belt, sequentially passing through each outbound station and releasing the loaded parcels at the appropriate locations (Boysen et al., 2019a, ).
-These systems are highly appreciated for its efficiency and stability, while suffering from the inflexibility and significant space occupation (Boysen et al.,, 2023).
+Conventional sorting systems commonly employ conveyor-based sorters , where the actuators move along with the conveyor belt, sequentially passing through each outbound station and releasing the loaded parcels at the appropriate locations (BOYSEN2019796).
+These systems are highly appreciated for their efficiency and stability, while suffering from the inflexibility and significant space occupation (BOYSEN2023582).
 (a) A dual-layer RSS solution
@@
 Figure 1: The application of RSS (source: www.geekplus.com)
-A novel sorting system utilizing mobile robots attracts increasing attention in recent years, which is called the robotic sorting system (RSS) (Zou et al.,, 2021).
+A novel sorting system utilizing mobile robots attracts increasing attention in recent years, namely the robotic sorting system (RSS) (zou2021robotic).
 Figure 1(a) presents a dual-layer solution of RSS.
-The site is arranged with numerous delivery ports, each associated with a specific sorting category, gathering all the parcels of that category.
+The site is arranged with numerous delivery ports, each associated with a specific sorting category, gathering all parcels of that category.
 A delivery port can be a bin, a cage cart, or a chute in a dual-layer structure.
@@
   1. 1.
-    The robot receives a parcel at a loading station, along with the destination information;
+    The robot receives a parcel at a loading station, along with destination information;
   2. 2.
@@
   3. 3.
-    When reaching the target outlet, the robot releases the loaded parcel;
+    Upon reaching the target outlet, the robot releases the loaded parcel;
   4. 4.
     The robot returns to the waiting zone behind a loading station and becomes idle.
-E-commerce order fluctuate during special selling seasons; therefore, warehouse throughput must be designed to accommodate these variations.
-Unlike conveyor-based sorting systems, the independent and modular nature of robotic operations contributes to the flexibility and scalability of RSS, making it well-suited to dynamic demands of e-commerce (Azadeh et al.,, 2017).
-During off-hours, warehouse managers can easily configure the control software to change the status of loading stations, the correspondences between outlets and destinations, and the aisle network topology (Xu et al.,, 2022).
-To leverage the flexibility, robotics companies introduce the innovative Robots-as-a-Service (RaaS) business model, allowing logistics clients to rent robots as needed.
+E-commerce orders fluctuate during special selling seasons; therefore, warehouse throughput must be designed to accommodate these variations.
+Unlike conveyor-based sorting systems, the independent and modular nature of robotic operations contributes to the flexibility and scalability of RSS, making it well-suited to the dynamic demands of e-commerce (pub99983).
+During off-hours, warehouse managers can easily configure the control software to change the status of loading stations, the correspondences between outlets and destinations, and the aisle network topology (XU2022102808).
+To leverage this flexibility, robotics companies have introduced the innovative Robots-as-a-Service (RaaS) business model, allowing logistics clients to rent robots as needed.
 RaaS eliminates the high costs associated with purchasing and maintaining robots and peripheral equipment, enabling warehouse managers to adjust the number of rented robots based on demand fluctuations.
@@
 The problem of finding conflict-free trajectories for all robots is known as the multi-agent path finding (MAPF) problem.
-The MAPF problem has been proven to be NP-hard (Yu and LaValle,, 2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in the instability of RSS efficiency.
-Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in RSS, particularly when dealing with a substantial parcel flow.
-Zou et al., (2021) proposed several closed queuing network (CQN) models to quantitatively analyze the impact of congestion and designed an algorithm to estimate throughput.
+The MAPF problem has been proven to be NP-hard (Yu_LaValle_2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in instability of RSS efficiency.
+poms13626 analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in RSS, particularly when dealing with a substantial parcel flow.
+zou2021robotic proposed several closed queueing network (CQN) models to quantitatively analyze the impact of congestion and designed an algorithm to estimate throughput.
 Validated through numerical experiments, these CQN models accurately reflect the performance of a real case from Deppon Express.
 However, the models primarily focus on the process of agents transferring between queues rather than on traffic flow and do not fully consider traffic management methods to mitigate conflicts.
-This aspect makes them less suited for analyzing the efficiency of MAPF solvers, such as the method proposed in this paper.
+This makes them less suitable for analyzing the efficiency of MAPF solvers, such as the method proposed in this paper.
 On the other hand, other existing studies often overlooked issues related to robot coordination, relying on the idealized assumption that there are no conflicts or deadlocks.
@@
 First, we propose a novel robot traffic management method to enhance system throughput, ensuring stable and efficient operations in multi-robot collaboration settings while overcoming the complexities of robot interactions.
-Second, we develop a precise model to quantify system efficiency in RSS, enabling accurate throughput estimation to clarify the relationship between system configurations and sorting capacity.
+Second, we develop a precise model to quantify system efficiency in RSS, enabling accurate throughput estimation to reveal the relationship between system configurations and sorting capacity.
 Third, we highlight the need for management strategies that leverage RSS flexibility to provide cost-effective responses to fluctuating demands.
@@
 To fully utilize global information while considering robot operational patterns and safety distance regulations, we adapt a network-level control strategy for coordinating autonomous vehicles and propose a collision-free spatio-temporal path planning method for RSS.
-An efficient heuristic algorithm is proposed to enhance the scalability.
+An efficient heuristic algorithm is introduced to enhance the scalability.
 Second, building upon this control strategy, we derive an estimation formula for throughput in RSS given system configurations, which serves as an efficiency constraint in the planning problem.
-Finally, we propose an optimization model aimed at minimizing the initial investment and average operations costs under fluctuating demands.
+Finally, we propose an optimization model that aims to minimize initial investment and average operational costs under fluctuating demands.
 The solution includes layout design and resource allocation recommendations, specifically on how to adjust the number of robots and workers.
 Experiments provide insights into the cost structure of RSS under varying unit price conditions.
-To the best of our knowledge, this paper is among the first to incorporate the traffic issues of robots during operations into the configuration design of the RSS system.
+To the best of our knowledge, this paper is among the first to incorporate robot traffic issues during operations into RSS configuration design.
 Our intention is to bridge operational and strategic planning in multi-robot coordination contexts, thereby contributing a comprehensive theoretical framework to support decision-making in such complex and flexible systems.
@@
 In Section 4, we introduce an innovative traffic management framework to coordinate multiple robots in a warehouse setting.
-Based on the proposed framework, section 5 presents a system throughput estimation formula, which provides quantitative measurement of layout configuration’s impact to system efficiency.
-In section 6, we demonstrate the superiority of our proposed traffic management framework compared with benchmarks, as well as validate the accuracy of throughput estimation formula.
+Based on the proposed framework, Section 5 presents a system throughput estimation formula, which quantifies the impact of layout configuration on system efficiency.
+Section 6 demonstrates the superiority of our proposed traffic management framework compared with benchmark methods and validates the accuracy of the throughput estimation formula.
 To emphasize the impact of traffic issues on RSS, experiments are conducted to show that queueing network model has estimation biases for throughput in certain scenarios.
-Moving on to section 7, we propose a layout optimization model to minimize the total costs and introduce an efficient solution approach.
-Section 8 presents a sensitivity analysis to investigate the optimal layout design and the corresponding cost structure under different unit cost, and insights distilled from the results.
+Moving on to Section 7, we propose a layout optimization model to minimize total costs and introduce an efficient solution approach.
+Section 8 presents a sensitivity analysis to investigate the optimal layout design and the corresponding cost structure under different unit costs, and insights distilled from the results.
 Finally, section 9 provides a summary of the entire paper.
 2 Literature Review
-The cost savings in labor through automation are particularly crucial in the picking process, and increased sorting efficiency enables parcel companies to offer convenient and low-cost same-day delivery services (Dekhne et al.,, 2019).
-However, the operations research community has scarcely explored the realm of robotized sorting systems, which represent the latest advance of warehouse automation.
-Among the existing literature, studies mainly focus on the efficiency of small parcel sorting systems in distribution centers.
+The cost savings in labor through automation are particularly crucial in the picking process, and improved sorting efficiency enables parcel companies to offer convenient and low-cost same-day delivery services (dekhne2019automation).
+However, the operations research community has scarcely explored the realm of robotized sorting systems, which represent the latest advances in warehouse automation.
+Among the existing literature, studies focus mainly on the efficiency of small parcel sorting systems in distribution centers.
 This body of work delves into the comprehensive assessment of transportation efficiency within the robot fleet under various system configurations.
-In the majority of these studies, robots’ sorting tasks are modeled as services in queuing network models, and the total throughput is the key performance evaluation metrics (Zou and Chen,, 2020; Zi and Gao, 2020a, ; Zi and Gao, 2020b, ; Zou et al.,, 2021; Xu et al.,, 2022).
-Zi and Gao, 2020a ; Zi and Gao, 2020b assumed that each robot moves independently.
-Zou and Chen, (2020) introduced a semi-open queuing network (SOQN) model and subsequently improved its theoretical aspects in their follow-up research (Zou et al.,, 2021), simplifying it to the closed queueing model (CQN) by assuming a sufficient high arrival rate of parcels.
-A queueing model takes into account the queuing behavior of robots at loading stations and outlets, partially addressing the issue of robot occupancy on the aisle.
-Furthermore, their second study compares how the minimum total costs of the system varied under different network topology and different unit cost of system components, given the target sorting efficiency.
-Xu et al., (2022) further investigated an RSS with a parcel-to-loading-station assignment mechanism, which could be regarded as a pre-sorting strategy.
-This study removes the constraint on the number of robots, resulting in an open queuing network (OQN) model.
-Experimental results demonstrates that the introduction of pre-sorting effectively reduces the average travel distance of the robots but might lead to extra congestion upstream at the same time.
-Liu et al., 2019b and Tan et al., (2021) focused on the task assignment problem with the objective of minimizing the sorting makespan.
-Both adopts a travel time model (closed-form travel time expression, defined by Azadeh et al., (2017)), and the second research formulates a mixed-integer programming model.
-Compared to queuing network models, travel time models exhibit significant distortion when the number of robots is high, as they allow multiple robots to simultaneously occupy a facility.
-Boysen et al., (2023) conducted research on order sorting system, which involves additional tasks including Piece-to-Order assignment and Order-to-Collection Point assignment.
-Therefore, the primary focus of the referred study is not solely on scheduling and control of robots, and the travel time of robots to a specific outlet is assumed to be constant.
-Shi et al., (2021)’s study, on the other hand, investigates a human-robot hybrid sorting system, in which the scale of robot sorting is fixed, while the manual capacity is adjustable according to the demand.
-Regrettably, in this context, the considerable flexibility inherent in the robotic system has not been fully harnessed.
+In the majority of these studies, robot sorting tasks are modeled as services in queuing network models, and total throughput is the key performance evaluation metric (9532245; 9216812; 9102081; zou2021robotic; XU2022102808; fang2025dynamic).
+9216812; 9102081 both assumed that each robot moves independently.
+9532245 introduced a semi-open queuing network (SOQN) model and subsequently improved its theoretical aspects in their follow-up research (zou2021robotic), simplifying it to the closed queueing model (CQN) by assuming a sufficiently high parcel arrival rate.
+Their queueing model took into account the queuing behavior of robots at loading stations and outlets, partially addressing the issue of robot occupancy on the aisle.
+Furthermore, their second study compared how the minimum total costs of the system varied under different network topologies and different unit costs of system components, given the target sorting efficiency.
+fang2025dynamic further emphasized the importance of dynamic robot routing and destination assignment policies in handling severe heterogeneity and fluctuations in sorting demand.
+They proposed an SOQN model to analyze network congestion and an MIP model to reassign the destination-outlet mapping in response to drastic changes in sorting demand.
+XU2022102808 investigated an RSS with a parcel-to-loading-station assignment mechanism, which could be regarded as a pre-sorting strategy.
+This study removed the constraint on the number of robots, resulting in an open queueing network (OQN) model.
+Experimental results demonstrated that the introduction of pre-sorting effectively reduces the average travel distance of the robots but might simultaneously lead to additional upstream congestion.
+liu2019multi and TAN2021101279 focused on the task assignment problem with the objective of minimizing the sorting makespan.
+Both adopt a travel time model based on the closed-form travel time expression proposed by pub99983, and the latter formulates a mixed-integer programming model.
+Compared to queuing network models, travel time models exhibit significant deviation when the number of robots is high, as they allow multiple robots to simultaneously occupy a facility.
+BOYSEN2023582 conducted research on an order sorting system, which involves additional tasks including Piece-to-Order assignment and Order-to-Collection Point assignment.
+Therefore, the primary focus of their study is not solely on scheduling and control of robots, and the travel time of robots to a specific outlet is assumed to be constant.
+poms13626, on the other hand, investigates a human-robot hybrid sorting system, in which the scale of robot sorting is fixed, while the manual capacity is adjustable according to the demand.
+Regrettably, in this context, the considerable flexibility inherent in the robotic system has not been fully exploited.
 Table 1: Literature summary on RSS investigation
 Reference System Modeling Objective Congestion & Deadlock
-Liu et al., 2019b parcel sortation TT, MIP MS, C $\times$
+liu2019multi parcel sortation TT, MIP MS, C $\times$
 […]
-Zi and Gao, 2020a parcel sortation Q T $\times$
-Zi and Gao, 2020b parcel sortation Q, MIP T, C $\times$
+9216812 parcel sortation Q T $\times$
+9102081 parcel sortation Q, MIP T, C $\times$
 […]
 […]
-Tan et al., (2021) parcel sortation TT, MIP MS $\times$
-Xu et al., (2022) parcel sortation OQN T $\times$
-Boysen et al., (2023) order sortation TT, MIP T $\times$
+TAN2021101279 parcel sortation TT, MIP MS $\times$
+XU2022102808 parcel sortation OQN T $\times$
+BOYSEN2023582 order sortation TT, MIP T $\times$
+[…]
 This paper parcel sortation TT, MIP T, TC $\checkmark$
@@
 Throughout the existing research, a limited number of studies consider the floor space cost while others relax the constraints on aisle resources.
-Furthermore, finding collision-free paths for multiple vehicles has been proved to be an NP-hard problem (Surynek,, 2010), and some state-of-the-art algorithms also struggle to completely eliminate the issue of road congestion resulting from an increasing number of robots.
-However, except the study by Shi et al., (2021), which employs a fitting method to predict the positive correlation between traffic flow and congestion, most of the studies neglect the mutual influence among multiple robots in aisle occupation for a compact system, leading to an misestimate of traffic capacity.
+Furthermore, finding collision-free paths for multiple vehicles has been proved to be an NP-hard problem (Surynek_2010), and some state-of-the-art algorithms also struggle to completely eliminate the issue of aisle congestion resulting from an increasing number of robots.
+However, except for the study by poms13626, which employs a fitting method to predict the positive correlation between traffic flow and congestion, most of the studies neglect the mutual influence among multiple robots in aisle occupation for a compact system, leading to a misestimate of traffic capacity.
 Table 1 presents a comprehensive overview of the literature reviewed on the RSS, highlighting its relevance to our paper.
 Some researchers recognized the limitations of queuing models in capturing the impact of aisle layout design on multi-robot systems and chose to address the traffic issues by modeling them as MAPF problems.
-Wagner and Choset, (2011) proposed a novel M* algorithm that allowed for the planning of a larger number of robot trajectories.
+wagner6095022 proposed a novel M* algorithm that allowed for the planning of a larger number of robot trajectories.
 However, they faced challenges due to significant memory requirements and time complexity, and could not guarantee a feasible solution.
-Some researchers designed priority-based algorithms and integrated task assignment to reduce waiting time (Nguyen et al.,, 2019; Liu et al., 2019a, ), sacrificing solution optimality in exchange for stability and computational efficiency.
-Li et al., (2021) introduced an RHCR algorithm framework to tackle the windowed MAPF problem, providing an efficient solution for managing newly arrived orders in warehouse operations.
+Some researchers designed priority-based algorithms and integrated them with task assignment to reduce waiting time (nguyen2019generalized; Liu2019path), sacrificing optimality for improved stability and computational efficiency.
+Li_Tinka2021 introduced an RHCR algorithm framework to tackle the windowed MAPF problem, providing an efficient solution for managing newly arrived orders in warehouse operations.
 However, despite their effectiveness, MAPF-based studies have faced challenges in deriving a closed-form expression for throughput estimation, which is crucial for creating reliable and efficient system configurations.
@@
 This study addresses these gaps by first developing a dedicated robot traffic management method for RSS at the operational level.
-Building on this method, we establish an optimal layout and configuration design model aimed at minimizing overall system costs in the long run, ensuring that the results are practically applicable in real-world warehouses.
+Building upon this method, we establish an optimal layout and configuration design model aimed at minimizing overall system costs in the long run, thus enhancing the applicability of the findings to real-world logistics operations.
 3 Problem Description
-In an RSS, three zones are deployed from center to periphery in the field (Wang et al.,, 2021): 1) sorting zone; 2) loading zone; 3) waiting zone, as shown in Figure 2.
+In an RSS, three zones are deployed from center to the periphery of the field (10.1007/978-981-16-8174-5_10): 1) sorting zone; 2) loading zone; 3) waiting zone, as shown in Figure 2.
 Parcels are loaded to robots by workers in the loading zone.
-Robots drop off each parcel to its target outlet in the sorting zone to fulfill a sorting task.
-Waiting zone is the space where robots idle and queue behind loading stations.
-The area of every zone is determined in the design phase, and remains unchanged regardless of sorting demands, number of robots and workers.
-Let $W_{l}$ and $W_{w}$ denote the width of loading zone and waiting zone, respectively, which are considered to be constant in this study.
-The decision of site planning mainly focuses on the design of sorting zone, particularly the configuration of the aisle network.
+Robots drop off each parcel at its target outlet in the sorting zone to fulfill a sorting task.
+The waiting zone is the space where robots idle and queue behind loading stations.
+The area of every zone is determined during the design phase and remains unchanged regardless of sorting demands or the numbers of robots and workers.
+Let $W_{l}$ and $W_{w}$ denote the widths of the loading zone and the waiting zone, respectively, both considered constant in this study.
+The decision of site planning mainly concerns the configuration of the aisle network in the sorting zone.
 Figure 2: Layout of RSS
 Without loss of generality, we make the following assumptions in this study:
-  • The parcels arrive in batch manner.
-When a parcel is loaded on a robot, the next parcel will arrive at the loading station immediately.
-That is, the loading time is negligible, and this parameter is considered implicitly in the loading rate of worker.
-  • The arrival rate of parcels destinating for each outlet at each loading station is the same (When the sorting demand is unevenly distributed, this assumption can be achieved by increasing or decreasing the number of outlets corresponding to each destination.
-Additionally, outlets with extremely low and high arrival rates can be placed together to balance the flow);
-  • The time for robots dropping parcels into outlets is negligible (To satisfy this assumption, robots need to release goods in advance based on its speed before reaching the target outlet, similar to the behavior of conveyor-based sorters.
-If it is difficult to achieve due to technical and cost issues, a 2-second drop-off time per delivery must be introduced (Zou et al.,, 2021));
-  • The drop-off operation by robots around one outlet does not affect each other, i.e., each outlet allows at most four robots releasing parcels simultaneously;
-  • Each robot can only carry one parcel during one delivery process;
-  • Loading stations are evenly distributed in the loading zone;
-  • The transportation of parcels after being collected in the outlet is not considered in this research.
-  • The battery charging needs of robots are not considered.
+  • The parcels arrive in batches.
+When a parcel is loaded onto a robot, the next parcel will arrive at the loading station immediately.
+That is, the loading time is negligible and is implicitly captured in the loading rate of the workers.
+  • The arrival rate of parcels destined for each outlet at each loading station is identical.
+(When sorting demand is unevenly distributed, this can be achieved by adjusting the number of outlets corresponding to each destination.
+Additionally, outlets with extremely low or high arrival rates may be grouped together to balance the flow.)
+  • Robots release parcels without slowing down or stopping, and drop-off time is considered negligible.
+(To satisfy this assumption, robots must release parcels in advance based on their speed before reaching the target outlet, similar to the behavior of conveyor-based sorters.);
+  • Drop-off operations around an outlet are independent, i.e., each outlet allows at most four robots releasing parcels simultaneously.
+  • Each robot can carry only one parcel during one delivery process;
+  • Once a robot exits the network through an exit node, it immediately joins the queue of the associated loading station without repositioning between loading stations.
+Travel time within the waiting and loading zones is neglected.
+  • The transportation of parcels after they are collected at the outlet is not considered.
+  • Battery charging requirements of robots are not considered.
 (a) Sample gird layout of a compact $4+4$ aisle network
 (b) Graph representation of sample layout
-Figure 3: Illustration of aisle network in sorting zone
+Figure 3: Illustration of aisle network in the sorting zone
 To better evaluate the performance of this system with high loads and complicated sorting demands, the microscopic traffic management of robots in the sorting zone needs careful investigation.
@@
 We use the term “outlet” to refer to the special cells designated for collecting parcels delivered by robots.
-Outlets are evenly distributed throughout the sorting zone, divided by aisles.
+Outlets are evenly distributed throughout the sorting zone, seperated by aisles.
 This layout allows robots to deliver parcels to outlets on both sides of the aisle and ensures that the outlets are arranged with sufficient density.
-Each loading station in the loading zone is connected to a pair of dedicated entrance and exit, which are adjacent to each other.
-Therefore, the number of loading stations cannot exceed the the number of aisles.
+Each loading station in the loading zone is connected to a pair of adjacent entrance and exit nodes.
+Therefore, the number of loading stations cannot exceed the number of aisles.
+A dedicated worker is required at each loading station to perform parcel loading operations.
+A loading station is defined as active if a worker is assigned and engaged in tasks at that station.
 Figure 3(b) shows the graph representation of the sample layout in Figure 3(a).
-In the graph, each cell that makes up the aisles is modeled as either a conflict node (blue dot), unloading node (red hexagon), or entrance/exit node (green/yellow square), with edges of length $D$ connecting adjacent nodes.
+In the graph, each cell that makes up the aisles is modeled as either a conflict node (blue dot), an unloading node (red hexagon), or an entrance/exit node (green/yellow square), with edges of length $D$ connecting adjacent nodes.
 Conflict nodes are at the intersections of aisles, unloading nodes are where robots deliver parcels, and entrance/exit nodes are at aisle ends.
@@
 Robots possess the capability of in-place rotation, with angular velocity denoted as $\omega_{r}$.
-To ensure collision-free movement of robots, the robot management system needs to avoid conflicts in path finding.
-Common conflicts, as illustrated in Figure 4, include the following four types (Stern et al.,, 2019): (a) Swapping conflict occurs when two robots are planned to swap locations at the same time.
-It does not exist in a one-way network; (b) Vertex conflict occurs when more than two robots are planned to occupy the same vertex; (c) Following conflict occurs when one robot is planned to occupy a vertex that was occupied by another robot in the previous timestep.
-The latter is stationary or moving at a lower speed; (d) Deadlock conflict occurs among multiple robots, with each robot having a following conflict with another robot.
+To ensure collision-free movement of robots, the robot management system needs to avoid conflicts in path planning.
+Common conflicts, as illustrated in Figure 4, include the following four types (stern2019multi): (a) Swapping conflict occurs when two robots are planned to swap locations at the same time.
+It does not exist in a one-way network; (b) Vertex conflict occurs when more than two robots are planned to occupy the same vertex; (c) Following conflict occurs when one robot is planned to occupy the vertex currently occupied by another robot, while the latter is either stationary or moving at a lower speed; (d) Deadlock conflict involves multiple robots, each trying to maintain the minimum headway with another, resulting in mutual blockage.
 Among these conflicts, both swapping conflicts and deadlock conflicts necessitate a re-planning of paths.
 The occurrence of either can result in significant losses in system efficiency.
+In this study, the RSS layout planning problem is structured as a two-stage cost-minimization model subject to throughput constraints, integrating long-term layout design with operational-level traffic control of robots:
+  1. 1.
+    Site planning stage (strategic level): The objective is to minimize the total cost, including facility cost $C_{f}$ and expected operations cost $C_{o}$ across different operating periods $\sigma\in\mathcal{S}$:
+    $\min\ C_{d}=C_{f}+C_{o}$ (1)
+    Each period $\sigma$ is associated with a required throughput $T^{\sigma}$, representing changes in sorting demand.
+Decision variables include the number of horizontal aisles $n_{h}$, vertical aisles $n_{v}$, loading stations $n_{l}$, and the number of workers $n_{w}^{\sigma}$ and robots $n_{r}^{\sigma}$ allocated in each period $\sigma$.
+Layout and resource configurations must meet the throughput requirements:
+    $\tilde{T}_{O}(n_{h},n_{v},n_{w}^{\sigma},n_{r}^{\sigma})\geq T^{\sigma}\qquad\forall\sigma\in\mathcal{S}$ (2)
+    where $\tilde{T}_{O}$ is the estimated system throughput.
+A trade-off exists between $C_{f}$ and $C_{o}$: larger networks increase facility costs and travel distances, while smaller layouts may cause congestion and reduce network capacity.
+  2. 2.
+    Operations stage (operational level): Given a fixed layout and resource configurations, the goal is to coordinate robot traffic efficiently to achieve stable and predictable system throughput.
+The robot traffic management method regulates robot movements and trajectory assignments, directly determining the system throughput.
+The estimation of system throughput $\tilde{T}_{O}$ serves as a critical link between operational performance and upper-level planning.
 Figure 5: Overview of research methodology
-Planing problems related to RSS can be divided into two stages: the site planning stage and operations stage.
-In the site planning stage, warehouse manager concentrates on the division of zones for different functions in the warehouse.
-The area of an RSS should be large enough to fulfill the designated throughput level in peak season, and the number of outlets should be no fewer than the number of individual sorting destinations.
-The decision variables of site planning stage are the number of horizontal and vertical aisles $n_{h},n_{v}$ and the number of loading stations $n_{l}$; in the operations stage, warehouse manager should adaptively adjust the number of workers and robots to minimize operations costs, given expected sorting demands.
-Since a warehouse is usually rented for years, it is advisable to keep the zone division unchanged during its operation.
-The decision variables of operations stage are the number of workers and robots in operating period $\sigma$, noted as $n_{w}^{\sigma},n_{r}^{\sigma}$, respectively.
-The goal of RSS planning problem is to minimize the total costs, which consists of two parts: the facility costs $C_{f}$, and the operations costs $C_{o}$.
-The size of the site holds a trade-off between $C_{f}$ and $C_{o}$: A small network leads to lower capacity, thereby increasing the waiting time of robots and raising the rental cost of robots in operations costs; whereas, an overly large network leads to increased site costs and longer travel distance.
-An overview of research methodology is shown in Figure 5.
-In sections 4-6, we present the traffic management of robots, conduct system efficiency analysis, and validate the performance and accuracy of estimation formula, respectively.
-In section 7, a layout design optimization model is formulated to minimize the total costs of RSS given the demand.
+Figure 5 provides an overview of this integrated framework.
+In Sections 46, we introduce the robot traffic management approach, analyze the system efficiency, and validate the performance and accuracy of the proposed throughput estimation method.
+In Section 7, we formulate the layout optimization model to minimize the total cost of RSS under given sorting demand.
 4 Management of High-density Robot Traffic in RSS
-Managing a large number of robots within a dense network poses significant challenges, especially when balancing system throughput performance with the computational load of algorithms under unpredictable traffic conditions.
-To address multi-robot path finding and eliminate traffic gridlock within an admissible computational duration, this study adopts an innovative autonomous vehicle management scheme, called Rhythmic Control (RC) (Lin et al.,, 2021).
-It enables the planning of uninterrupted scheduling for robot fleets.
+Managing a large number of robots within a dense network poses significant challenges, especially when balancing system throughput with the computational load of algorithms under unpredictable traffic conditions.
+To address the multi-robot path finding problem and eliminate traffic gridlock within an acceptable computational time, this study adopts an innovative autonomous vehicle management scheme called Rhythmic Control (RC) (RCII), which enables uninterrupted scheduling of robot fleets.
 In this section, we begin with the key concepts inherited from the original RC: the virtual platoon and the cycle.
-By leveraging the reservation mechanism, we propose a new centralized framework called Rhythmic Control in Sorting Scenario (RC-S).
-RC-S presents high orderliness and serves as the foundation for theoretical analysis of system efficiency, which will be discussed in detail in section 5.
+By leveraging the reservation mechanism of RC, we propose a new centralized framework called Rhythmic Control for the Sorting Scenario (RC-S).
+RC-S exhibits a high degree of orderliness and serves as the foundation for the theoretical analysis of system efficiency, which will be discussed in detail in Section 5.
 4.1 Virtual platoon and cycle
-Lin et al., (2021) investigated a method of incorporating rhythm into traffic management at the macro-traffic level and introduced the concept of virtual platoon (VP).
-VPs represent the spatiotemporal slot that is generated in a rhythmic manner and keep uniform linear motion.
-As shown in Figure 6(a), VPs move along a designated spatiotemporal trajectory and maintain a constant safe distance between each other.
-Figure 6(b) shows the spatiotemporal trajectory of VPs in a four-lane network, without any collision or deceleration.
-We discretize time based on the generation time of VPs at the entrance, and each time interval is noted as a cycle.
+RCII investigated a method of incorporating rhythm into traffic management at the macro-traffic level and introduced the concept of virtual platoon (VP).
+A VP is a reserved spatio-temporal slot along a lane, characterized by a uniform linear trajectory and generated at fixed intervals.
+Each VP could be identified by its corresponding lane and the interval in which it is released.
+As shown in Figure 6(a), VPs move along a designated spatio-temporal trajectory and maintain a constant safe distance between each other.
+Figure 6(b) shows the spatio-temporal trajectory of VPs in a four-lane network, without any collision or deceleration.
+We discretize time based on the generation time of VPs at the entrance, and each time interval is referred to as a cycle.
 The duration of each time interval is the cycle length.
-By staggering the entry times of VPs at different entrances within the same cycle, VPs pass one intersection alternately with relatively short headway and the intersection capacity could be maximized.
-Whenever a vehicle needs to pass through a lane, it is mandated to follow the movement of one VP in that lane until it reaches its destination and exits the lane.
+By staggering the entry times of VPs at different entrances within the same cycle, VPs pass a intersection alternately with relatively short headway and the intersection capacity could be maximized.
+Whenever a vehicle needs to pass through a lane, it is required to follow the movement of a VP in that lane until it reaches its destination and exits the lane.
 Before entering the network, the vehicle must wait for an unoccupied VP at the entrance.
-It first sends its destination to the control system, which then plans a conflict-free route and reserves one unit of capacity from the corresponding VPs in different segments of the route.
+It first sends its destination to the control system, which then plans a conflict-free route and reserves one unit of capacity in the corresponding VPs in different segments of the route.
 The capacity is released when the vehicle leaves each VP.
@@
 (a) Virtual platoons in a four-lane network
-(b) Trajectory of virtual platoons in different directions. (Lin et al.,, 2021)
+(b) Trajectory of virtual platoons in different directions. (RCII)
 Figure 6: Concept of Rhythmic Control
 RC is essentially a traffic control method, similar to a combination of globally coordinated traffic signal control and road network flow control.
-Within this framework, it is essential to clarify the movement patterns of each entity for integration into a sorting-specific robot management system.
-For detailed information on the RC framework, we refer readers to Lin et al., (2021).
-The following points provide a foundational overview critical to our study on the implementation of the RC-S scheme:
-  • A VP represents a time–space slot that is available for being occupied by vehicles and is generated in a rhythmic manner in each entrance of the network.
-  • VPs move in a straight line at a constant speed while maintaining a fixed distance, ensuring that the total number of VPs present in the road network remains the same at any given moment.
-  • After entering the network, vehicles need to continuously move along with a specific VP or transfer from one VP to another until exiting through an exit.
-  • At any given moment, the number of vehicles accommodated by a VP cannot exceed its capacity.
-If the required VP capacity for vehicle travel is insufficient, the vehicle needs to queue and wait at the entrance until entering the network in a subsequent cycle.
+Within this framework, it is essential to clarify the movement patterns of each entity so that RC can be integrated into a sorting-specific robot management system.
+For detailed information on the RC framework, we refer the reader to RCII.
+The following points provide a foundational overview that is critical to our study on the implementation of the RC-S scheme:
+  • A virtual platoon (VP) represents a spatio-temporal slot that can be reserved and occupied by robots and is generated in a rhythmic manner at each entrance of the network.
+  • VPs move in a straight line at a constant speed while maintaining a fixed spacing, ensuring that the total number of VPs present in the aisle network remains constant.
+  • After entering the network, a robot must continuously follow a specific VP or transition to another VP until it exits.
+  • The capacity of each VP is one, which means at most one robot can occupy it at any time.
+  • A robot may enter the network only when the required VP capacity is available; otherwise, it must wait for a later cycle.
 4.2 Concept of RC-S
 In order to embody RC in the RSS, we regulate the behavior of robots, and propose the RC-S scheme, which can be regarded as a centralized MAPF solver.
-To focus on the behavior of each individual robot, we discretize the map into grids, defining each grid where robots can pass through as a node.
-Given $(n_{v},n_{h})$, the network in sorting zone is represented by a directed graph $\mathcal{G}=(\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{e},\mathcal{E})$, where $\mathcal{V}_{c}$, $\mathcal{V}_{u}$, $\mathcal{V}_{e}$ and $\mathcal{E}$ stand for the set of conflict nodes, unloading nodes, entrance/exit nodes and edges, respectively.
-To reduce intersections, we adopt a one-way network, where aisles from different directions are arranged in an interleaved pattern, as shown in Figure 7.
-Based on the network structure, each VP can accommodate at most one AGV.
-Let $\tau_{e}$ and $\tau_{c}$ denote the fixed travel time of VPs on each unit cell in the grid map and the cycle length of rhythmic control, respectively.
 Under RC-S, the behaviors of robots are limited to three actions: moving straight, dropping off parcel, and turning.
-Robots enter the network with fixed interval $\tau_{c}$, thus the upper bound of robot flow on each grid is $1/\tau_{c}$.
+We discretize the map into grid-cells, defining each grid-cell where robots can pass through as a node.
+Given $(n_{v},n_{h})$, the network in the sorting zone is represented by a directed graph $\mathcal{G}=(\mathcal{V}=\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{en}\cup\mathcal{V}_{ex},\mathcal{E})$, where $\mathcal{V}_{c}$, $\mathcal{V}_{u}$, $\mathcal{V}_{en}$, $\mathcal{V}_{ex}$ and $\mathcal{E}$ stand for the set of conflict nodes, unloading nodes, entrance nodes, exit nodes and edges, respectively.
+To reduce potential conflicts and increase the density of outlets, we adopt a one-way network, where aisles from different directions are arranged in an interleaved pattern, as shown in Figure 7.
+Based on the network structure, each VP can accommodate at most one robot.
+Let $\tau_{e}$ and $\tau_{c}$ denote the fixed travel time of VPs on each grid-cell and the cycle length of rhythmic control, respectively.
+That is, at each entrance node, VPs enter the network at fixed intervals of $\tau_{c}$.
 Figure 7: Illustration of RC-S. (a)-(d) are the four phases in one cycle.
-In each cycle, the distribution of VPs could be divided into four phases (Figure 7(a)-(d)), and the duration of each phase is $\tau_{e}$. We have:
+In each cycle, the distribution of VPs could be divided into four phases (Figure 7(a)-(d)), and the duration of each phase is $\tau_{e}$.
+Each VP moves forward one grid-cell along its assigned direction after $\tau_{e}$, transitioning the system from one distribution phase to the next, in the sequence: (a) $\rightarrow$ (b) $\rightarrow$ (c) $\rightarrow$ (d) $\rightarrow$ (a) $\rightarrow\dots$.
+We could obtain:
 $\tau_{c}=4\tau_{e}$ (3)
-$\tau_{e}$ and $\tau_{c}$ are determined by the maximum velocity of robot $v_{max}$ and the maximum loading rate of each loading station $r_{l}$, respectively:
+$\tau_{e}$ and $\tau_{c}$ are restricted by the maximum velocity of robot $v_{max}$ and the maximum loading rate of each loading station $r_{l}$, respectively:
 $\tau_{e}\geq\frac{D}{v_{max}}$ (4)
 $\tau_{c}\geq\frac{1}{r_{l}}$ (5)
-where $D$ is the length of each grid.
-At phase(a) of a cycle, new VPs are released at the entrance of each aisle; at phase(c), VPs leave sorting zone through the exit of each aisle.
-Thus, if we do not count the VPs at the exit of each aisle, then the number of VPs in the network remains the same throughout the entire sorting process.
-Intersections are occupied by VPs at phases(a) and (c); each outlet could be served by four VPs in four different directions at phases(b) or (d).
+where $D$ is the length of each grid-cell.
+At phase(a) of a cycle, new VPs are released at the entrance node of each aisle; at phase(c), VPs leave the sorting zone through the exit node of each aisle.
+Thus, if we do not count the VPs at the exit nodes, then the number of VPs in the network remains the same throughout the entire sorting process.
+Each intersection is occupied by two different VPs at phases (a) and (c) within a cycle, respectively.
+Each outlet could be served by four VPs in four different directions at phases(b) or (d).
 The fixed number of VPs and the regularity in their phases allow us to manage and analyze this system from a global perspective.
@@
 $v_{VP}$ $=\frac{D}{\tau_{e}}$ (6)
-Figure 8: Illustration of turning and transfer process for one robot (red dot)
-RSS does not require auxiliary turning lanes since robots are capable to rotate within the intersection grid.
-The process of turning and transfer between VPs for robots are demonstrated in Figure 8.
-As shown in the figure, the robot to turn first decelerates to zero, stops following its original VP, rotates 90 degrees at the intersection, joins a new VP and follows the movement of it.
-Turning action entails one cycle for both the original and subsequent VPs, leading to a reduction in traffic capacity.
-During this process, the parameters of RC-S needs to satisfy the following conditions:
+Figure 8: Illustration of turning and transfer process for a robot
+Figure 9: (a) Intersection (b) Trajectory of a turning robot
+RSS does not require auxiliary turning lanes, since robots are able to rotate within the intersection grid-cell.
+The process of robot turning and transfer between VPs is demonstrated in Figure 8.
+As shown in the figure, the turning robot first decelerates to zero, stops following its original VP, rotates 90 degrees at the intersection, joins a new VP, and then follows its movement.
+The turning action occupies one cycle in both the original and the subsequent VPs, leading to a reduction in traffic capacity.
+During this process, the parameters of RC-S must satisfy the following conditions:
 $d(\tau_{e},v_{VP},v_{max},c_{max})\geq 2D$ (7)
 $\tau_{e}\geq\frac{\pi}{2\omega_{r}}$ (8)
-Constraint (5) implies that the robot can catch up with the VP in a turn, where $d(\tau_{e},v_{VP},v_{max},c_{max})$ denotes the maximum travel distance a robot can cover in acceleration/deceleration process.
-The expression is derived in detail in Appendix B; Constraint (6) imposes limits on the rotation speed $\omega_{r}$, ensuring that the robot can complete a 90-degree turn within one phase.
-In summary, using expressions (1)-(6) and the kinematic parameters of the robots, we could determine the parameters for RC-S.
-The trajectory of robot in one delivery task should start from one loading station, pass the target outlet corresponding to the destination of parcel, and end at another loading station.
-We define feasible path in RC-S:
+Constraint (7) imposes a restriction on $\tau_{e}$ based on the robot’s maximum acceleration/deceleration rate $c_{max}$.
+The function $d(\tau_{e},v_{VP},v_{max},c_{max})$ represents the maximum distance that a robot can cover during acceleration and deceleration, and its derivation is provided in Appendix B.
+Constraint (8) sets a minimum requirement on $\tau_{e}$ to ensure that the robot can complete a 90-degree turn within one phase, given its rotation speed $\omega_{r}$.
+As a result, using Expressions (3)–(8) together with the kinematic parameters of the robots, we could determine the parameters for RC-S.
+The trajectory of a robot for a single delivery task starts at a loading station, passes the target outlet corresponding to the parcel’s destination, and ends at another loading station.
+We define a feasible path under the RC-S scheme as follows:
 DEFINITION 1.
-A feasible path in RC-S is a sequence of VPs and their occupying cycles.
-The trajectories traveled by these VPs during the occupied cycles are connected end-to-end, linking two active loading stations at the periphery of the network.
-Definition 1 defines the candidate spatio-temporal paths for robots under RC-S, where an active loading station indicates that a worker is currently engaged in tasks.
-When a robot performs a sorting task, it moves from one loading station $i$ to another loading station $j$, $i,j\in\mathcal{L}$, where $\mathcal{L}$ is the set of all loading stations.
+A feasible path in RC-S is a sequence of VP–cycle pairs, denoted by $\{(p_{1},l_{1}),(p_{2},l_{2}),\ldots,(p_{n},l_{n})\}$, where $(p_{i},l_{i})$ indicates that the robot occupies VP $p_{i}$ during cycle $l_{i}$.
+The sequence must satisfy the following conditions:
+  1. 1.
+    Start and end points: The first VP $p_{1}$ originates from an entrance node $\nu_{en}\in\mathcal{V}_{en}$ in cycle $l_{1}$, and the last VP $p_{n}$ ends at an exit node $\nu_{ex}\in\mathcal{V}_{ex}$ in cycle $l_{n}$.
+  2. 2.
+    Temporal consistency: For every consecutive pair $(p_{i},l_{i})$ and $(p_{i+1},l_{i+1})$:
+    • If $p_{i+1}=p_{i}$, then $l_{i+1}=l_{i}+1$ (the robot remains in the same VP).
+    • Otherwise, $l_{i+1}=l_{i}$ (the robot transitions between VPs).
+  3. 3.
+    Spatial connectivity: If $p_{i}\neq p_{i+1}$, then $p_{i}$ and $p_{i+1}$ must intersect at a common conflict node $\nu_{c}$ from different directions, and the corresponding entry times $t_{i}$ and $t_{i+1}$ must satisfy:
+    $\tau_{c}\cdot(l_{i}-1)\leq t_{i}<t_{i+1}<\tau_{c}\cdot\left(l_{i}+\frac{1}{2}\right),\quad t_{i+1}=t_{i}+\frac{\tau_{c}}{2}$
+    where $\tau_{c}$ denotes the cycle length.
+PROPOSITION 1.
+Given $\mathcal{G}=(\mathcal{V},\mathcal{E})$, each feasible path under the RC-S control scheme can be represented as a sequence of node–cycle pairs $\{(\nu_{i},l_{i})\}$.
+The proof of Proposition 1 is provided in Appendix B.
+Definition 1 specifies the candidate spatio-temporal trajectory patterns for robots under RC-S.
+Condition 1 ensures that the trajectory connects valid start and end nodes on the network boundary.
+Conditions 2 and 3 guarantee continuity and compliance with RC-S by restricting robot movements in accordance with the RC-S rules.
+Moreover, since robots are not permitted to stop except when turning, feasible paths cannot contain transfers between VPs in the same direction.
+Proposition 1 enables the transformation of VP reservations into node–cycle pairs, forming the basis for the optimization model presented later.
+When a robot performs a sorting task, it moves from one loading station $i$ to another loading station $j$, with $i,j\in\mathcal{L}$, where $\mathcal{L}$ denotes the set of all loading stations.
 Let $\mathcal{O}$ represent the set of outlets.
-Then, we denote the set of all feasible path as $\mathcal{R}=\cup_{i\in\mathcal{L},j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(i,k,j)$, where $\mathcal{R}(i,k,j)$ is the subset containing the paths start from loading station $i$, pass by outlet $k$ and end at loading station $j$.
-To simplify the notation, we use $\mathcal{R}_{i,k}$ to represent the set of all feasible paths originating from station $i$ and passing by outlet $k$, and ^$\hat{\mathcal{R}}_{i}$ to represent the set of all feasible paths heading for station $i$, that is, $\mathcal{R}_{i,k}=\cup_{j\in\mathcal{L}}\mathcal{R}(i,k,j)$ and ^$\hat{\mathcal{R}}_{i}=\cup_{j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(j,k,i)$.
-In each cycle, the feasible path assignment problem (FPA) for all the loaded robots is formulated as follows.
-$\underset{x_{i,k}^{r},\hat{x}_{i,k}}{\min}\ \sum_{i\in\mathcal{L}}\sum_{k\in\mathcal{O}}(\sum_{r\in\mathcal{R}_{i,k}}c_{i,k}^{r}x_{i,k}^{r}+\hat{c}_{i,k}\hat{x}_{i,k})$
-$s.t.\quad$ $\sum_{i\in\mathcal{L}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}\delta_{i,k}^{r,\nu,l}x_{i,k}^{r}\leq N_{\nu}^{l}$ $\forall\nu\in\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{e},l\in\mathcal{C}$ (7)
-$\hat{x}_{i,k}+\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=d_{i,k}$ $\forall i\in\mathcal{L},k\in\mathcal{O}$ (8)
-$\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=\sum_{j\in\mathcal{L}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{j,k}\cap\hat{\mathcal{R}}_{i}}x_{j,k}^{r}$ $\forall i\in\mathcal{L}$ (9)
-$x_{i,k}^{r},\hat{x}_{i,k}\in\{0,1\}$ $\forall i\in\mathcal{L},k\in\mathcal{O},r\in\mathcal{R}_{i,k}$ (10)
-There are two sets of variables: $x_{i,k}^{r}$ equals one if the feasible path $r\in\mathcal{R}_{i,k}$ is reserved in current cycle by a robot, and zero otherwise; $\hat{x}_{i,k}$ equals one if the robot at loading station $i$ is heading for outlet $k$ and is kept waiting until the next cycle, and zero otherwise.
-In the objective function, $c_{i,k}^{r}$ is the travel time on path $r\in\mathcal{R}_{i,k}$, and $\hat{c}_{i,k}$ is the penalty of delay.
-The objective function is to minimize the total costs in this cycle.
-Constraint (7) indicates that in each cycle, at most one robot can be present at node $\nu$, where $N_{\nu}^{l}$ is the updated remaining capacity in node $\nu$ in cycle $l$, $N_{\nu}^{l}\leq 1$.
-Parameter $\delta_{i,k}^{r,\nu,l}$ equals one if feasible path $r$ will occupy node $\nu$ in cycle $l$ and zero otherwise.
-Constraint (8) describes the path allocation results for a robot at loading station $i$, where $d_{i,k}$ equals one if the robot is heading for outlet $k$, and zero otherwise.
-Constraint (9) is the conservation of flow constraint, ensuring a stable number of robots at each loading station.
-In large-scale network, the number of feasible paths rapidly increases.
-It is difficult to solve FPA within a reasonable amount of time.
-Here we propose a heuristic algorithm to generate a feasible solution in a relatively short time, as shown in Algorithm 1.
-To expand the search depth, feasible paths starting at some future cycle are allowed to be reserved.
-We further define the entry cycle:
+We denote by $\mathcal{R}$ the set of all feasible paths, that is, $\mathcal{R}=\bigcup_{i\in\mathcal{L},j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(i,k,j)$, where $\mathcal{R}(i,k,j)$ is the subset containing feasible paths starting from loading station $i$, passing outlet $k$, and ending at loading station $j$.
+To simplify the notation, we use $\mathcal{R}_{i,k}$ to represent the set of all feasible paths originating from station $i$ and passing outlet $k$, and $\hat{\mathcal{R}}_{i}$ to represent the set of all feasible paths terminating at station $i$, that is, $\mathcal{R}_{i,k}=\cup_{j\in\mathcal{L}}\mathcal{R}(i,k,j)$ and $\hat{\mathcal{R}}_{i}=\cup_{j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(j,k,i)$.
+The number of paths in $\mathcal{R}(i,k,j)$ is theoretically infinite, since without additional constraints, each feasible path could include an arbitrary number of turns.
+We limit the total number of feasible paths by allowing at most three turns and prohibiting robots from stopping within the network except for turning.
+These restrictions ensure that $\mathcal{R}$ is finite and that each feasible path uses no more than four distinct VPs.
+The rationale behind this choice will be explained in the analysis of the average travel distance in the next section.
+To expand the search space, feasible paths starting at future cycles should also be considered. We further define the entry cycle:
 DEFINITION 2.
-The entry cycle of a specific feasible path is the earliest cycle at which all VPs in the path are unoccupied.
-Algorithm 1 A heuristic method to implement RC-S
-0:  Load the reservation table that records the status of VPs on the map within $N_{p}$ cycles. Set the maximum search range in each cycle as $N_{t}$. Calculate feasible path set $\mathcal{R}$.
-0:  At the beginning of cycle $t$, add the first robot of each queue at entrance to the candidate list.
-0:  If the candidate list is empty, go to step 5; else, find the robot with the longest waiting time in the candidate list, obtain its origin $i$ and target outlet $k$.
-Calculate the entry cycle of each feasible path in $\mathcal{R}_{i,k}$, and choose the feasible path with the smallest entry cycle.
-0:  If the entry cycle is less than $N_{t}$, reserve the VPs according to the feasible path, and assign the path to the robot; else, add the waiting time of the robot by $\tau_{c}$.
-0:  Remove the robot from the candidate list and go to step 2.
-0:  Record the newly generated VPs in the reservation table, labeled as unoccupied. Remove the VPs which have left the network. Output the assignment of feasible path.
-So far, we have identified the essential components required to integrate heuristics into the RC-S framework: recording the status of VPs ever existed, and for each delivery task, finding a feasible path with an entry cycle, and reserving the VPs in the space-time map.
-To improve robot transporting efficiency and ensure the accessibility of each outlet, we allow up to three turns in each delivery route under this scheme.
-Compared with classic MAPF methods, it avoids potential path conflicts from the design stage and greatly reduces the search range while maximizing intersection capacity.
-The optimal strategy of RC-S requires selecting a combination of feasible path and entry time for each robot, and checking the occupancy status of VPs.
-Algorithm 1 has a complexity of $O(N_{t}\cdot\lvert R\rvert)$ in each cycle.
-Moreover, when the algorithm is forcibly terminated midway, it can still produce a feasible solution, where robots that are not served will wait until the next cycle.
-The performance of Algorithm 1 will be validated in section 6.
+The entry cycle of a specific feasible path is the earliest cycle in which all VPs in the path are unoccupied.
+A feasible path corresponds to a specific spatio-temporal trajectory pattern, represented by the sequence of VP reservations it occupies.
+The entry cycle denotes the earliest possible release cycle for the trajectory pattern, given the current VP reservation status.
+Before the beginning of each cycle, the RC-S scheme performs path planning for all loaded robots that are waiting at entrance nodes and do not yet have assigned paths.
+These robots execute their sorting tasks based on the spatio-temporal trajectories specified by their assigned feasible paths.
+The corresponding feasible path assignment (FPA) problem is formulated as an integer programming model with binary decision variables.
+The objective is to minimize the total completion time of all unassigned tasks in the current cycle, together with penalties for any task that cannot be assigned a feasible path.
+To reduce computational complexity, VP reservations from previously assigned paths are treated as fixed input.
+Consequently, the current planning process is based on the real-time availability of VP resources, leading to a greedy insertion approach.
+The main notation used in the FPA model is listed in Table LABEL:Notataion_fpa.
+Table 2: Notations and explanation
+Notations Explanation
+Sets
+$\mathcal{V}$ set of all nodes
+$\mathcal{L}$ set of all loading stations
+$\mathcal{O}$ set of all outlets
+$\mathcal{C}$ set of all cycles starting from the current cycle
+$\mathcal{R}_{i,k}$ set of feasible path patterns starting from loading station $i\in\mathcal{L}$ and passing by outlet $k\in\mathcal{O}$, independent of starting cycle
+$\mathcal{R}_{i}^{(0)}$ Set of feasible paths available to loading station $i\in\mathcal{L}$ at the beginning of current cycle, based on the current destination outlet of the pending sorting task.
+If no task is pending, $\mathcal{R}_{i}^{(0)}=\emptyset$
+$\hat{\mathcal{R}}_{i}$ set of feasible paths ending at loading station $i\in\mathcal{L}$
+Parameters
+$c_{i}^{r}$ completion time of feasible path $r\in\mathcal{R}_{i}^{(0)}$, measured in number of cycles
+$\hat{c}_{i}$ penalty for delaying the assignment of a feasible path to a sorting task from loading station $i\in\mathcal{L}$ until the next cycle
+$\delta_{i}^{r,t,\nu,l}$ incidence between feasible path $r\in\mathcal{R}_{i}^{(0)}$ starting at cycle $t\in\mathcal{C}$, node $\nu\in\mathcal{V}$ and cycle $l\in\mathcal{C}$
+$N_{\nu}^{l}$ remaining capacity in node $\nu\in\mathcal{V}$ in cycle $l\in\mathcal{C}$
+$d_{i}$ sorting demand from loading station $i\in\mathcal{L}$, $d_{i}\in\{0,1\}$
+Variables
+$x_{i}^{r,t}$ decision on whether feasible path $r\in\mathcal{R}_{i}^{(0)}$ starting at cycle $t\in\mathcal{C}$ is reserved
+$\hat{x}_{i}$ decision on whether the sorting task from loading station $i\in\mathcal{L}$ should be postponed to the next cycle
+$\underset{x_{i}^{r,t},\hat{x}_{i}}{\min}\ \sum_{i\in\mathcal{L}}[\sum_{r\in\mathcal{R}_{i}^{(0)}}\sum_{t\in\mathcal{C}}(c_{i}^{r}+t)x_{i}^{r,t}+\hat{c}_{i}\hat{x}_{i}]$
+$s.t.\quad$ $\sum_{i\in\mathcal{L}}\sum_{r\in\mathcal{R}_{i}^{(0)}}\sum_{t\in\mathcal{C}}\delta_{i}^{r,t,\nu,l}x_{i}^{r,t}\leq N_{\nu}^{l}$ $\forall\nu\in\mathcal{V},l\in\mathcal{C}$ (9)
+$\hat{x}_{i}+\sum_{r\in\mathcal{R}_{i}^{(0)}}\sum_{t\in\mathcal{C}}x_{i}^{r,t}=d_{i}$ $\forall i\in\mathcal{L}$ (10)
+$x_{i}^{r,t},\hat{x}_{i}\in\{0,1\}$ $\forall i\in\mathcal{L},r\in\mathcal{R}_{i}^{(0)},t\in\mathcal{C}$ (11)
+We introduce two sets of binary decision variables.
+Variable $x_{i}^{r,t}$ equals one if the feasible path $r\in\mathcal{R}_{i}^{(0)}$ starting at cycle $t$ is assigned to the robot at loading station $i$, and zero otherwise.
+Variable $\hat{x}_{i}$ equals one if this robot is deferred to the next cycle, and zero otherwise.
+$\mathcal{R}_{i}^{(0)}$ denotes the set of all feasible paths available to loading station $i$ in the current cycle $t_{0}$.
+If a sorting task is pending with destination outlet $k$, then $\mathcal{R}_{i}^{(0)}=\mathcal{R}_{i,k}$; otherwise, $\mathcal{R}_{i}^{(0)}=\emptyset$.
+In the objective function, $c_{i}^{r}$ denotes the completion time of path $r$, and $\hat{c}_{i}$ denotes the penalty for delaying the assignment of a feasible path to a loaded robot until the next cycle.
+This penalty is introduced to prevent robots from being assigned paths with excessively long completion times.
+Constraints (9) enforce the spatio-temporal occupancy limits of each node $\nu$, where $N_{\nu}^{l}$ denotes the remaining capacity of node $\nu$ in cycle $l$.
+$N_{\nu}^{l}$ is dynamically updated based on the FPA decisions in previous cycles.
+Conflict nodes are allowed to accommodate up to two robots per cycle; all other nodes are restricted to one.
+The parameter $\delta_{i}^{r,t,\nu,l}$ equals one if feasible path $r$ starting at cycle $t$ will occupy node $\nu$ in cycle $l$, and zero otherwise.
+Constraints (10) ensure that exactly one assignment is made for each sorting task, where $d_{i}$ equals one if there is a sorting task pending at loading station $i$.
+Finally, Constraints (11) impose integrality conditions.
+It is essential to solve the FPA problem efficiently within each cycle.
+In large-scale networks, the number of feasible paths grows rapidly, making it challenging to obtain a solution within a reasonable computation time.
+We propose a heuristic algorithm to generate a feasible solution in a relatively short time, as shown in Algorithm 1.
+Algorithm 1 Heuristic Algorithm for Implementing RC-S
+0:  Initialize the reservation table that records the status of VP reservation for the upcoming $N_{record}$ cycles, labeled as unoccupied.
+Set the waiting time of each robot that is waiting to be assigned a feasible path as 0.
+Set the maximum search range in each cycle as $N_{search}$.
+Pre-compute the set of feasible path patterns with less than three turns as $\mathcal{R}=\bigcup_{i\in\mathcal{L},\ k\in\mathcal{O}}\mathcal{R}_{i,k}$.
+Initialize the candidate robot list as $\mathcal{S}_{r}=\emptyset$.
+Set cycle index $l=1$.
+0:  At the beginning of cycle $l$, for each loading station, if the first robot in the queue has not been assigned a feasible path, add it to the candidate list $\mathcal{S}_{r}$.
+0:  If $\mathcal{S}_{r}$ is empty, go to Step 5.
+Otherwise, select the robot $a$ with the longest waiting time from $\mathcal{S}_{r}$.
+Obtain its origin loading station $i$ and target outlet $k$.
+Search the shortest feasible path with the earliest entry cycle for robot $a$: for future cycle $t\in[l,l+N_{\text{search}}]$:    for feasible path $r\in\mathcal{R}_{i,k}$ (sorted by path duration in ascending order):       Check if reserving path $r$ starting at cycle $t$ is feasible under current reservation table.
+If feasible, record $(r,t)$ as the assigned path for robot $a$, and go to Step 3.
+end forend for
+0:  If a feasible pair $(r,t)$ was recorded in Step 2, then the VP-cycle occupation of this assignment is given by $\{(p_{i},l_{i}+t)\}$.
+Update the reservation table with $\{(p_{i},l_{i}+t)\}$, and assign the path to robot $a$.
+If no such pair is found, increase the waiting time of robot $a$ by $\tau_{c}$.
+0:  Remove robot $a$ from $\mathcal{S}_{r}$, and return to Step 2.
+0:  Add the newly generated VPs in the next cycle into the reservation table, labeled as unoccupied. Remove the VPs that have exited the network.
+0:  Set $l=l+1$. Update the status of all robots and assign parcels to empty robots queueing in loading zone. Return to Step 1.
+In Algorithm 1, we introduce a heuristic search method that iteratively identifies the shortest feasible path with the earliest entry cycle, as a substitute for solving the original FPA model.
+During initialization, the reservation table is constructed with dimensions $n_{h}\times n_{v}\times N_{record}$, since the aisle network consistently contains $n_{h}\times n_{v}$ VPs at any given time.
+The value of $N_{record}$ is required to be greater than the maximum duration of feasible paths to fully capture the impact of each assignment on future spatio-temporal resource availability.
+For each feasible path, we retain only the earliest admissible release cycle, as defined in Definition 2.
+By introducing a maximum search range $N_{search}$ in Step 3, we ensure that the number of pairs of feasible paths and entry cycles is finite, with an upper bound equal to the number of trajectory patterns $\lvert\mathcal{R}\rvert$ multiplied by $N_{search}$.
+In each cycle, VPs move according to the RC-S pattern, accompanied by the generation of new VPs and the departure of others.
+The reservation table must be updated accordingly, as shown in Step 5.
+In Step 6, the path-assignment results for the current cycle are executed by the robots, which then perform movement, parcel drop-off, queueing, and loading within the cycle.
+The algorithm then proceeds to the planning phase of the next cycle.
+[…]
 5 Performance Analysis of RC-S
@@
 A throughput estimation formula is proposed, serving as a constraint for the planning model in Section 7.
-Two important metrics in traffic analysis (Daganzo,, 2010) are the total vehicular distance traveled per second of operation $\bar{m}(veh\cdot m/s)$, and the average travel distance $\bar{l}$.
-The relationship among system throughput $\tilde{T}_{O}$ and these two metrics is shown in equation (11):
+Two important metrics in traffic analysis (DAGANZO2010434) are the total vehicular distance traveled per second of operation, $\bar{m}(veh\cdot m/s)$, and the average travel distance, $\bar{l}$.
+The relationship between system throughput $\tilde{T}_{O}$ and these two metrics is shown in Equation (12):
 $\tilde{T}_{O}=\frac{\bar{m}}{\bar{l}}$ (12)
-Under the mechanism of RC-S, the first metric $\bar{m}$ is numerically consistent with the product of the number of occupied VPs, denoted by $n^{occupied}_{VP}$, and the speed of each VP, denoted by $v_{VP}$:
-$\bar{m}=n^{occupied}_{VP}\cdot v_{VP}$ (12)
-The speed of each VP could be obtained by expression (4)-(5) with the parameters of RC-S.
-For the number of occupied VPs, an intuitive idea is that it is related to the number of workers engaged in loading parcels in the loading zone and the spatial utilization under RC-S.
-We introduce two factors: workforce factor $\kappa$, representing the impact of the number of workers on the quantity of VPs that can be occupied; attenuation factor $\beta$, representing the utilization of VPs by the feasible paths generated by RC-S.
-In addition, the number of available robots $n_{r}$ also limits the upper bound of occupied VPs.
-Above all, the expression for calculating $n^{occupied}_{VP}$ can be formulated as:
-$n^{occupied}_{VP}=\min\{\kappa\cdot\beta\cdot n_{VP},n_{r}\}$ (13)
+Under the mechanism of RC-S, robots are required to follow VPs for movement, so their average speed within the sorting zone is approximately equal to $v_{VP}$.
+Although acceleration and deceleration inevitably occur during turning, we limit the maximum number of turns in each feasible path and mitigate their impact on the average robot speed through the design of the robot turning and VP transfer process, as described in Section 4.
+Therefore, the first metric $\bar{m}$ can be estimated as the product of the number of occupied VPs, denoted by $n^{\text{occupied}}_{VP}$, and the speed of each VP, denoted by $v_{VP}$:
+$\bar{m}\approx n^{\text{occupied}}_{VP}\cdot v_{VP}$ (13)
+The speed of each VP can be obtained from Expressions (6)–(7), given the parameters of RC-S.
+For the number of occupied VPs, an intuitive assumption is that it depends on the number of workers engaged in loading parcels in the loading zone and on the spatial utilization under RC-S.
+We therefore introduce two factors: the workforce factor $\kappa$, representing the impact of the number of workers on the quantity of VPs that can be occupied; and the attenuation factor $\beta$, representing the utilization of VPs by the feasible paths generated by RC-S.
+In addition, the number of available robots $n_{r}$ also limits the maximum number of occupied VPs.
+Combining these effects, the expression for calculating $n^{\text{occupied}}_{VP}$ can be formulated as:
+$n^{\text{occupied}}_{VP}=\min\left\{\kappa\cdot\beta\cdot n_{VP},n_{r}\right\}$ (14)
 where $n_{VP}$ denotes the total number of VPs in the network, which is constant given the scale of the network.
@@
 $n_{VP}=n_{h}\cdot n_{v}$ (15)
-Now we investigate the forms of the two factors.
+We now investigate the forms of the two factors $\kappa$ and $\beta$.
 Workforce factor
-This factor represents the impact of the number of workers on the quantity of VPs that can be occupied.
-For each aisle, the newly generated VPs at the entrance could be occupied by loaded robots when its corresponding loading station is active and worker is engaged in tasks.
-The quantity of such VPs is a critical determinant of the maximum throughput.
+The workforce factor $\kappa$ represents the impact of the number of workers on the number of VPs that can be occupied.
+It is determined by the network size $(n_{h},n_{v})$ and the number of workers $n_{w}$.
+For each aisle, the newly generated VPs at the entrance can be occupied by loaded robots when the corresponding loading station is active and the worker at that station is engaged in tasks.
+The number of such VPs is a critical determinant of the maximum throughput.
 Consider, for instance, a scenario where eight workers are stationed in an RSS with a $6+6$ aisle network.
-To ensure the balance of traffic flow and mitigate congestion at the aisle network’s periphery, workers are are evenly divided into four groups and positioned along the four edges of the site, as depicted in Figure 9.
-Four horizontal aisles and four vertical aisles are served by workers.
-We merge the area covered by these aisles and obtain the blue polygon.
-Figure 9: The impact of number of workers on the VPs that can be occupied
-For ease of discussion, the notation $\alpha$ will be employed to denote the ratio of workers to the network’s maximum worker capacity.
+To ensure balanced traffic flow and mitigate congestion at the periphery of the aisle network, the workers are evenly divided into four groups and positioned along the four edges of the site, as shown in Figure 10.
+Four horizontal aisles and four vertical aisles are served by these workers.
+We merge the areas covered by these aisles and obtain the blue polygon.
+Figure 10: The impact of the number of workers on the VPs that can be occupied
+For ease of discussion, we use $\alpha$ to denote the ratio of the number of workers to the network’s maximum worker capacity.
 $\alpha=\frac{n_{w}}{n_{h}+n_{v}}$ (16)
-The proportion of the blue region, noted as the workforce factor $\kappa$, could be obtained by:
+The proportion of the blue area, referred to as the workforce factor $\kappa$, is derived by:
 $\kappa=1-{(1-\alpha)}^{2}$ (17)
@@
 PROPOSITION 2.
-If either (i) the candidate feasible paths in RC-S include no more than two turns, or (ii) the average travel distance of a task is less than $(2-\alpha)min\{n_{v},n_{h}\}$, then an upper bound of the proportion of occupied VPs is $\kappa$;
-The proof is presented in Online Appendix B. Proposition 1 provides with support for estimating the maximized network throughput. Next, we analyze the utilization rate of VPs using the RC-S scheme.
+If either (i) the candidate feasible paths in RC-S include no more than two turns, or (ii) the average travel distance of a task is less than $(2-\alpha)\min\{n_{v},n_{h}\}$, then an upper bound on the proportion of occupied VPs is $\kappa$.
+The proof is presented in Online Appendix B.
+Proposition 2 provides theoretical support for estimating the maximum network throughput.
+Next, we analyze the utilization rate of VPs under the RC-S scheme.
 Attenuation factor
-This factor represents the utilization of VPs by the feasible paths generated by RC-S.
-We begin with examine the path reservation strategy within the framework of RC-S .
-In a first-come-first-served basis, the system allocates to each newly requested sorting task an available robot, assigning it the currently shortest viable route.
-This robot then proceeds through the network along a predefined spatio-temporal path, ensuring a collision- and queue-free delivery within the aisle network.
-However, this spatio-temporal reservation mechanism obligates the robot to adhere strictly to its assigned spatio-temporal trajectory, precluding any alterations like deceleration or temporary halts.
+The attenuation factor $\beta$ represents the utilization of VPs by the feasible paths generated by RC-S.
+We begin by examining the path reservation strategy within the framework of RC-S.
+On a first-come-first-served basis, the system allocates an available robot to each newly requested sorting task, assigning it the currently shortest viable route.
+This robot then proceeds through the network along a predefined spatio-temporal path, ensuring collision- and queue-free delivery within the aisle network.
+However, this spatio-temporal reservation mechanism obligates the robot to adhere strictly to its assigned spatio-temporal trajectory, precluding any alterations such as deceleration or temporary halts.
 Such inflexibility can lead to underutilization of traffic capacity, as VPs that are momentarily unoccupied cannot always be sequentially linked to construct a continuous path if they do not intersect at the same point in time.
-To quantify the impact of this reduction, we introduce an attenuation factor $\beta$, representing the spatial utilization.
-Figure 10: The ”pile up” process of spatiotemporal occupation under RC
-Figure 11: The loss of capacity in different network
-The attenuation factor $\beta$ rises with the increase of $(n_{h}+n_{v})$.
-Under the RC-S protocol, the aisle’s spatiotemporal occupation showcases a ”pile-up” effect, as depicted in Figure 10.
+To quantify the impact of this reduction, we introduce the attenuation factor $\beta$, representing the spatial utilization.
+Figure 11: The ”pile-up” process of spatio-temporal occupation under RC
+Figure 12: The loss of capacity in different networks
+The attenuation factor $\beta$ increases with $n_{h}+n_{v}$.
+Under the RC-S protocol, the spatio-temporal occupancy of an aisle exhibits a ”pile-up” effect, as depicted in Figure 11.
 The figure’s horizontal axis marks the spatial position within an aisle, while the vertical axis tracks discretized time intervals.
 Each robot’s transit through this path necessitates occupying a VP, visualized as a blue bar in the illustration.
-The collective occupation by different robots generates the ”pile-up” phenomenon.
+The collective occupancy by different robots generates the ”pile-up” phenomenon.
 The capacity loss at any moment is represented by the unoccupied area beneath the upper boundary of this accumulation.
-When comparing networks of varying scales, i.e., those with shorter versus longer aisle lengths, we maintain a continuous approach: assuming that new sorting tasks emerge with starting points evenly spaced along the boundary and target destinations uniformly spread across the network.
-In a stable operation, the flow distribution of paths under consistent control logic remains analogous across different network sizes, with variations only in total path lengths.
+When comparing networks of varying scales, i.e., those with shorter versus longer aisle lengths, we adopt a continuous approach: we assume that new sorting tasks emerge with starting points evenly spaced along the boundary and target destinations uniformly spread across the network.
+In stable operation, the flow distribution of paths under consistent control logic remains analogous across different network sizes, with variations only in the total path lengths.
 Accordingly, the pattern of unoccupied spaces beneath the upper boundary remains consistent across scales.
-Figure 11 illustrates this by comparing capacity losses across aisles of diverse lengths.
-Longer aisles tend to exhibit extended periods of spatiotemporal occupation, potentially exacerbating capacity loss — the unoccupied intervals shown in Figure 11.
-In consequence, we can reasonably infer that the absolute amount of lost capacity increases quadratically with the number of aisles $(n_{h}+n_{v})$, and the proportion of the lost traffic capacity should grow linearly (because the number of available VPs increases linearly with the length of aisle).
-In summary, the expression of attenuation factor $\beta$ is as follows:
+Figure 12 illustrates this by comparing capacity losses across aisles of different lengths.
+Longer aisles tend to exhibit extended periods of spatio-temporal occupancy, potentially exacerbating capacity loss–the unoccupied intervals shown in Figure 12.
+Consequently, we can reasonably infer that the absolute amount of lost capacity increases quadratically with the number of aisles $(n_{h}+n_{v})$, and that the proportion of the lost traffic capacity grows linearly (because the number of available VPs increases linearly with the length of the aisles).
+In summary, the expression of the attenuation factor $\beta$ is as follows:
 $\beta(n_{h},n_{v})=\frac{1}{a+b\cdot(n_{h}+n_{v})}$ (18)
 where $a$ and $b$ are two parameters.
-Parameter $a$ is slightly greater than 1 and $b$ is relatively small, indicating that for a small-scale network, the attenuation factor should be close to 1 because there are fewer conflict points in one feasible path.
-We use simulation to obtain the real performance of RC-S and apply linear regression to calibrate the value of the parameters in the attenuation coefficient.
-Experiments show that a proper value is $a=1.4$, $b=0.012$.
-Note that, due to the presence of the continuity assumption, there is a certain bias in $\beta$ when the number of aisles is small.
-Specifically, in smaller networks, this factor may overestimate the system’s throughput, especially in scenarios where the activation rate of loading stations is low.
+Parameter $a$ is slightly greater than 1 and $b$ is relatively small, indicating that for a small-scale network, the attenuation factor should be close to 1 because there are fewer conflict points in a feasible path.
+We use simulations to obtain the actual performance of RC-S and apply linear regression to calibrate the values of the parameters in the attenuation factor.
+Experiments show that appropriate values are $a=1.4$ and $b=0.012$.
+Note that, due to the continuity assumption, there is a certain bias in $\beta$ when the number of aisles is small.
+Specifically, in smaller networks, this factor may overestimate the system throughput, especially in scenarios where the activation rate of loading stations is low.
 This is because some path shapes may not exist in such networks.
 Experimental validation of this effect will be presented in the subsequent section.
-To this end, we derive the final expression of number of occupied VPs:
-$n^{occupied}_{VP}(\alpha,n_{r})=\min\{\frac{1-{(1-\alpha)}^{2}}{a+b(n_{h}+n_{v})}\cdot n_{h}n_{v},\ n_{r}\}$ (18)
+To this end, we derive the final expression for the number of occupied VPs:
+$n^{\text{occupied}}_{VP}(\alpha,n_{r})=\min\left\{\frac{1-{(1-\alpha)}^{2}}{a+b(n_{h}+n_{v})}\cdot n_{h}n_{v},\ n_{r}\right\}$ (19)
 LEMMA 1.
-$n^{occupied}_{VP}(\alpha,n_{r})$ is non-decreasing and concave for all $\alpha$ and $n_{r}$ in its domain.
-The proof of Lemma 1 can be accomplished by verifying that the partial derivatives of the two terms in the minimum function are non-negative and the Hessian matrices are both semi-negative definite.
-Lemma 1 indicates that opening more loading stations will increase the capacity of VPs within the network, thereby raising the upper limit of total throughput.
+$n^{\text{occupied}}_{VP}(\alpha,n_{r})$ is non-decreasing and concave in both $\alpha$ and $n_{r}$ over its domain.
+The proof of Lemma 1 can be established by verifying that the partial derivatives of the two terms in the minimum function are non-negative and that the Hessian matrices are both negative semi-definite.
+Lemma 1 indicates that opening more loading stations will increase the capacity of VPs within the network, thereby raising the upper limit of the total throughput.
 However, the marginal returns of constructing new loading stations will gradually diminish.
 Average travel distance
-We now derive the average travel distance $\bar{l}$ of a sorting task, as the remaining part of the throughput estimation formula.
-Based on the mechanism of RC-S, an appropriate selection of paths is based on the principles of evenly distributed sorting demand, workload balancing of workers, and minimizing travel distance.
-Workload balancing principle ensures that loading stations do not experience starvation, thus preserving the stability of the system.
-It also aligns better with real-world scenarios.
+We now derive the average travel distance $\bar{l}$ of a sorting task as the remaining component of the throughput estimation formula.
+Under the mechanism of RC-S, appropriate path selection follows the principles of evenly distributed sorting demand, balanced worker workload, and minimized travel distance.
+The workload-balancing principle ensures that loading stations do not experience starvation, thus preserving the stability of the system.
+It also aligns well with real-world scenarios.
 To maintain a balanced flow of robots between loading stations, we arrange the workers at the centers of each side, covering a length equal to $\alpha$ times the side length.
-Figure 12 illustrates this layout, where outlets are divided into four types of areas based on their positions.
-Sorting demand are uniformly and continuously distributed across the entire area.
-As mentioned earlier, each time a robot turns, it occupies additional space in the VP fleet, so it is necessary to restrict the maximum permissible robot’s turning times in one delivery.
+Figure 13 illustrates this layout, where the sorting zone is divided into four types of areas based on the spatial locations of the outlets.
+Sorting demand is assumed to be uniformly and continuously distributed across the entire area.
+As mentioned earlier, each time a robot turns, it occupies additional space in the VP fleet, so it is necessary to restrict the maximum permissible number of turns in one delivery.
 However, from the diagram, it can be observed that to ensure the outlets in Area 3 are reachable from each loading station, paths with at least three turns are required.
-According to our simulation experiments, introducing paths with four or more turns did not effectively improve the sorting throughput.
+According to our simulation experiments, introducing paths with four or more turns does not effectively improve the sorting throughput.
 Therefore, we prioritize paths with fewer turns and do not consider paths that require more than three turns.
-We denote the path length serving area $i$ as $l_{i}$, with the proportion of this area represented by $p_{i}$.
-We will next present the rationale for their calculation and the formulas.
-For a detailed derivation process, please refer to the appendix B.
-Figure 12: Dividing areas according to the layout
-l
+We denote the path length serving Area $i$ by $l_{i}$, with the proportion of this area represented by $p_{i}$.
+We next present the rationale for their calculation and the corresponding formulas.
+For a detailed derivation, please refer to Appendix B.
+Figure 13: Division of areas according to the layout
 Figure 14: Service paths for different categories
-Firstly, for area 1, servicing the outlets requires one change of direction, as shown in Figure 13(a1).
-Based on the balance of workload, the turning points should have an equal probability of occurring within the central area.
-We then perform a weighted accumulation based on the length and then take the average.
-Consequently, we obtain expressions for the area proportion and the expected path length:
+First, for Area 1, servicing the outlets requires one change of direction, as shown in Figure 14(a1).
+Assuming balanced workload, the turning points should have equal probability of occurring within the central area.
+We then perform a weighted accumulation based on the length and take the average.
+Consequently, we obtain the following expressions for the area proportion and the expected path length:
 $p_{1}$ $=\alpha(1-\alpha)$ (20)
-$E[l_{1}]$ $=2D\cdot[\frac{\frac{9+a^{2}}{6}(n_{h}^{2}+n_{v}^{2})-n_{h}n_{v}-\frac{1}{3}}{n_{h}+n_{v}}+1]$ (20)
-For paths that pass by the areas of area 2, their endpoints are located at either the lower or upper loading stations, as shown in Figure 13(b1) and (b2), respectively.
-Let $(x_{1},y_{1})$, $(x_{2},y_{2})$ denote the coordinates of the two turning points in a path.
-$x_{1}$ and $x_{2}$ are uniformly distributed over the length covered by the workers, while the $y_{1}$ and $y_{2}$ are identical and uniformly distributed across the area.
-Notably, when x1 = x2, the path actually does not involve any turns.
-Let the robot choose the nearest top or bottom exit point when making the second turn.
+$E[l_{1}]$ $=2D\cdot\left[\frac{\frac{9+\alpha^{2}}{6}(n_{h}^{2}+n_{v}^{2})-n_{h}n_{v}-\frac{1}{3}}{n_{h}+n_{v}}+1\right]$ (21)
+For paths that serve Area 2, their endpoints are located at either the lower or upper loading stations, as shown in Figure 14(b1) and (b2), respectively.
+Let $(x_{1},y_{1})$ and $(x_{2},y_{2})$ denote the coordinates of the two turning points in a path.
+$x_{1}$ and $x_{2}$ are uniformly distributed over the length covered by the workers, while $y_{1}$ and $y_{2}$ are identical and uniformly distributed across the area.
+Notably, when $x_{1}=x_{2}$, the path actually does not involve any turns.
+We assume that the robot chooses the nearest top or bottom exit point when making the second turn.
 The expressions for $p_{2}$ and $E[l_{2}]$ are as follows:
 $p_{2}$ $=\alpha(1-\alpha)$ (22)
-$E[l_{2}]$ $=2D\cdot[\frac{\alpha(n_{h}^{2}+n_{v}^{2})}{3(n_{h}+n_{v})}-\frac{2}{3\alpha(n_{h}+n_{v})}+\frac{3n_{h}n_{v}}{2(n_{h}+n_{v})}]$ (22)
-For Area 3, as previously mentioned, only paths involving three turns can service it.
+$E[l_{2}]$ $=2D\cdot\left[\frac{\alpha(n_{h}^{2}+n_{v}^{2})}{3(n_{h}+n_{v})}-\frac{2}{3\alpha(n_{h}+n_{v})}+\frac{3n_{h}n_{v}}{2(n_{h}+n_{v})}\right]$ (23)
+For Area 3, as previously mentioned, only paths involving three turns can serve it.
 Area 3 includes four individual small areas.
-The two small areas on the bottom, as their service paths shown in Figure 13(c1), have a length equal to the paths with one turn in Figure 13(a1); the two areas on the top, with service paths depicted in Figure 13(c2), involve two extra segments of path in the y-direction (or in the x-direction when the starting point is on the left or right side).
+The two small areas at the bottom, whose service paths are shown in Figure 14(c1), have a length equal to that of the paths with one turn in Figure 14(a1); the two areas at the top, with service paths depicted in Figure 14(c2), involve two extra segments of path in the $y$-direction (or in the $x$-direction when the starting point is on the left or right side).
 As a result, the expressions for $p_{3}$ and $E[l_{3}]$ are as follows:
 $p_{3}$ $={(1-\alpha)}^{2}$ (24)
-$E[l_{3}]$ $=2D\cdot(\frac{n_{h}+n_{v}}{2}+\frac{(1+\alpha)}{4}\cdot\frac{n_{h}n_{v}}{n_{h}+n_{v}})$ (24)
-Finally, for Area 4, which is in the center, we assume it is equally served by the paths of area 1 and area 2.
+$E[l_{3}]$ $=2D\cdot\left(\frac{n_{h}+n_{v}}{2}+\frac{(1+\alpha)}{4}\cdot\frac{n_{h}n_{v}}{n_{h}+n_{v}}\right)$ (25)
+Finally, for Area 4, which is in the center, we assume that it is equally served by the paths of Areas 1 and 2.
 $p_{4}$ $={\alpha}^{2}$ (26)
@@
 $\bar{l}(n_{h},n_{v},\alpha)=\sum_{i=1}^{4}p_{i}E[l_{i}]$ (28)
-Now, we further investigate the properties of equation (27).
+We now further investigate the properties of Equation (28).
 LEMMA 2.
-For uniformly distributed sorting demands, if the length-to-width ratio is less than 2 and $min\{n_{h},n_{v}\}\geq 4$, (i) as $\alpha$ increases, $\bar{l}(n_{h},n_{v},\alpha)$ decreases initially and then increases; (ii) when the site size is fixed ($n_{h}\cdot n_{v}=\textit{constant}$), a square-shaped site has a smaller average travel distance; (iii) when the site is square-shaped, the range of average travel distance could be obtained by $2D\cdot n_{h}<\bar{l}(n_{h},n_{v},\alpha)<2D\cdot\frac{9}{8}n_{h}$.
-Proofs for Lemma 2 is provided in Appendix B.
-The condition within the lemma indicates that the conclusion is applicable to an RSS where the length-to-width ratio does not exceed two.
-This is a reasonable setup in practice, as excessively elongated sites accommodate fewer outlets under the same area.
-Lemma 2(i) states that the efficiency of sorting by robots exhibits a trend of first increasing and then decreasing with the increase in the number of workstations.
-This is because, under the worker load balancing criterion, the average distance from the corner loading stations to each outlet is greater than the stations located in the middle of the four sides.
-Lemma 2(ii) suggests that, under the permissible conditions, designing the aisle network in a square shape reduce the demand for robots.
-As aforementioned in section 3, the density of outlets remains constant, thus the distribution of outlets in RSS does not need to be altered as well.
-Lastly, Lemma 2(iii) provides the range of average distance to be covered in a single delivery within RC-S.
-This will serve as a foundation for the efficiency analysis in the following sections.
-Combining equations (15)-(27), the estimation formula of sorting throughput under given configuration $(n_{h},n_{v},n_{w},n_{r})$ is as follows:
-$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})=\frac{D\cdot n^{occupied}_{VP}(\frac{n_{w}}{n_{h}+n_{v}},n_{r})}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{w}}{n_{h}+n_{v}})}$ (28)
-If we fix the values of $n_{h}$ and $n_{v}$, the upper bound of system throughput could be calculated by setting the value of $n_{r}$ to infinity and assigning workers to all loading stations:
-$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})\leq\frac{D\cdot n^{occupied}_{VP}(\frac{n_{l}}{n_{h}+n_{v}},\infty)}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$ (29)
-We use $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ to denote this upper bound in the right side of inequality (29). Here we investigate the properties of this bound.
+For uniformly distributed sorting demands, if the length-to-width ratio is less than 2 and $\min\{n_{h},n_{v}\}\geq 4$, then (i) as $\alpha$ increases, $\bar{l}(n_{h},n_{v},\alpha)$ first decreases and then increases; (ii) when the site size is fixed ($n_{h}\cdot n_{v}=\text{constant}$), a square-shaped site has a smaller average travel distance; (iii) when the site is square-shaped, the range of average travel distance is given by $2D\cdot n_{h}<\bar{l}(n_{h},n_{v},\alpha)<2D\cdot\frac{9}{8}n_{h}$.
+The proof of Lemma 2 is provided in Appendix B.
+The conditions in the lemma indicate that its conclusions is applicable to an RSS where the length-to-width ratio does not exceed two.
+This is a reasonable setup in practice, as excessively elongated sites accommodate fewer outlets for the same area.
+Lemma 2(i) states that the sorting efficiency of robots exhibits a trend of first increasing and then decreasing with the increase as the number of loading stations increases.
+This is because, under the worker load-balancing criterion, the average distance from the corner loading stations to each outlet is greater than that from loading stations located in the middle of the four sides.
+Lemma 2(ii) suggests that, under the permissible conditions, designing the aisle network in a square shape can reduce the demand for robots.
+As mentioned in Section 3, the outlet density remains constant, thus the distribution of outlets in RSS does not need to be altered.
+Lastly, Lemma 2(iii) provides the range of the average distance to be covered in a single delivery within RC-S.
+This serves as a foundation for the efficiency analysis in the following sections.
+Combining Equations (16)–(28), the estimation formula for the sorting throughput under a given configuration $(n_{h},n_{v},n_{w},n_{r})$ is as follows:
+$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})=\frac{D\cdot n^{\text{occupied}}_{VP}(\frac{n_{w}}{n_{h}+n_{v}},n_{r})}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{w}}{n_{h}+n_{v}})}$ (29)
+If we fix the values of $n_{h}$ and $n_{v}$, the upper bound of the system throughput can be obtained by setting $n_{r}$ to infinity and assigning workers to all loading stations:
+$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})\leq\frac{D\cdot n^{\text{occupied}}_{VP}(\frac{n_{l}}{n_{h}+n_{v}},\infty)}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$ (30)
+We use $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ to denote this upper bound on the right-hand side of Inequality (30). We now investigate the properties of this bound.
 PROPOSITION 3.
-$\frac{\partial\tilde{T}_{M}(\cdot)}{\partial n_{l}}$ has one zero point within the valid range of $n_{l}$.
-Denote the value as $n_{l}^{(0)}$, then $\tilde{T}_{M}(\cdot)$ is monotonically increasing in the interval $(0,n_{l}^{(0)}]$ and monotonically decreasing in the interval $(n_{l}^{(0)},n_{h}+n_{v}]$.
-The proof of Proposition 2 can be derived by incorporating Lemma 1 and Lemma 2, which are depicted in detail in Appendix B.
-Proposition 2 reveals that under the RC-S robot control scheme, the maximum throughput has a critical point, and this critical point is not achieved when all loading stations are activated, due to workload balancing and uniformly distributed demands.
-When the proportion of active loading stations approaches 1, the growth rate of $n^{occupied}_{VP}$ will be slower than the growth rate of the service path length, resulting in a decrease in throughput.
-We will verify the formula through experiments in the next section.
+$\frac{\partial\tilde{T}_{M}(\cdot)}{\partial n_{l}}$ has a unique zero point within the valid range of $n_{l}$.
+Denote this value by $n_{l}^{(0)}$, then $\tilde{T}_{M}(\cdot)$ is monotonically increasing on the interval $(0,n_{l}^{(0)}]$ and monotonically decreasing in the interval $(n_{l}^{(0)},n_{h}+n_{v}]$.
+The proof of Proposition 3 can be established by combining Lemma 1 and Lemma 2, and is presented in detail in Appendix B.
+Proposition 3 reveals that under the RC-S robot control scheme, the maximum throughput has a critical point, and that this critical point is not attained when all loading stations are activated, due to workload balancing and uniformly distributed demand.
+When the proportion of active loading stations approaches 1, the growth rate of $n^{\text{occupied}}_{VP}$ becomes slower than the growth rate of the service path length, resulting in a decrease in throughput.
+We will verify this formula through experiments in the next section.
 6 Numerical Validation
 6.1 Comparative analysis: RC-S versus benchmark traffic management framework
-In this section, we compare our proposed RC-S against cooperative A* (CA*) (Silver,, 2021) for traffic management in a warehouse, to better illustrate the superiority of RC-S in terms of AGV service distance and system throughput.
-CA* is a framework based on a simple prioritized-planning scheme: each agent is first assigned a unique priority, and based on the order of priorities, algorithm find the shortest path for each agent that avoids conflicts with agents of higher priority.
-We use SIPP (Phillips and Likhachev,, 2011), an efficient variant of location-time A*, as the lower-level solver of CA*.
-Moreover,other well-known MAPF algorithm frameworks, such as PBS (Ma et al.,, 2019) and RHCR (Li et al.,, 2021), are found less efficient compared to CA* in large-scale settings ($n_{r}>100$) of our experiments, primarily due to the increased likelihood of cycle conflicts within the RSS framework, and thus are excluded from this comparative analysis.
-Table 2: RC-S parameters in simulation
+In this section, we compare our proposed RC-S with cooperative A* (CA*) (Silver_2021) for traffic management in a warehouse, in order to illustrate the superiority of RC-S in terms of robot travel distance and system throughput.
+CA* is a framework based on a simple prioritized-planning scheme: each agent is first assigned a unique priority, and, in order of these priorities, the algorithm finds the shortest path for each agent that avoids conflicts with agents of higher priority.
+We use SIPP (phillips2011sipp), an efficient variant of location-time A*, as the lower-level solver of CA*.
+Moreover, other well-known MAPF frameworks, such as PBS (ma2019searching) and RHCR (Li_Tinka2021), are found to be less efficient than CA* in large-scale settings ($n_{r}>100$) of our experiments, primarily due to the increased likelihood of cycle conflicts within the RSS framework, and are thus excluded from this comparative analysis.
+Table 3: RC-S parameters used in the simulations
 $D$ (m) $\tau_{e}$ (s) $\tau_{c}$ (s) $v_{VP}$ (m/s)
 1 0.5 2 2
-We compare the performance in two scenarios, 12+12 aisles and 20+20 aisles.
+We compare the performance in two scenarios: 12+12 aisles and 20+20 aisles.
 To ensure uniformity in the experimental setup, we use directed maps in all tests to prevent swapping conflicts.
-We activate all possible loading station locations for the experiments, and parcel targets are evenly distributed among outlets.
+All potential loading-station locations are activated in the experiments, and parcel destinations are evenly distributed among outlets.
 The RC-S parameters, detailed in Table 3, are based on sorting robot data from the Geekplus company.
@@
 The simulations do not consider acceleration and deceleration phases.
-We conduct 50 repeated experiments for each scenario.
-In each experiment, we set a warm-up period of 30 minutes and then record the system performance for 60 minutes of continuous operations.
-Our algorithms and simulations are coded in Python 3.11, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM.
+We conduct 50 independent experiments for each scenario.
+In each experiment, we set a warm-up period of 30 minutes and then record the system performance for 60 minutes of continuous operation.
+Our algorithms and simulations are implemented in Python 3.11, and all experiments are conducted on a personal computer running Windows 11 with an Intel i7-9700F CPU and 16 GB of RAM.
 All processes are run single-threaded.
 Table 4: Results on 12+12 aisle network
-$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time per cycle(ms)
+$n_{r}$ Throughput ($\times 10^{3}$/h) Service time (s) Run-time per cycle(ms)
 RC-S CA* RC-S CA* RC-S CA*
@@
 Table 5: Results on 20+20 aisle network
-$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time per cycle (ms)
+$n_{r}$ Throughput ($\times 10^{3}$/h) Service time (s) Run-time per cycle (ms)
 RC-S CA* RC-S CA* RC-S CA*
@@
 400 33.66 24.37 23.41 23.92 37.40 363.45
-Tables 3 and 4 report the system throughput, average service time and average run-time.
-The comparative results indicate that the run-time of RC-S is consistently lower than that of CA*, and in larger-scale cases, it is even less than $1/10$ of the latter.
+Tables 4 and 5 report the system throughput, average service time, and average run-time.
+The comparative results indicate that the run-time of RC-S is consistently lower than that of CA*, and in large-scale cases, it can be less than one tenth of that of CA*.
 Moreover, RC-S always assigns shorter service paths to robots, resulting in an average service time reduction of 10.3% compared to CA*.
-In most experiments, RC-S demonstrates superior throughput compared to CA*, with the exception observed in a 20+20 aisles scenario with $n_{r}=400$.
-This observation underscores the suitability of the RC-S algorithm for environments like RSS, characterized by a compact network and numerous outlets as obstacles.
-The difference in algorithm performance is primarily due to the following improvements: (1) RC-S optimizes by selecting from candidate spatio-temporal paths, while CA* and similar cell-based search algorithms consider variable waiting times at each cell and allow robots to return to previously visited cells; (2) RC-S limits the total number and maximum density of robots and is free of conflicts, while high robot density around outlets in benchmark algorithm leads to significant resource expenditure on conflict resolution.
+In most experiments, RC-S demonstrates higher throughput than CA*, with the only exception observed in the 20+20–aisle scenario with $n_{r}=400$.
+This observation underscores the suitability of the RC-S algorithm for RSS, which is characterized by a compact network and numerous outlets acting as obstacles.
+The difference in algorithm performance is primarily due to the following factors: (1) RC-S optimizes by selecting from candidate spatio-temporal paths, whereas CA* and similar cell-based search algorithms consider variable waiting times at each cell and allow robots to return to previously visited cells; and (2) RC-S imposes limits on both the total number and the maximum density of robots and remains conflict-free, whereas the high robot density around outlets in the benchmark algorithm leads to significant resource expenditure on conflict resolution.
 6.2 Throughput estimation formula validation
-This section validates the throughput estimation formula proposed in section 5 across different active ratios of loading stations.
-We consider 5 scenarios with different network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively.
-The number of robots is set as the number of activated VPs in the aisle network, plus 5 per loading station to ensure the stability of operation.
-Other experimental parameters are specified in section 6.1 in detail.
+This section validates the throughput estimation formula proposed in Section 5 for different active ratios of loading stations.
+We consider five scenarios with different network sizes: $n_{h}=n_{v}=12,14,16,18,20$, respectively.
+The number of robots is set to be equal to the number of activated VPs in the aisle network, plus five robots per loading station to ensure stable operation.
+Other experimental parameters are specified in detail in Section 6.1.
 Figure 15: Analytical formula compared with simulation results.
-Figure 15: Robot flow distribution in simulation, $n_{h}=n_{v}=20$
-Figure 14 shows the trend of the average service distance as a function of the number of active workstations when $n_{h}=n_{v}=20$.
-It indicates that the formula and simulation results present the similar trend of first decreasing and then increasing.
+Figure 16: Robot flow distribution in simulation, $n_{h}=n_{v}=20,n_{w}=40$
+Figure 15 shows the trend of the average service distance as a function of the number of active loading stations when $n_{h}=n_{v}=20$.
+It indicates that the formula and the simulation results exhibit a similar trend of first decreasing and then increasing.
 This suggests that the service frequency of each robot is highest when about half of the loading stations are activated.
-Figure 14 shows the relative error between formula and simulation results under five scenarios.
-It can be seen that the error is always within ±$\pm 8\%$, indicating a good fit.
-Figure 14 shows the trend of the system throughput as a function of the number of activated workstations when $n_{h}=n_{v}=20$.
-The trend demonstrates a clear marginal effect of the number of activated loading stations: when the system is operating at the maximum capacity, there is a certain degree of resource waste.
-Figure 14 shows the relative error between formula and simulation results for five cases, which remains within ±$\pm 15\%$.
-Figure 15 depicts the average traffic flow distribution during a stable operation of the system in simulation with the maximum number of workers.
+Figure 15 shows the relative error between the formula and the simulation results under five scenarios.
+It can be seen that the error is always within $\pm 8\%$, indicating a good fit.
+Figure 15 shows the trend of the system throughput as a function of the number of activated loading stations when $n_{h}=n_{v}=20$.
+The trend demonstrates a clear marginal effect of the number of activated loading stations: when the system is operating near its maximum capacity, there is a certain degree of resource waste.
+Figure 15 shows the relative error between the formula and the simulation results for five cases, which remains within $\pm 15\%$.
+Figure 16 depicts the average traffic flow distribution during stable operation of the system in simulations with the maximum number of workers.
 The darker the color of the aisle segments, the higher the traffic load.
-We notice that the corner loading stations have almost no incoming or outgoing traffic flow.
-This indicates that Algorithm 1, in practical execution, cannot effectively balance the distribution of robots in the loading stations in the extreme situation.
-6.3 Efficiency underestimating in queueing models
-Queueing models have been widely used in analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different work pieces (John A.
-Buzacott,, 1993).
+We observe that the corner loading stations have almost no incoming or outgoing traffic flow.
+This indicates that Algorithm 1, in practical execution, cannot effectively balance the distribution of robots over loading stations in this extreme situation.
+6.3 Underestimation of efficiency in queueing models
+Queueing models have been widely used in the analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different workpieces (buzacott1993stochastic).
 In robotic systems, robots function similarly to pallets in traditional manufacturing systems, transferring workpieces from one machine to another for processing.
-However, we find that queueing models could not capture the efficiency of robots accurately when the number of robots has not reached the network’s capacity limit.
-We apply the CQN model established in Zou et al., (2021), where a two-tier RSS model with closest Robot-to-Loading-Station assignment rule is modified to meet the assumptions in section 3.
-Eight scenarios of large-scale network with few robots are tested, using the same experimental setup as in section 6.2.
-The results are shown in Table 5, where Eq.(28) represents the proposed estimation formula.
+However, we find that queueing models could not accurately capture robot efficiency when the number of robots has not reached the network’s capacity limit.
+We apply the CQN model established in zou2021robotic, in which a two-tier RSS with a closest Robot-to-Loading-Station assignment rule is modified to meet the assumptions in Section 3.
+Eight scenarios of large-scale networks with few robots are tested, using the same experimental setup as in Section 6.2.
+The results are shown in Table 6, where Eq.
+(29) represents the proposed estimation formula.
 The CQN model significantly underestimates system throughput in scenarios with few robots.
 It lacks the capability to accurately depict traffic capacity under low flow conditions with RC-S control and overestimates the impact of conflicts under low-demand conditions, thereby underestimating individual robot efficiency.
-Although the CQN model provides accurate estimates when the network approaches capacity, it results in unnecessary expenses for surplus robots during off-peak periods.
+Although the CQN model provides accurate estimates when the network approaches capacity, it leads to unnecessary costs due to surplus robots during off-peak periods.
 Table 6: Error in throughput estimation under RC-S control
@@
 $n_{r}$ 50 100 200 300 50 100 200 300
-Throughput (×$\times 10^{3}$/h) Simulation 8.43 16.25 28.60 32.90 7.27 14.30 27.01 36.49
+Throughput ($\times 10^{3}$/h) Simulation 8.43 16.25 28.60 32.90 7.27 14.30 27.01 36.49
 CQN 5.98 11.87 22.98 31.47 4.96 9.88 19.52 28.61
@@
 Eq.(29) 7.12 3.65 -9.52 -2.92 9.92 8.41 3.04 -7.10
-we also conduct a experiments on a real case of Deppon Express described in Zou et al., (2021), where the network scale is $n_{h}+n_{v}=18+6$, with 108 outlets, 6 loading stations and 170 robots.
-The The aisle width $D$, maximum robot speed $v_{max}$, and worker loading rate $r_{l}$ were set according to the data provided by Deppon Express.
-We compared the performance of RC-S with the results from Zou et al., (2021), as shown in Table 6, where $T_{D}$, $T_{CQN}$ and $T_{RCS}$ represent the throughput of the Deppon system, the CQN model and the simulation under RC-S scheme, respectively.
-Combined with the experimental results in section 6.1, RC-S scheme significantly improves the operational efficiency of robots, thereby enhancing the overall system throughput.
-Meanwhile, the CQN model effectively captures the congestion in the RSS system implemented by Deppon Express, but underestimates the traffic capacity of the network when the RC-S scheme is applied.
+We also conduct an experiment on a real case of Deppon Express described in zou2021robotic, where the network scale is $n_{h}+n_{v}=18+6$, with 108 outlets, 6 loading stations, and 170 robots.
+The aisle width $D$, maximum robot speed $v_{max}$, and worker loading rate $r_{l}$ are set according to the operational data provided by Deppon Express.
+We compare the performance of RC-S with the results from zou2021robotic, as shown in Table 7, where $T_{D}$, $T_{CQN}$, and $T_{RCS}$ represent the throughput of the Deppon system, the CQN model and the simulation under the RC-S scheme, respectively.
+Together with the experimental results in Section 6.1, this real-case study shows that the RC-S scheme significantly improves the operational efficiency of robots, thereby enhancing the overall system throughput.
+Meanwhile, the CQN model effectively captures congestion in the RSS system implemented by Deppon Express but underestimates the traffic capacity of the network when the RC-S scheme is applied.
 Table 7: Real case validation
@@
 6.4 Evaluating RC-S across diverse scenarios
-In real sorting centers, due to cost constraints and demand arrival rates, loading stations are often not spread across the entire perimeter of the RSS, resulting in only partial availability of network entrances and exits.
-The placement of outlets is also restricted by site factors, such as load-bearing columns or the presence of certain equipment.
-Additionally, demand arrival rates are heterogeneous and are not uniformly distributed across each outlet.
+In real sorting centers, due to cost constraints and demand arrival rates, loading stations are often not distributed across the entire perimeter of the RSS, resulting in only partial availability of network entrances and exits.
+The placement of outlets is also restricted by site-specific factors, such as load-bearing columns or the presence of certain equipment.
+Additionally, demand arrival rates are heterogeneous and are not uniformly distributed across outlets.
 In this section, we evaluate the performance of RC-S across various application scenarios and test the impact of different system configurations on efficiency by simulation.
-For each set of experiments, we evaluate the overall performance across 50 test runs.
-Each run includes a 30-minute warm-up period, followed by an observation period where average data is recorded over one hour.
-We first compared the impact of different numbers of active loading stations on average throughput.
-We consider five scenarios with varying network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively.
-The number of robots is set to be sufficient in each scenario, ensuring that it does not become a bottleneck for sorting process.
+For each set of experiments, we evaluate overall performance across 50 independent runs.
+Each run includes a 30-minute warm-up period, followed by an observation period during which average performance is recorded over one hour.
+We first compare the impact of different numbers of active loading stations on average throughput.
+We consider five scenarios with varying network sizes: $n_{h}=n_{v}=12,14,16,18,20$, respectively.
+The number of robots is set to be sufficient in each scenario, ensuring that it does not become a bottleneck for the sorting process.
 Other RC-S parameters are specified in detail in Table 3.
 Figure 17 shows the trend of throughput across the five scenarios.
-It can be observed that large-scale networks not only have a higher throughput upper bound, but their slope—representing the average efficiency per loading station—is also generally higher.
+It can be observed that large-scale networks not only have a higher upper bound on throughput but also exhibit steeper slopes, representing higher average efficiency per loading station.
 This is because, under RC-S, larger networks provide more space to distribute incoming traffic across different aisles, thereby reducing congestion at each entrance.
-Figure 16: The effect of different numbers of active loading stations
+Figure 17: Effect of different numbers of active loading stations on throughput
 (a) Division of outlet areas in the network
@@
 Figure 18: The effect of closing outlets in the network
-We next evaluate the effect of closing some outlets on system throughput.
-The network size is set to $n_{h}=n_{v}=20$, with other experimental parameters remaining constant.
+We next evaluate the effect of closing a subset of outlets on system throughput.
+The network size is set to $n_{h}=n_{v}=20$, with other experimental parameters kept constant.
 We consider four scenarios for closing outlets: (1) randomly closing outlets; (2) closing outlets in the corners; (3) closing outlets along the edges; (4) closing outlets in the center.
-To control variates, the number of closed outlets in each scenario is set to 1/9 of the total.
+To control for other factors, the number of closed outlets in each scenario is set to one ninth of the total.
 Scenarios (2), (3), and (4) correspond to Areas 1, 2, and 3 in Figure 18(a), respectively.
-Results from 50 repeated experiments for each scenario are shown in Figure 17(b).
+Results from 50 independent experiments for each scenario are shown in Figure 18(b).
 The green dashed line represents the average throughput when all outlets are open.
-It can be observed that randomly closing outlets has no significant impact on system throughput.
+It can be observed that randomly closing outlets does not have a significant impact on system throughput.
 Closing outlets in Areas 1 and 2 results in a decrease in average throughput, while closing outlets in Area 3 leads to a slight increase in average throughput.
-It is intuitive that closing outlets in a specific area causes an uneven distribution of demand, leading to imbalanced network traffic and increased traffic load in other areas of the network.
+Intuitively, closing outlets in a specific area causes an uneven distribution of demand, leading to imbalanced network traffic and increased traffic load in other areas of the network.
 However, from another perspective, it reduces robot turns within the closed area, thereby improving the utilization of spatio-temporal resources.
-Since most shortest paths pass through the central area, closing central outlets in scenario 4 results in higher throughput compared to the normal case.
+Since most of the shortest paths pass through the central area, closing central outlets in scenario 4 results in higher throughput compared to the normal case.
 Figure 19: The effect of heterogeneous demand distribution
 In addition, we evaluate the effect of heterogeneous demand distribution on system efficiency.
-In real sorting centers, sorting destinations with higher demand arrival rates are often assigned multiple outlets to minimize differences in the visit frequency for each outlet; however, this cannot entirely eliminate demand heterogeneity.
+In real sorting centers, sorting destinations with higher demand arrival rates are often assigned multiple outlets to minimize differences in the visit frequency between outlets; however, this cannot completely eliminate demand heterogeneity.
 We assume there are two types of outlets in the RSS: one type has an average demand arrival rate twice that of the other, with both following a normal distribution.
@@
 The two types of outlets are spatially uniformly distributed throughout the network.
-Keeping other experimental parameters constant, we adjust the proportion of the two types of outlets, and the system performance are shown in Figure 18.
-The results indicate that with a fixed total demand and varying proportions of high-demand outlets, RC-S maintains efficiency, and the impact of demand heterogeneity on system throughput is negligible.
-This is achieved by leveraging demand arrival rates for different destinations and balancing the visit frequency to each area, ensuring a well-distributed traffic flow across the network.
-Table 7: RC-S performance in large-scale system
+Keeping other experimental parameters constant, we adjust the proportion of the outlet types, and the resulting system performance is shown in Figure 19.
+The results indicate that, with a fixed total demand and varying proportions of high-demand outlets, RC-S maintains its efficiency, and the impact of demand heterogeneity on system throughput is negligible.
+This is achieved by leveraging demand arrival rates for different destinations and balancing the visit frequency to each area, ensuring a well-distributed traffic flow throughout the network.
+Table 8: RC-S performance in large-scale systems
 $n_{h},n_{v}$ Throughput Avg. path Avg. time Max. time # Feasible paths
@@
 30 5.517 37.261 0.1004 0.2820 406,856
-[…]
+Finally, we evaluate the performance of RC-S in large-scale systems.
+The experimental scenarios include five large-scale networks with $n_{h}=n_{v}=22,\ 24,\ 26,\ 28$, and $30$, respectively.
+Among these, the $30+30$ network covers an actual area of over 5,000 square meters, significantly larger than most sorting systems in modern warehouses.
+For example, a real case from Deppon Express mentioned in the study by zou2021robotic has a network with a scale of $18+6$, approximately one-quarter of the $30+30$ scenario (assuming the same aisle width and loading station space occupation).
+The results are shown in Table 8, including system throughput, the average and maximum run-time per cycle, and the number of precomputed feasible paths associated with the network size.
+As the network scale increases, the total number of feasible paths grows, and the path selection time per cycle gradually increases.
+However, the maximum run-time remains acceptable and is much shorter than the length of each cycle.
 7 Layout Design Optimization
-The layout design problem aims at seeking for a balance between facility costs and long-term operating costs.
+The layout design problem aims to balance facility costs and long-term operating costs.
 A larger facility will achieve higher maximum throughput; however, it will also increase construction and equipment costs.
-Considering transportation costs and supply stability, we analyze the following scenario: The warehouse signs a long-term RaaS contract with a robotics company, with each year divided into several operating periods denoted by $\sigma$.
+Considering transportation costs and supply stability, we analyze the following scenario: A warehouse signs a long-term RaaS contract with a robotics company, with each year divided into several operating periods denoted by $\sigma$.
 The proportion of each period’s duration within the year is represented by $\theta^{\sigma}$, and the planned throughput level during each period is $T^{\sigma}$.
-The facility costs $C_{f}$ and operations costs $C_{o}$ are calculated as follows.
+The facility cost $C_{f}$ and expected operations cost $C_{o}$ are calculated as follows:
 $C_{f}=P_{s}\cdot[2D\cdot(n_{h}-1)+W_{w}+W_{l}]\cdot[2D\cdot(n_{v}-1)+W_{w}+W_{l}]+P_{l}\cdot n_{l}$ (31)
 $C_{o}=\sum_{\sigma\in\mathcal{S}}\theta^{\sigma}(P_{w}n_{w}^{\sigma}+P_{r}n_{r}^{\sigma})$ (32)
-Parameter $P_{s}$ and $P_{l}$ in equation (30) denote the discounted site rental cost per square meter and equipment cost per loading station, while $P_{w}$ and $P_{r}$ in equation (31) denote the discounted unit labor cost and unit rental cost of robot.
+Parameters $P_{s}$ and $P_{l}$ in Equation (31) denote the discounted site rental cost per square meter and equipment cost per loading station, respectively, while $P_{w}$ and $P_{r}$ in equation (32) denote the discounted unit labor cost and the unit rental cost of a robot.
 The objective of the problem is modeled as follows:
-$C_{d}(T^{\sigma},\theta^{\sigma},N_{o})=C_{f}+C_{o}$ (32)
-where parameter $N_{o}$ denotes the minimum number of outlets that should be covered.
+$C_{d}=C_{f}+C_{o}$ (33)
 The trade-off between these two types of costs in the objective function is the focus of the site planning stage.
-We formulate the layout design problem (LDP) as following integer programming model.
+We formulate the layout design problem (LDP) as the following integer programming model:
 $\underset{n_{h},n_{v},n_{l},n_{w}^{\sigma},n_{r}^{\sigma}}{\min}\ C_{d}$
 […]
-$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{\sigma},n_{r}^{\sigma})\geq T^{\sigma}\qquad\forall\sigma\in\mathcal{S}$ (33)
+$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{\sigma},n_{r}^{\sigma})\geq T^{\sigma},\qquad\forall\sigma\in\mathcal{S}$ (34)
 $(n_{h}-1)(n_{v}-1)\geq N_{o}$ (35)
 $n_{l}\leq n_{h}+n_{v}$ (36)
-$n_{w}^{\sigma}\leq n_{l}\qquad\forall\sigma\in\mathcal{S}$ (36)
-$n_{h}=2k_{h}\ ,\quad k_{h}\in\mathcal{Z}_{+}$ (37)
-$n_{v}=2k_{v}\ ,\quad k_{v}\in\mathcal{Z}_{+}$ (38)
-$n_{w}^{\sigma},n_{r}^{\sigma}\in\mathcal{Z}_{+}^{2\left|\mathcal{S}\right|}$ (39)
-Constraint (33) ensures that the system throughput meets predicted sorting requirements in different periods.
-Constraint (34) restricts the lower bound of the site size to accommodate an adequate number of outlets.
-Constraint (35) represents the maximum number of loading stations accommodated in the aisle network under RC-S.
+$n_{w}^{\sigma}\leq n_{l},\qquad\forall\sigma\in\mathcal{S}$ (37)
+$n_{h}=2k_{h}\ ,\qquad k_{h}\in\mathcal{Z}_{+}$ (38)
+$n_{v}=2k_{v}\ ,\qquad k_{v}\in\mathcal{Z}_{+}$ (39)
+$n_{w}^{\sigma},n_{r}^{\sigma}\in\mathcal{Z}_{+},\qquad\forall\sigma\in\mathcal{S}$ (40)
+Constraint (34) ensures that the system throughput meets the sorting requirements in different periods.
+Constraint (35) imposes a lower bound on the site size to accommodate an sufficient number of outlets $N_{o}$.
+Constraint (36) represents the maximum number of loading stations that can be accommodated in the aisle network under RC-S.
 Constraint (37) limits the maximum number of workers to be stationed.
-Constraints (37)-(38) ensure the number of aisles even.
-The remaining constraint (39) requires the variables to be positive integers.
+Constraints (38)-(39) ensure that the numbers of aisles are even.
+The remaining constraint (40) requires the decision variables to be positive integers.
 It is difficult to solve LDP because of the non-linearity of the inequality constraints, especially for large-scale problems.
-In this research, we apply the method of penalty successive linear programming (PSLP) to solve the LP-relaxation of LDP, which exhibits good robustness and convergence properties for large-scale problems (Bazaraa et al.,, 2013).
-Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem along with the penalty function, and utilizes the concept of trust region (updated at each iteration) to control the step size.
-To ensure the constraints continuous and differentiable, we introduce additional constraints requiring that the number of robots is less than the number of available VPs, shown as follows.
-$n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\quad\forall\sigma\in\mathcal{S}$ (40)
+In this research, we apply the method of penalty successive linear programming (PSLP) to solve the LP relaxation of LDP, which exhibits good robustness and convergence properties for large-scale problems (bazaraa2013nonlinear).
+Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem augmented with the penalty function, and utilizes the concept of a trust region (updated at each iteration) to control the step size.
+To ensure that the constraints continuous and differentiable, we introduce additional constraints requiring that the number of robots is less than the number of available VPs, given by
+$n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\quad\forall\sigma\in\mathcal{S}.$ (41)
 At the end of the algorithm, the optimal solution of the relaxed problem is converted into a feasible solution for the original problem by rounding up.
-The details of PSLP algorithm are shown in appendix C.
+The details of the PSLP algorithm are shown in Appendix C.
 8 Numerical Examples for Layout Design
-In this section, we conduct a sensitivity analysis on the site rental cost per square meter $P_{s}$, and labor cost per man-month $P_{w}$, focusing on the trend of total costs under different scenarios as the sorting throughput level changes.
+In this section, we conduct a sensitivity analysis on the site rental cost per square meter $P_{s}$ and the labor cost per man-month $P_{w}$, focusing on the trend of total costs under different scenarios as the sorting throughput level changes.
 Due to increased competition in the robot market, prices remain relatively stable across different companies.
-Consequently, we do not investigate variations in robot rental cost during the experiments, nor do we examine equipment cost at loading stations.
-We consider a 5-year investment plan with a monthly interest rate $\gamma_{0}=0.5\%$, and then the discounted costs of different components in LDP are calculated by:
+Consequently, we do not investigate variations in the robot rental cost in the experiments, nor do we examine the equipment cost at loading stations.
+We consider a five-year investment plan with a monthly interest rate $\gamma_{0}=0.5\%$, and the discounted costs of different components in LDP are calculated as:
 $P_{i}=\sum_{t=0}^{60}\frac{M_{i}}{{(1+\gamma_{0})}^{t}},\quad i=s,l,w,r$ (42)
-where $M_{i}$ denotes the unit cost of component $i$ in a month.
+where $M_{i}$ denotes the unit cost of component $i$ per month.
 We estimate the monthly warehousing rental cost using JD Logistics’ financial reports for the first three quarters of 2023.
-Labor cost is sourced from data related to warehousing job recruitment in Beijing, obtained from website for job posts (58.com).
-The data regarding the equipment cost of loading stations and the rental cost of robots are obtained from Geekplus company.
-We assume that the warehouse only expands its sorting capacity during a few major shopping events, meaning there are two typical levels of sorting demand, namely the average off-peak season demand ${T}^{L}$ and the average peak season demand ${T}^{H}$, with ${T}^{L}=0.8\cdot{T}^{H}$.
-The ratios of the two periods are set as $(\theta^{L},\theta^{H})=(5/6,1/6)$, which means that each quarter typically includes a two-week-long shopping event.
-The default values of parameters is shown in table 8 and the configurations of RC-S keeps the same with table 2.
-Our algorithms are coded in Python 3.11 with the solver Gurobi Optimizer 9.0, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM.
+The labor cost is sourced from data related to warehousing job recruitment in Beijing, obtained from a job-posting website (58.com).
+The data regarding the equipment cost of loading stations and the rental cost of robots are obtained from the Geekplus company.
+We assume that the warehouse only expands its sorting capacity during a few major shopping events, meaning that there are two typical levels of sorting demand, namely the average off-peak season demand ${T}^{L}$ and the average peak season demand ${T}^{H}$, with ${T}^{L}=0.8\cdot{T}^{H}$.
+The ratios of the two periods are set as $(\theta^{L},\theta^{H})=(5/6,1/6)$, which means that each quarter typically includes a two-week shopping event.
+The default values of the parameters are shown in Table 9 and the configurations of RC-S remain the same as in Table 3.
+Our algorithms are implemented in Python 3.11 with the solver Gurobi Optimizer 9.0, and all experiments are conducted on a personal computer running Windows 11 with an Intel i7-9700F CPU and 16 GB of RAM.
 All processes are run single-threaded.
-Table 8: Default values of parameters in section 8
+Table 9: Default parameter values in Section 8
 $M_{s}$ (CNY / $m^{2}\cdot$ mo) $M_{l}$ (CNY/mo) $M_{w}$ (CNY/mo) $M_{r}$ (CNY/mo) $W_{w}$ (m) $W_{l}$ (m)
 10 400 5,000 200 5 5
-8.1 Sensitivity Analysis on the site rental cost
+8.1 Sensitivity analysis on site rental cost
 We begin by investigating the impact of site rental cost under two different scenarios: one with $N_{o}=100$ (scenario 1) and another with $N_{o}=400$ (scenario 2).
-The monthly rental cost of site floor space per square meter varies from $10\ CNY$ to $30\ CNY$.
-The results are presented in Figure 19(a) and Figure 19(b).
+The monthly rental cost of floor space per square meter varies from 10 CNY to 30 CNY.
+The results are presented in Figures 20(a) and  20(b).
 The findings reveal that:
@@
 (b) $N_{o}=400$ (scenario 2)
-Figure 19: Comparison of total costs and facility costs proportion on different site rental cost
-Table 9: Results of scenario $M_{s}=10,N_{o}=100$
+Figure 20: Comparison of total costs and facility cost proportions on different site rental costs
+Table 10: Results for scenario $M_{s}=10$, $N_{o}=100$
 ${T}^{H}$ $(n_{h},n_{v})$ $(n_{w}^{L},n_{r}^{L})$ $(n_{w}^{H},n_{r}^{H})$ $C_{f}$ $C_{o}$ $C_{d}$ Proportion of
-(×$(\times{10}^{6})$ (×$(\times{10}^{6})$ (×$(\times{10}^{6})$ site rental cost (%)
+$(\times{10}^{6})$ $(\times{10}^{6})$ $(\times{10}^{6})$ site rental cost (%)
 3000 (10,12) (2,10) (3,13) 0.53 0.67 1.20 38.80
@@
 30000 (30,30) (29,302) (49,401) 3.42 11.72 15.14 15.88
-  • When the target throughput level is low, the total costs in scenario 1 is significantly lower than that in scenario 2.
+  • When the target throughput level is low, the total cost in scenario 1 is significantly lower than that in scenario 2.
 This is because, to accommodate more outlets, the latter requires renting a larger initial space.
 In both scenarios, the curve of the optimal system cost exhibits a turning point $T^{*}$.
-When $T^{H}>T^{*}$, the cost escalation rate increase significantly.
+When $T^{H}>T^{*}$, the cost escalation rate increases significantly.
 The turning point indicates that the initial space in the current scenario is no longer sufficient to accommodate a larger throughput, necessitating an expansion of the aisle network scale.
-In scenario 1, this turning point occurs at $T^{*}=9,000$; whereas in scenario 2, the turning point is at $T^{*}=21,000$.
-  • The proportion of facility costs increases as $M_{s}$ rises, particularly remarkable at lower sorting throughput levels.
+In scenario 1, this turning point occurs at $T^{*}=9{,}000$, whereas in scenario 2, the turning point is at $T^{*}=21{,}000$.
+  • The proportion of facility costs increases as $M_{s}$ increases, and this effect is particularly pronounced at lower sorting throughput levels.
 As the throughput level becomes higher, this proportion gradually decreases and stabilizes after the turning point $T^{*}$ in each case.
-By observing the system configurations in each dataset, we find that this phenomenon arises because, at higher throughput demands, efficiency improvements necessitate expanding the space, which leads to the increase in the site rental cost.
-Table 9 shows the optimal system configuration and corresponding costs in the case of $M_{s}=10$ and $N_{o}=100$.
-When $T^{H}>T^{*}$, the optimal solutions show not only assigning more workers and robots to loading stations but also simultaneously expanding the size of the network to alleviate traffic pressure.
-  • Comparing the data for $T^{H}=15,000$ and $T^{H}=30,000$ in Table 9, the total costs for the latter are nearly three times those for the former.
-This indicates that due to the simultaneous growth of $C_{f}$ and $C_{o}$, the total costs growth rate of the RSS system exceeds the throughput growth rate.
-From the results of study in site rental cost, the derived managerial insights are summarized as follows.
+By examining the system configurations in each dataset, we find that this phenomenon arises because, at higher throughput demands, efficiency improvements necessitate expanding the space, which leads to an increase in the site rental cost.
+Table 10 shows the optimal system configuration and the corresponding costs in the case of $M_{s}=10$ and $N_{o}=100$.
+When $T^{H}>T^{*}$, the optimal solutions not only assign more workers and robots to the loading stations, but also expand the size of the network to alleviate traffic pressure.
+  • Comparing the data for $T^{H}=15{,}000$ and $T^{H}=30{,}000$ in Table 10, the total cost for the latter is nearly three times that for the former.
+This indicates that due to the simultaneous growth of $C_{f}$ and $C_{o}$, the growth rate of the total cost exceeds the growth rate of throughput.
+From the results of the site rental cost analysis, the derived managerial insights are summarized as follows.
 INSIGHT 1.
-The advantage of RSS is evident in its low initial investment in scenarios with both throughput demand and sortation category low.
+The advantage of RSS is evident in its low initial investment in scenarios where both throughput demand and the number of sortation categories are low.
 INSIGHT 2.
-High density lead to decreased efficiency of workers.
-Instead of running RSS at full capacity, it is more cost-effective to appropriately expand the site size and reduce the proportion of activated loading stations.
+High density leads to decreased worker efficiency.
+Instead of running an RSS at full capacity, it is more cost-effective to appropriately expand the site size and reduce the proportion of activated loading stations.
 (a) $N_{o}=100$ (scenario 1)
 (b) $N_{o}=400$ (scenario 2)
-Figure 20: Comparison of total costs and labor costs proportion on different unit labor cost
-8.2 Sensitivity Analysis on the labor cost
-In a similar manner as our previous experiments, we now explore the influence of different labor costs on the total costs in two scenarios with $N_{o}=100$ and $N_{o}=400$.
-The monthly labor cost of each worker varies from $5,000\ CNY$ to $12,000\ CNY$.
-The proportions of labor costs and robot rental costs are both visualized in Figure 20(a) and Figure 20(b).
-Our observations are concluded as follows:
-  • As $T^{H}$ increases, the proportion of labor costs gradually rises and eventually stabilizes.
+Figure 21: Proportions of labor cost and robot rental cost under different unit labor costs
+8.2 Sensitivity analysis on labor cost
+Similarly to our previous experiments, we now explore the influence of different labor costs on the total cost in two scenarios with $N_{o}=100$ and $N_{o}=400$.
+The monthly labor cost per worker varies from 5,000 CNY to 12,000 CNY.
+The proportions of labor cost and robot rental cost are both visualized in Figures 21(a) and 21(b).
+Our observations are summarized as follows:
+  • As $T^{H}$ increases, the proportion of labor cost gradually increases and eventually stabilizes.
 In every scenario, the proportion of labor costs in the optimal solution for $T^{H}=30{,}000$ exceeds 50%.
@@
   • Comparing the two scenarios, it is evident that in larger-scale sorting sites (scenario 2), the initial increase in the proportion of robot rental cost is faster, making it challenging for RSS systems to balance the advantages of high scalability and low cost.
-From the result of study in labor cost, we summarize the following managerial insight:
+From the labor cost analysis, we summarize the following managerial insight:
 INSIGHT 3.
-At lower levels of throughput, facility costs are the primary expense of RSS, whereas at higher levels of throughput, labor costs becomes the primary expense.
-Additionally, the cost structure undergoes notably changes with varying sorting demands.
+At lower levels of throughput, facility cost is the primary expense of RSS, whereas at higher levels of throughput, labor cost becomes the primary expense.
+Additionally, the cost structure undergoes notable changes with varying sorting demand.
 In scenarios with high sorting demand, the proportion and growth rate of robot rental cost become more significant.
 This reveals a distinctive feature of RSS systems: Robots, as carriers of parcels, become less efficient in larger sites because they have to cover increased distances.
-Merely enhancing management methods is insufficient to tackle the challenge of diminishing traffic efficiency.
-Today’s commonly used conveyor sorters, equipped with high-power motors, necessitate substantial initial investments but experience only slight increments in total costs as sorting demands grow (Zou et al.,, 2021).
-Furthermore, the optimal system configurations of a conveyor sorter are not easily affected by fluctuations in the unit price of cost components (Russell and Meller,, 2003), which makes it suitable for long-term operation in large-scale sorting scenarios.
+Merely enhancing management methods is insufficient to address the challenge of diminishing traffic efficiency.
+Today’s commonly used conveyor sorters, equipped with high-power motors, require substantial initial investments but experience only slight increases in total cost as the sorting demand grows (zou2021robotic).
+Furthermore, the optimal system configurations of a conveyor sorter are not easily affected by fluctuations in the unit prices of cost components (russell2003cost), which makes it suitable for long-term operation in large-scale sorting scenarios.
 Therefore, we conclude the following insight regarding the application of RSS:
 INSIGHT 4.
-Due to the traffic issues of robots, RSS is suitable for application in small-scale scenarios, such as distribution centers at the end of the supply chain.
+Due to the traffic limitations of robots, RSS is suitable for application in small-scale scenarios, such as distribution centers at the end of the supply chain.
 9 Conclusion
-This study conducts a comprehensive analysis of the RSS in modern warehouses, encompassing the underlying robot traffic management, efficiency analysis, as well as the cost composition and layout design of the system.
-By incorporating the throughput estimation, this research sheds light on the prediction of performance and resource allocation in modern warehouses or factories that utilize numerous robots, promoting the ongoing development of automation in logistics industry.
+This study conducts a comprehensive analysis of RSS in modern warehouses, encompassing robot traffic management, system efficiency, and the cost composition and layout design of the system.
+By incorporating throughput estimation, this research provides guidance for performance prediction and resource allocation in modern warehouses or factories that utilize numerous robots, promoting the ongoing development of automation in the logistics industry.
 In the operations stage, efficient robot scheduling emerges as a crucial element for unlocking the business value of RSS.
-We propose an innovative RC-S scheme, which serves as a framework for managing a considerable number of robots simultaneously and efficiently.
-We provide a detailed description of the composition of phases and the method for setting parameters in RC-S.
-Additionally, we present a mathematical programming model FPA to minimize the travel costs at each cycle, and a heuristic algorithm as online solver.
-Subsequently, a theoretical analysis of the efficiency of RC-S is conducted, exploring the impact of various system configurations on throughput.
-In the validation section, we first compare our proposed traffic management framework with classical cooperative A* algorithm.
+We propose an innovative RC-S scheme, which serves as a framework for managing a large number of robots simultaneously and efficiently.
+We provide a detailed description of the phase structure and the parameter-setting method in RC-S.
+Additionally, we present a mathematical programming model, FPA, to minimize the travel cost at each cycle, and a heuristic algorithm as an online solver.
+We then conduct a theoretical analysis of the efficiency of RC-S, exploring the impact of various system configurations on throughput.
+In the validation section, we first compare our proposed traffic management framework with the classical cooperative A* algorithm.
 The simulation results indicate that our control method achieves a higher level of performance and computational efficiency.
 Furthermore, we validate the throughput estimation formula.
-Results show that queueing models can experience significant distortion in some scenarios, while our model fits the real traffic flows of robots more precisely.
-In the site planning stage, warehouse managers need to make decisions that involve a trade-off between the initial investment of facility $C_{f}$, and the discounted operations costs $C_{o}$.
-We analyze the cost composition of the RSS, and propose a layout design optimization model LDP that minimizes the total system costs.
-Through a detailed sensitivity analysis examining the cost dynamics of site rental and labor, we investigate cost proportions under distinct throughput levels.
-Specifically, at lower throughput, the rental cost for the site represents a considerable proportion of the total costs.
-However, at higher throughput, the expenditure on labor emerges as the predominant cost element.
-Results also validate the discussions on the properties of our model.
-The key insights can guide managers in understanding the investment and returns of applying the RSS, thereby reducing the total costs of the warehouse.
+The results show that queueing models can produce significant errors in some scenarios, whereas our model fits the real traffic flows of robots more precisely.
+In the site planning stage, warehouse managers need to make decisions that involve a trade-off between initial facility investment $C_{f}$ and discounted operations costs $C_{o}$.
+We analyze the cost composition of RSS and propose a layout design optimization model, LDP, that minimizes total system costs.
+Through a detailed sensitivity analysis examining the cost dynamics of site rental and labor, we investigate cost proportions at different throughput levels.
+Specifically, at lower throughput, the rental cost for the site represents a considerable proportion of the total cost.
+However, at higher throughput, labor expenditure becomes the predominant cost component.
+The results also support our analysis of the model properties.
+Key insights can guide managers in understanding the investment and returns associated with the application of RSS, thus reducing the total cost of the warehouse.
 For future work, there are two directions to explore.
-Firstly, optimizing layout design considering non-uniform demand and various outlet distributions would further improve the model’s effectiveness.
-Secondly, extending the analysis to cover the entire process from order fulfillment to sorting in modern warehouses would provide valuable insights for enhancing system performance.
+First, optimizing layout design by considering non-uniform demand and various outlet distributions would further improve the model’s effectiveness.
+Second, extending the analysis to cover the entire process from order fulfillment to sorting in modern warehouses would provide valuable insights for enhancing system performance.
 Declaration of generative AI and AI-assisted technologies in the writing process
@@
 $\mathcal{G}$ graph constituted of all nodes and links
+$\mathcal{V}$ set of all nodes
 $\mathcal{V}_{c}$ set of conflict nodes
@@
 $\mathcal{O}$ set of all outlets
-$\mathcal{C}$ set of all cycles
-$\mathcal{R}$ set of all feasible path
-$\mathcal{R}(i,j,k)$ set of feasible path connecting loading station $i$ and $j$, passing outlet $k$
-$\mathcal{R}_{i},k$ set of feasible path starting at loading station $i$ and passing by outlet $k$
-^$\hat{\mathcal{R}}_{i}$ set of feasible path ending at loading station $i$
+$\mathcal{C}$ set of all cycles starting from the current cycle
+$\mathcal{R}$ set of all feasible paths
+$\mathcal{R}(i,j,k)$ set of feasible paths connecting loading station $i\in\mathcal{L}$ and $j\in\mathcal{L}$, passing outlet $k\in\mathcal{O}$
+$\mathcal{R}_{i,k}$ set of feasible path patterns starting from loading station $i\in\mathcal{L}$ and passing by outlet $k\in\mathcal{O}$, independent of starting cycle
+$\mathcal{R}_{i}^{(0)}$ Set of feasible paths available to loading station $i$ at the beginning of current cycle, based on the current destination outlet of the pending sorting task.
+If no task is pending, $\mathcal{R}_{i}^{(0)}=\emptyset$
+$\hat{\mathcal{R}}_{i}$ set of feasible paths ending at loading station $i\in\mathcal{L}$
 $\mathcal{I}_{x}$ set of locations of loading stations on the top and bottom sides
@@
 $P_{r}$ cost of robot per vehicle per month
-${T}^{\sigma}$ target throughput level in operating period $\sigma$
-$\theta^{\sigma}$ Ratio of operating period $\sigma$
+${T}^{\sigma}$ target throughput level in operating period $\sigma\in\mathcal{S}$
+$\theta^{\sigma}$ Ratio of operating period $\sigma\in\mathcal{S}$
 $N_{o}$ minimum number of outlets required to meet the sorting category demands
@@
 $c_{max}$ maximum acceleration/deceleration of robots
+$\omega_{r}$ rotation speed of robots
 $r_{l}$ maximum loading rate of a loading station
-$c_{i,k}^{r}$ travel time of feasible path $r\in R_{i,k}$
-$\hat{c}_{i,k}$ penalty of delay of a sorting task from loading station $i$ to outlet $k$
-$\delta_{i,k}^{r,\nu,l}$ incidence between feasible path $r\in R_{i,k}$, node $\nu$ and cycle $l$
-$N_{\nu}^{l}$ remaining capacity in node $\nu$ in cycle $l$
-$d_{i,k}$ sorting demand from loading station $i$ to outlet $k$
+$c_{i}^{r}$ completion time of feasible path $r\in\mathcal{R}_{i}^{(0)}$, measured in number of cycles
+$\hat{c}_{i}$ penalty for delaying the assignment of a feasible path to a sorting task from loading station $i\in\mathcal{L}$ until the next cycle
+$\delta_{i}^{r,t,\nu,l}$ incidence between feasible path $r\in\mathcal{R}_{i}^{(0)}$ starting at cycle $t\in\mathcal{C}$, node $\nu\in\mathcal{V}$ and cycle $l\in\mathcal{C}$
+$N_{\nu}^{l}$ remaining capacity in node $\nu\in\mathcal{V}$ in cycle $l\in\mathcal{C}$
+$d_{i}$ sorting demand from loading station $i\in\mathcal{L}$, $d_{i}\in\{0,1\}$
 Variables
-$x_{i,k}^{r}$ decision on whether feasible path $r\in R_{i,k}$ is reserved in current cycle
-$\hat{x}_{i,k}$ decision on whether the robot located at loading station $i$ should wait until the next cycle, with the target outlet $k$
+$x_{i}^{r,t}$ decision on whether feasible path $r\in\mathcal{R}_{i}^{(0)}$ starting at cycle $t\in\mathcal{C}$ is reserved
+$\hat{x}_{i}$ decision on whether the sorting task from loading station $i\in\mathcal{L}$ should be postponed to the next cycle
 $n_{h},n_{v}$ number of horizontal and vertical aisles in sorting zone
 $n_{l}$ number of loading stations in loading zone
-$n_{w}^{\sigma}$ number of workers in operating period $\sigma$
-$n_{r}^{\sigma}$ number of robots in operating period $\sigma$
+$n_{w}^{\sigma}$ number of workers in operating period $\sigma\in\mathcal{S}$
+$n_{r}^{\sigma}$ number of robots in operating period $\sigma\in\mathcal{S}$
 Appendix B: Proofs
 B.1 Proof of Proposition 1
+Assume that the number of vertical and horizontal aisles in the network is $(n_{v},n_{h})$.
+We illustrate how each VP–cycle reservation in a feasible path can be translated into a set of node–cycle pairs, using the VPs moving along a horizontal aisle as an example.
+For a horizontal aisle $a$, we denote the sequence of nodes along the aisle (from entrance node to exit node) as $\mathcal{V}_{aisle}^{(a)}=\{\nu_{1}^{(a)},\nu_{2}^{(a)},\nu_{3}^{(a)},\ldots\}$.
+The total number of nodes in this aisle is:
+$\lvert\mathcal{V}_{ailse}^{(a)}\rvert=2n_{v}+1$
+Under the RC-S control scheme, each cycle is divided into four phases, and each phase has a duration of $\tau_{e}$.
+A VP advances to the next node along its path after each phase.
+As illustrated in Figure 7 (Section 4), a newly generated VP enters the entrance node of its aisle during the second phase of the cycle in which it is released.
+In the cycle when it exits the network, it reaches the exit node during the second phase.
+Thus, for a VP $p_{0}$ generated in cycle $l_{0}$ along aisle $a$, the set of nodes it occupies in each cycle $l_{i}\in[l_{0},l_{0}+\frac{n_{v}}{2}-1]$ is given by:
+$(p_{0},l_{i})\iff\begin{cases}\{(\nu_{1}^{(a)},l_{i}),(\nu_{2}^{(a)},l_{i}),(\nu_{3}^{(a)},l_{i})\},&\text{if }\ l_{i}=l_{0}\\ \{(\nu_{4(l_{i}-l_{0})}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+1}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+2}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+3}^{(a)},l_{i})\},&\text{if }\ l_{0}<l_{i}<l_{0}+\frac{n_{v}}{2}\\ \{(\nu_{2n_{v}}^{(a)},l_{i}),(\nu_{2n_{v}+1}^{(a)},l_{i})\},&\text{if }\ l_{i}=l_{0}+\frac{n_{v}}{2}\\ \emptyset,&\text{otherwise.}\end{cases}$
+Similarly, for a VP $p_{0}$ moving along a vertical aisle $a$, we have:
+$(p_{0},l_{i})\iff\begin{cases}\{(\nu_{1}^{(a)},l_{i}),(\nu_{2}^{(a)},l_{i}),(\nu_{3}^{(a)},l_{i})\},&\text{if }\ l_{i}=l_{0}\\ \{(\nu_{4(l_{i}-l_{0})}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+1}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+2}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+3}^{(a)},l_{i})\},&\text{if }\ l_{0}<l_{i}<l_{0}+\frac{n_{h}}{2}\\ \{(\nu_{2n_{h}}^{(a)},l_{i}),(\nu_{2n_{h}+1}^{(a)},l_{i})\},&\text{if }\ l_{i}=l_{0}+\frac{n_{h}}{2}\\ \emptyset,&\text{otherwise.}\end{cases}$
+$\square$
+B.2 Proof of Proposition 2
 (i):
@@
 $\square$
-B.2 Derivation of Maximum Travel Distance in Constraint 5
+B.3 Derivation of Maximum Travel Distance in Constraint (7)
 We first examine the acceleration process of the robot after a turn.
@@
 The speed curve $v_{1}(t)$ and the associated travel distance $d_{1}$ are formulated as follows:
-$v_{1}(t)=\left\{\begin{aligned} &c_{max}t,\quad t\leq\tau_{e}+\frac{v_{VP}}{2c_{max}}\\ &v_{VP}+c_{max}(2\tau_{e}-t),\quad\tau_{e}+\frac{v_{VP}}{2c_{max}}<t\leq 2\tau_{e}\\ \end{aligned}\right.$
+[…]
 $d_{1}=\int_{0}^{2\tau_{e}}v_{1}(t)=\frac{v_{VP}^{2}}{4c_{max}}+\tau_{e}v_{VP}+c_{max}\tau_{e}^{2}$
@@
 The speed curve $v_{1}(t)$ and the associated travel distance $d_{1}$ are formulated as follows:
-$v_{2}(t)=\left\{\begin{aligned} &c_{max}t,\quad t\leq\frac{v_{max}}{c_{max}}\\ &v_{max},\quad\frac{v_{max}}{c_{max}}<t\leq 2\tau_{e}-\frac{v_{max}-v_{VP}}{c_{max}}\\ &v_{VP}+c_{max}(2\tau_{e}-t),\quad 2\tau_{e}-\frac{v_{max}-v_{VP}}{c_{max}}<t\leq 2\tau_{e}\\ \end{aligned}\right.$
+[…]
 $d_{2}=\int_{0}^{2\tau_{e}}v_{2}(t)=2\tau_{e}v_{max}-\frac{2v_{max}^{2}-2v_{max}v_{VP}+v_{VP}^{2}}{2c_{max}}$
 Above all, the maximum travel distance is expressed by:
-$d(\tau_{e},v_{VP},v_{max},c_{max})=\left\{\begin{aligned} &d_{1},\quad 2v_{max}-v_{VP}\geq 2\tau_{e}c_{max}\\ &d_{2},\quad 2v_{max}-v_{VP}<2\tau_{e}c_{max}\\ \end{aligned}\right.$
-Similarly, the analysis for the acceleration process before a turn is consistent with that of the deceleration process, yielding the same result.
+[…]
+Similarly, the analysis for the deceleration process before a turn is consistent with that of the acceleration process, yielding the same result.
 $\square$
-B.3 Derivation of Average Travel Distance in Section 5
+B.4 Derivation of Average Travel Distance in Section 5
 Let $I_{x}$ represents the set of x-coordinates of loading stations on the top and bottom sides, while $I_{y}$ represents the set of y-coordinates of loading stations on the left and right sides.
@@
 Area 3: We first calculate the basic average path length.
-In accordance with workload balance, the starting and ending points of the path should be equally likely to fall on each workstation.
+In accordance with workload balance, the starting and ending points of the path should be equally likely to fall on each loading station.
 Then add the detour lengths for the two different scenarios multiplied by their respective probabilities, according to the path allocation rules.
@@
 $=$ $2D\cdot(\frac{n_{h}+n_{v}}{2}+\frac{(1+\alpha)}{4}\cdot\frac{n_{h}n_{v}}{n_{h}+n_{v}})$
-B.4 Proof of Lemma 2
+B.5 Proof of Lemma 2
 We first prove (i). By equation (28), the expression of average travel distance of a sorting task is as follows:
@@
 $\square$
-B.5 Proof of Proposition 2
+B.6 Proof of Proposition 3
 We first obtain the expression of $\tilde{T}_{M}(n_{h},n_{v},n_{l})$:
@@
 Appendix C: Penalty Successive Linear Programming
-The classic penalty successive linear programming (PSLP) enjoys good robustness and convergence properties for large-scale problems (Bazaraa et al.,, 2013).
+The classic penalty successive linear programming (PSLP) enjoys good robustness and convergence properties for large-scale problems (bazaraa2013nonlinear).
 Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem along with the penalty function, and utilizes the concept of trust region (updated at each iteration) to control the step size.
@@
 The subproblem in iteration $k$ are formed as follows:
-[…]
-[…]
+$\textbf{(LP-S)}(\omega_{k},\Delta_{k})$
+$\underset{d}{\min}\ \nabla{C_{d}(\omega_{k})}^{T}\cdot d+\mu\cdot(\sum_{i=1}^{2\left|S\right|+2}y_{i})$
 $s.t.\quad$ $y_{i}\geq g_{i}(\omega_{k})+\nabla{g_{i}(\omega_{k})}^{T}d,\quad i=1,2,\cdots,2\left|S\right|+2$ (C-1)
@@
 $-(n_{h}-1)(n_{v}-1)+N_{o}\leq 0$ (C-4)
-[…]
+$\underset{\sigma}{max}\{n_{w}^{\sigma}\}-n_{h}-n_{v}\leq 0$ (C-5)
 $n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\qquad\forall\sigma\in\mathcal{S}$ (C-6)
 $-\frac{D\cdot n_{r}^{\sigma}}{\tau_{e}\cdot E[l(n_{h},n_{v},\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})]}+{T}^{\sigma}\leq 0\qquad\forall\sigma\in\mathcal{S}$ (C-7)
+To avoid waste of resources, the number of loading station is always equal to the peak number of workers, thus $n_{l}$ is replaced by $\underset{\sigma}{max}\{n_{w}^{\sigma}\}$ in the constraints.
+The details of solving PLSP is shown in Algorithm C-1.
+It could be roughly divided into 2 steps in each iteration: (1) Generate LP-S and obtain the step size $d_{k}$, determine whether to stop; (2) Adjust step bounds for the next iteration.
+Algorithm C-1 The PLSP algorithm to solve model LDP
 […]
-Algorithm C-1 The PLSP algorithm to solve model LDP
-0:    An initial feasible solution $\omega_{0}$onfidence intervals $0\leq\Delta_{LB}\leq\Delta_{1}$arameters $0\leq\rho_{0}\leq\rho_{1}\leq\rho_{2}\leq 1$, $\phi\leq 1$Large enough constant $\mu$Maximum iterations $N_{m}$
 0:  The optimal value of layout decision variables, $\omega$
@@
 3:   repeat
-[…]
+4:    Solve $\textbf{LP-S}(\omega_{k},\Delta_{k})$ to obtain $d_{k}$
 5:    if $C_{EL_{k}}(\omega_{k})-C_{EL_{k}}(\omega_{k}+d_{k})=0$ or $\lceil\omega_{k}+d_{k}\rceil=\omega_{k}$ then